### 25 Mar 2026 v2

# From the Stochastic Embedding Sufficiency Theorem to a Superspace Diffusion Framework

Abstract
The forward derivation of stochastic differential equations in individual physical domains (Brownian motion, quantum mechanics, chemical kinetics) has proceeded independently for over a century without generalising across disciplines. A generalisation of Takens’ delay-coordinate embedding theorem to stochastic systems, the Stochastic Embedding Sufficiency Theorem, closes this methodological gap, as an inverse methodology enabling non-parametric recovery of both drift and diffusion fields from scalar time series without prior assumptions about the governing physics.
A blind recovery protocol, receiving only raw time series and sampling interval, is applied identically to nine physical domains: classical mechanics, statistical mechanics, nuclear physics, quantum mechanics, chemical kinetics, electromagnetism, relativistic quantum mechanics, quantum harmonic oscillator dynamics, and quantum electrodynamics. The pipeline recovers the governing equations of each domain blindly, with recovery errors quantified by bootstrap confidence intervals from 50 independent pipeline runs per domain. Fundamental physical constants (the Boltzmann constant, the Planck constant, the speed of light, the Fano factor, and the Van Kampen scaling exponent) emerge in both the drift and diffusion channels without prior specification.
The recovered diffusion coefficients, viewed across domains, constitute an empirical pattern, the $\sigma$-continuum, in which $k_{B}$, $\hbar$, and $c$ each play structurally distinct roles. The Gravitational Diffusion Theorem, derived from the fluctuation–dissipation theorem, the massless mode structure of linearised gravity, and gravitational self-coupling via the equivalence principle, each among the most precisely tested results in physics, determines the gravitational diffusion coefficient as one Planck length per square root of Planck time, with three independent uniqueness arguments converging on the dimensionless prefactor $\alpha=1$.
Four canonical axioms formalise the framework: three grounded in standard physics (Wheeler’s superspace, the derived fluctuation amplitude, and classical correspondence with general relativity) and one interpretive choice (epistemic probability). Physical time emerges as a monotone functional of the stochastic evolution. Within these axioms, the noise character, drift, covariance operator, and fluctuation amplitude are all uniquely determined by theorem (Appendix D), yielding the superspace diffusion hypothesis:
$\mathrm{d}g_{ij}=\mathcal{D}_{ij}[g]\,\mathrm{d}\tau+\ell_{P}\,\mathrm{d}W_{ij},$
where all coefficients are non-parametric, first-principles consequences of the axioms. Coarse-graining of the superspace Fokker–Planck equation via the Mori–Zwanzig projection yields predictions for gravitational acceleration at galactic scales that are testable against kinematic data.
Corresponding author: a.conradie@ucl.ac.uk
Keywords: stochastic differential equations, stochastic embedding sufficiency theorem, non-parametric inference, diffusion coefficient recovery, superspace diffusion, stochastic gravity, galactic kinematics
1 Introduction
1.1 Forward Derivations of Stochastic Differential Equations
The stochastic differential equation
$\mathrm{d}X=\mu(X)\,\mathrm{d}t+\sigma(X)\,\mathrm{d}W$ (1)
appears across physics. The drift $\mu(X)$ encodes deterministic dynamics; the diffusion coefficient $\sigma(X)$ encodes the amplitude of stochastic fluctuations; and $\mathrm{d}W$ is a Wiener increment [1]. Three landmark programmes established this equation in distinct domains.
Langevin [2] derived the equation of motion for a Brownian particle from the fluctuation–dissipation relation, building on Einstein’s kinetic theory of Brownian motion [3]. The diffusion coefficient $\sigma=\sqrt{2\gamma k_{B}T/m}$ was determined by the requirement of thermal equilibrium: the stochastic forcing must balance frictional dissipation at temperature $T$ [4, 5]. The derivation proceeded from known thermodynamics to stochastic dynamics: a forward problem.
Nelson [6] demonstrated that quantum mechanics could be recovered from classical mechanics supplemented by a diffusion process with $\sigma=\sqrt{\hbar/m}$. Starting from a stochastic differential equation with this specific amplitude, Nelson derived the Schrödinger equation [7] as a mathematical consequence. The derivation assumed the value of $\sigma$ from first principles and derived the quantum formalism. Again a forward problem, yielding quantum mechanics rather than presupposing it.
Gillespie [8] derived the Chemical Langevin Equation from discrete molecular reaction dynamics, extending his earlier exact stochastic simulation algorithm [9]. The diffusion coefficient $\sigma\propto\Omega^{-1/2}$ was determined by the system size $\Omega$ through the law of large numbers, consistent with Van Kampen’s system-size expansion [10]. The derivation proceeded from the known reaction network to the mesoscopic stochastic equation: a forward problem requiring complete specification of the underlying chemistry.
Each of these programmes was domain-specific. Each required prior knowledge of the governing physics to determine $\sigma$. None was data-driven: the value of $\sigma$ was derived from theory, not extracted from observation. Nelson could not consider values of $\sigma$ other than $\sqrt{\hbar/m}$. Gillespie could not extract $\sigma$ from a time series without first specifying the reaction network. The forward problem was solved independently in several domains, but could not, by construction, reveal whether the stochastic structure is universal through an inverse methodology.
1.2 Closing the Stochastic Gap
The reconstruction of dynamical systems from time series data has a long history. Takens’ embedding theorem [11] established that the attractor geometry of a deterministic system can be recovered diffeomorphically from scalar observations via delay-coordinate embedding [12]. This result underpins much of modern nonlinear time series analysis.
Takens’ theorem, however, does not accommodate stochasticity. It guarantees reconstruction of the deterministic attractor, but provides no mechanism for recovering the diffusion coefficient $\sigma(X)$ from data. Non-parametric methods for stochastic differential equations have addressed estimation of drift and diffusion from discrete observations, but require direct access to the state variable and typically assume a known parametric class [13, 14]. A generalisation treating stochasticity as dynamical structure to be recovered from scalar observations has been elusive.
The Stochastic Embedding Sufficiency Theorem (Appendix A; complete proof in the Supplementary Materials) closes this gap. Rather than requiring diffeomorphic injectivity on the attractor, the theorem establishes measure-theoretic injectivity on the correlation manifold: sufficient conditions under which both $\mu(X)$ and $\sigma(X)$ can be consistently estimated from scalar time series, without assuming a parametric form for either field and without prior knowledge of whether the system is deterministic or stochastic. Consequently, the drift and diffusion of a stochastic process can be extracted from observed data using a single, domain-agnostic pipeline. The inverse problem (given a time series, determine $\sigma$) becomes tractable.
The following paragraphs summarise the theorem’s architecture, where the complete proof occupies Appendix A and the Supplementary Materials.
The conceptual innovation.
Takens’ theorem reconstructs the deterministic attractor diffeomorphically from delay-coordinate embedding. Noise destroys this: distinct initial conditions can produce identical delay vectors under different realisations, and the diffeomorphism condition fails. The Stochastic Embedding Sufficiency Theorem takes a different approach. Rather than seeking pathwise reconstruction, it establishes that distributional reconstruction suffices: distinct initial conditions produce distinct probability distributions over delay vectors. Recovering the drift $\mu(X)$ and diffusion $\sigma(X)$ requires only that these distributions be distinguishable: a condition weaker than diffeomorphic injectivity but sufficient for consistent non-parametric estimation. The data determines the non-parametric model structure.
The Hörmander–Takens parallel.
In Takens’ theorem, the observation function must be generic to avoid degenerate projections. The stochastic extension requires an analogous non-degeneracy condition on the noise structure. Hörmander’s bracket-generating condition provides this: the Lie algebra generated by the drift and diffusion vector fields must span the tangent space at every point. When satisfied, the stochastic flow explores all directions even if the diffusion matrix is rank-deficient. The parallel is precise: Takens’ genericity prevents the observation from collapsing deterministic geometry, where Hörmander’s condition prevents the noise from confining the stochastic flow to a lower-dimensional submanifold.
Five mathematical foundations.
The proof rests on five mathematical foundations drawn from distinct traditions.
Foundation 1: Hörmander: hypoelliptic regularity. The bracket-generating condition guarantees that transition densities $p_{t}(x,y)$ are smooth and strictly positive for all $t>0$, even when the diffusion matrix is not full rank.
Foundation 2: Malliavin: non-degeneracy of the stochastic flow. The Malliavin covariance matrix is almost surely invertible under Hörmander’s condition, ensuring that the delay embedding map has full rank in a measure-theoretic sense.
Foundation 3: Varadhan–Léandre: transition density separation. The central result: distinct initial conditions $x\neq x^{\prime}$ produce distinct transition densities for all $\tau>0$. The proof uses semigroup bisection and the short-time heat kernel asymptotic $\lim_{t\to 0}t\log p_{t}(x,z)=-d(x,z)^{2}/2$, where $d$ is the sub-Riemannian distance.
Foundation 4: Frostman: measure geometry of the collision set. The parametric transversality theorem establishes that the collision set has controlled codimension. The Frostman covering argument converts this codimension into measure zero under the invariant measure $\mu_{\infty}$, bridging smooth geometry and fractal geometry.
Foundation 5: Stone: non-parametric estimation consistency. The consistency of $k$-nearest-neighbour estimators on the correlation manifold, with convergence rate $O_{P}((k/N)^{\beta/m^{*}})+O(\Delta t)$, provides the bridge from the mathematical sufficiency result to the computational pipeline.
The synthesis.
The five foundations converge in the Probabilistic Uplift Theorems (Supplementary Materials, Theorems 4.5–4.6). Hörmander guarantees the target functions are smooth, Malliavin guarantees the embedding is non-degenerate, Varadhan–Léandre guarantees the estimation targets are single-valued, Frostman guarantees single-valuedness under the invariant measure, and Stone provides the convergence rate. The synthesis produces the operational conclusion: given a scalar time series of sufficient length from an SDE satisfying Hörmander’s condition, the pipeline recovers both $\mu(X)$ and $\sigma(X)$ with convergence guaranteed.
The nine-domain validation (Section 3) serves a dual purpose: it demonstrates that the pipeline recovers known physics and confirms the theorem’s predictions empirically: that consistent recovery of both drift and diffusion is achievable from scalar observations across dynamical regimes ranging from deterministic through stochastic.
1.3 Scope and Structure
This paper applies the inverse methodology, the data-driven recovery of governing physics, to nine physical domains spanning classical mechanics, statistical mechanics, nuclear physics, quantum mechanics, chemical kinetics, electromagnetism, relativistic quantum mechanics, quantum field theory, and quantum electrodynamics. All data is synthetically generated from established governing equations, providing exact theoretical values against which the blind recovery is validated.
The recovered diffusion coefficients, viewed across domains, reveal a systematic pattern: the $\sigma$-continuum. The physical constants of each domain ($k_{B}$, $\hbar$, $c$, the Van Kampen exponent, the Fano factor) emerge in the recovered fields without prior specification. The pattern extends from $\sigma=0$ (deterministic domains) through thermal, Poisson, quantum, and chemical regimes to relativistic and quantum field theory scales.
Derivation from the relativistic and quantum field theory results yields a specific superspace diffusion hypothesis, formalised by four canonical axioms within which the gravitational diffusion coefficient, the drift, the covariance operator, and physical time itself are all fixed. The resulting framework contains non-parametric, first-principles predictions within the stated axioms. At galactic scales, the Fokker–Planck equation on superspace generates predictions for gravitational acceleration that are testable against kinematic data including rotation curves, velocity dispersions, wide binaries, galaxy clusters, gravitational lensing, and cosmic voids.
The paper is structured as follows. Section 2 describes the recovery pipeline, the blind protocol, the data generation procedures, and the statistical validation framework. Section 3 presents the nine-domain results with statistical rigour and introduces the $\sigma$-continuum. Section 4 derives the gravitational diffusion coefficient from the results in three stages: relativistic structure (§4.2, summarising Appendix B), the gravitational diffusion coefficient (§4.3, summarising Appendix C), and axiomatic formalisation with uniqueness proofs (§4.4, summarising Appendix D). Section 5 states the conclusions. Appendices A–D provide the mathematical derivations: the Stochastic Embedding Sufficiency Theorem (A), the relativistic transformation properties (B), the gravitational diffusion coefficient (C), and the uniqueness theorems (D).
2 Materials and Methods
2.1 The Stochastic Differential Equation
The central object of this study is the stochastic differential equation stated in Eq. (1). The drift $\mu$ and diffusion $\sigma$ are both unknown a priori and are to be recovered from time series data using the methodology described below.
2.2 The Stochastic Embedding Pipeline
A non-parametric reconstruction framework recovers $\mu(X)$ and $\sigma(X)$ directly from a scalar time series $\{y(t_{1}),y(t_{2}),\ldots,y(t_{N})\}$ without assuming a parametric model for either field. The framework extends Takens’ delay-coordinate embedding [11] to accommodate stochastic dynamics, drawing on the Stochastic Embedding Sufficiency Theorem (Appendix A). Whereas Takens’ theorem guarantees diffeomorphic reconstruction of deterministic attractors, the Stochastic Embedding Sufficiency Theorem establishes measure-theoretic injectivity on the correlation manifold, sufficient for consistent estimation of both drift and diffusion. The complete proof is given in the Supplementary Materials.
The reconstruction proceeds in three stages.
Stage 1: Embedding
The scalar time series $y(t)$ is embedded into a $d$-dimensional state space via the delay map
$\mathbf{Y}(t)=\bigl[y(t),\;y(t-\tau),\;y(t-2\tau),\;\ldots,\;y(t-(d-1)\tau)\bigr],$ (2)
where $d$ is the embedding dimension and $\tau$ is the delay. The embedding dimension is selected using the $E_{1}$ statistic of Cao [15]: $E_{1}(d)$ measures the ratio of nearest-neighbour distances in successive embedding dimensions, and saturation of $E_{1}$ indicates that the correlation manifold has been unfolded. For non-oscillatory systems (C1–C7), $\tau=1$ (one sampling interval) is used. For oscillatory systems (C8, C9), where the sampling interval is much shorter than the dynamical timescale, $\tau$ is set to the first zero-crossing of the autocorrelation function, ensuring that the delay embedding spans a meaningful portion of the phase space (Table 1). The embedding dimension $d$ is domain-specific and determined by $E_{1}$ saturation.
Cao’s $E_{2}$ statistic is computed alongside $E_{1}$ for characterisation of the dynamical regime. Departure of $E_{2}$ from unity indicates stochastic content in the time series; $E_{2}\approx 1$ indicates deterministic dynamics. The $E_{2}$ statistic is used for classification only and does not influence the embedding dimension.
Following delay embedding, singular value decomposition (SVD) [16] projects the embedded vectors onto a low-rank subspace retaining at least 95% of the total variance. This projection reduces the effective dimensionality of the correlation manifold while preserving the geometric structure relevant to nearest-neighbour queries.
Stage 2: Local Geometry on the Correlation Manifold
At each point $\mathbf{Y}(t_{i})$ on the correlation manifold, the $k$ nearest neighbours are identified using a KD-tree index [17] with $O(\log N)$ query complexity. For each query point, the forward increments $\Delta y_{j}=y(t_{j}+\Delta t)-y(t_{j})$ of its $k$ neighbours are collected, and the local drift and diffusion are estimated via the Kramers–Moyal conditional moments:
$\hat{\mu}(\mathbf{Y}_{i})=\frac{1}{k\,\Delta t}\sum_{j\in\mathcal{N}_{i}}\Delta y_{j},\qquad\hat{\sigma}^{2}(\mathbf{Y}_{i})=\frac{1}{k\,\Delta t}\sum_{j\in\mathcal{N}_{i}}\bigl(\Delta y_{j}\bigr)^{2}\;-\;\hat{\mu}(\mathbf{Y}_{i})^{2}\,\Delta t,$ (3)
where $\mathcal{N}_{i}$ denotes the set of $k$ nearest neighbours of the $i$-th embedded vector and $\Delta t$ is the sampling interval. The subtraction of $\hat{\mu}^{2}\Delta t$ in the diffusion estimator removes the $O(\Delta t)$ bias from the drift contribution (see Supplementary Materials, Remark 3.12, for the full derivation from Itô–Taylor expansion). No parametric form is assumed for $\mu$ or $\sigma$; both fields emerge from the local statistics of the data.
Stage 3: Validation
The recovered drift and diffusion fields are validated through autonomous free-run forecasting. Starting from a short seed sequence (typically 10 time steps drawn from the true series), the Euler–Maruyama scheme [18, 19]
$y(t_{n+1})=y(t_{n})+\hat{\mu}\bigl(\mathbf{Y}(t_{n})\bigr)\,\Delta t+\hat{\sigma}\bigl(\mathbf{Y}(t_{n})\bigr)\,\sqrt{\Delta t}\;\xi_{n},\qquad\xi_{n}\sim\mathcal{N}(0,1),$ (4)
generates stochastic trajectories using only the pipeline-recovered fields. For stochastic systems, an ensemble of 250 independent realisations is generated, and the fraction of time steps for which the true trajectory lies within the 95% confidence interval of the ensemble is computed. This 95%-CI coverage serves as the primary validation metric, where nominal coverage is 95%.
2.3 The Blind Recovery Protocol
A strict methodological separation prevents contamination between the data generation and blind recovery analysis. The pipeline receives only the raw time series $\{y(t_{i})\}$ and the sampling interval $\Delta t$. No physical parameters, governing equations, boundary conditions, or domain labels are provided. The embedding dimension, the dynamical regime classification (deterministic or stochastic), and the functional forms of $\mu(X)$ and $\sigma(X)$ are determined entirely by the pipeline.
This strict separation ensures that the recovery of physical constants from $\hat{\mu}$ and $\hat{\sigma}$ constitutes a data-driven inverse derivation: the physics emerges from the data, not from prior knowledge supplied to the algorithm.
2.4 Data Generation
All time series analysed in this study are synthetically generated from known physics. This is stated explicitly: no experimental or observational data is used. Synthetic generation from established governing equations provides exact first-principles values of $\mu$ and $\sigma$ against which the blind recovery is validated.
Nine physical domains are tested, spanning classical mechanics, statistical mechanics, nuclear physics, quantum mechanics, chemical kinetics, electromagnetism, relativistic quantum mechanics, and quantum field theory. Table 1 summarises the governing equation, simulation method, series length, and theoretical diffusion coefficient for each domain.
Table 1: Data generation summary for the nine domain validations. NN denotes the number of time series samples provided to the pipeline; $k$ denotes the number of nearest neighbours used in Stage 2; $\tau_{\mathrm{embed}}$ is the embedding lag in samples. For oscillatory systems (C8, C9), $\tau_{\mathrm{embed}}$ is set to the first zero-crossing of the autocorrelation function rather than unity, ensuring that the delay-coordinate embedding spans a meaningful portion of the dynamical phase space. All data is synthetically generated from known physics.
Test Domain Simulation Theoretical $\sigma$ NN $k$ $\tau_{\mathrm{embed}}$
C1 Classical mechanics Numerical integration of Kepler orbit with GR precession 0 (deterministic) 36 525 20 1
C2 Statistical mechanics Euler–Maruyama Langevin equation $\sqrt{2\gamma k_{B}T/m}$ 100 000 200 1
C3 Nuclear physics Poisson sampling $\sqrt{\mu}$ (Fano =1=1) 10 000 200 1
C4 Quantum mechanics Spectral Schrödinger evolution (wavepacket width) $\sqrt{\hbar/m}$ 2 001 30 1
C5 Chemical kinetics Tau-leaping birth–death process $\propto\Omega^{-1/2}$ 1 001 30 1
C6 Electromagnetism Gaussian wavepacket propagation (centroid tracking) 0 (deterministic) 500 30 1
C7 Relativistic QM Euler–Maruyama Nelson SDE with relativistic dispersion $\sigma_{0}/\gamma$ 2 000 30 1
C8 Quantum field theory Euler–Maruyama damped QHO at zero-point $\sqrt{\gamma\hbar\omega/m}$ 20 000 100 $\tau_{\mathrm{ACF}}$
C9 QED (photon field) Euler–Maruyama QHO with $m_{\mathrm{eff}}=\hbar\omega/c^{2}$ $c\sqrt{\gamma}$ 20 000 100 $\tau_{\mathrm{ACF}}$
The nine domains are selected to span a progression of increasing complexity: from deterministic systems ($\sigma=0$; C1, C4, C6) through constant additive noise (C2) and state-dependent diffusion (C3, C8) to systems with both state-dependent drift and state-dependent diffusion (C5), relativistic corrections (C7), and the quantum field theory regime where fundamental constants cancel from the diffusion coefficient (C9).
2.5 Statistical Validation Framework
For each domain, the following statistical measures are reported.
The recovery error quantifies agreement between the pipeline-recovered quantity and its theoretical value as a percentage: $\varepsilon=|\hat{\theta}-\theta_{\mathrm{true}}|/|\theta_{\mathrm{true}}|\times 100\%$.
Bootstrap confidence intervals (95%) are the primary statistical assessment. For each domain, 50 independent pipeline runs are performed, each with independently generated synthetic data. Each run yields a normalised ratio $\hat{\theta}/\theta_{\mathrm{true}}$. The pairwise bootstrap (10 000 resamples of the 50 ratios with replacement) gives a distribution of the sample median; the 2.5th and 97.5th percentiles of this distribution define the 95% confidence interval [20]. If the interval contains 1.0, the recovered quantity is consistent with theory at the stated confidence level.
For each domain, both the drift $\hat{\mu}$ and diffusion $\hat{\sigma}$ are extracted and reported where the physics admits a meaningful comparison (dual-channel reporting). For deterministic systems (C1, C4, C6), the drift channel recovers the governing dynamical parameter while the diffusion channel confirms $\hat{\sigma}/\mathrm{signal}\ll 1$. For stochastic systems, the diffusion channel recovers $\sigma$ or a derived physical constant, and the drift channel recovers the deterministic component of the dynamics.
For domains testing scaling exponents (C5: Van Kampen exponent; C7: relativistic suppression slope), the pipeline is run at multiple parameter values within each simulation, and the log–log slope is computed per simulation before bootstrapping. The multiplicative $k$-NN estimator bias, which is approximately uniform across the parameter range, cancels in the slope [14].
The 95%-CI coverage of the free-run ensemble (Stage 3) is computed for stochastic systems. The fraction of time steps for which the true trajectory lies within the 2.5th–97.5th percentile band of the 250-realisation ensemble is reported, where values near 95% indicate well-calibrated uncertainty quantification.
3 Results
The blind recovery protocol of §2.3 was applied identically to all nine domains. Results are presented in order of increasing complexity, beginning with deterministic systems and progressing through additive and state-dependent stochastic dynamics to relativistic and quantum field theory regimes. All data is synthetically generated from known physics as described in §2.4.
3.1 Deterministic Domains: $\sigma=0$
C1: Classical Mechanics, Mercury’s Perihelion Precession
The pipeline was applied to two representations of Mercury’s orbital dynamics: a cumulative precession time series (linear trend at 574.61 arcsec/century, as predicted by general relativity [21]) and a full Keplerian orbit with secular precession.
For the cumulative precession series (Fig. 2a,b), SVD identified rank 1, and the recovered drift from a single primary run reproduced the theoretical precession rate to 0.02%. The median recovered diffusion was $\hat{\sigma}=7.19\times 10^{-14}$ arcsec, indistinguishable from zero at machine precision. Across 50 independent pipeline runs with small measurement noise, the drift-channel ratio $\hat{\theta}/\theta_{\mathrm{true}}$ had median $1.0000$ (within 1 ppm of unity) with bootstrap 95% CI [1.0000, 1.0000][1.0000,\,1.0000]; the diffusion channel confirmed $\hat{\sigma}/\mathrm{signal}=1.2\times 10^{-6}$.
For the Keplerian orbit (Fig. 2c,d), SVD identified rank 2 (consistent with a two-dimensional attractor), and autonomous free-run forecasting from a 10-point seed tracked the true orbit for 2.35 years (approximately 10 orbital periods) with RMSE $1.63\times 10^{-4}$ AU. The recovered $\hat{\sigma}=3.40\times 10^{-5}$ AU reflects $k$-nearest-neighbour discretisation on a curved attractor, not physical stochasticity: the free-run forecast remains deterministically accurate over the full prediction horizon.
Classical mechanics operates in the deterministic limit of Eq. (1): $\sigma=0$.
C4: Quantum Mechanics, Wavepacket Spreading
The width $\sigma(t)$ of a free Gaussian wavepacket evolving under the Schrödinger equation [7] was tracked as a scalar time series. The pipeline recovered a nonlinear drift $\hat{\mu}(\sigma)$ consistent with the analytical spreading law $\mathrm{d}\sigma/\mathrm{d}t=(\hbar/(2m\sigma_{0}))^{2}\cdot t/\sigma(t)$, and a diffusion field $\hat{\sigma}\approx 0$ (median $3.90\times 10^{-5}$), confirming that wavepacket spreading is a deterministic evolution.
The Planck constant was extracted from the recovered drift: $\hat{\hbar}=1.00026$ (error $0.026\%$ relative to the simulation value $\hbar=1$). Across six independent simulations with initial widths $\sigma_{0}\in[1,5]$, the recovered $\hat{\hbar}$ was stable to within $0.05\%$ (Fig. 3d). Fifty independent pipeline runs with added measurement noise yielded median $\hat{\hbar}/\hbar=1.0005$ with bootstrap 95% CI [1.0001, 1.0015][1.0001,\,1.0015] (Fig. 3e); the diffusion channel confirmed $\hat{\sigma}/\mathrm{signal}=9.5\times 10^{-3}$ (deterministic).
The wavepacket system has $\sigma=0$ (deterministic dynamics) but encodes $\hbar$ in the drift. This is the first indication that fundamental constants appear in the recovered fields.
C6: Electromagnetism, Wave Propagation
The electromagnetic wave propagation results are presented jointly with C7 (Klein–Gordon suppression) in §3.5 below, as the transition from deterministic wave propagation ($\sigma=0$) to relativistic diffusion suppression constitutes the pivotal connection between classical and quantum regimes. In summary: the pipeline recovered $\hat{\sigma}/\mathrm{signal}=4.3\times 10^{-7}$ and $\hat{c}/c_{\mathrm{true}}=1.0000$ with bootstrap 95% CI [1.0000, 1.0000][1.0000,\,1.0000] from 50 independent runs (Fig. 8a,b).
3.2 Additive Noise: Constant $\sigma$
C2: Statistical Mechanics, Brownian Motion
The velocity $v(t)$ of a Brownian particle governed by the Langevin equation $\mathrm{d}v=-\gamma v\,\mathrm{d}t+\sigma\,\mathrm{d}W$ [2, 4], with $\sigma=\sqrt{2\gamma k_{B}T/m}$, was simulated by Euler–Maruyama integration ($10^{6}$ steps, subsampled to $10^{5}$).
The pipeline recovered a linear drift $\hat{\mu}(v)=-\hat{\gamma}v$ (Fig. 4c) and a flat diffusion profile $\hat{\sigma}(v)\approx\mathrm{const.}$ (Fig. 4b), consistent with additive noise. The Boltzmann constant, extracted from the fluctuation–dissipation relation $k_{B}=m\hat{\sigma}^{2}/(2\hat{\gamma}T)$ [5, 22], was recovered as $\hat{k}_{B}T/(k_{B}T)_{\mathrm{true}}=0.984$ with bootstrap 95% CI [0.979, 0.989][0.979,\,0.989] from 50 independent simulations (Fig. 4f), a $1.56\%$ error. The drift channel yielded $\hat{\gamma}/\gamma_{\mathrm{true}}=0.971$ with bootstrap 95% CI [0.902, 0.998][0.902,\,0.998] (Fig. 4e). The wider drift-channel CI reflects the lower signal-to-noise ratio of local drift estimation relative to diffusion recovery at these parameters.
The 250-realisation free-run ensemble achieved 94.5% coverage of the true trajectory (nominal 95%), confirming well-calibrated uncertainty quantification.
Brownian motion encodes $k_{B}$ in $\sigma$ and $\gamma$ in $\mu$. Both fundamental constants are recovered from their respective channels.
3.3 State-Dependent Diffusion
C3: Nuclear Physics, Radioactive Decay
Radioactive count data ($\mu=3000$ counts per interval, N=10  [^0] N=10\,000 measurements) were generated by Poisson sampling. The theoretical prediction is $\sigma^{2}=\mu$ (Fano factor $F=1$ [23]): the diffusion coefficient depends on the state.
The pipeline recovered a state-dependent diffusion profile $\hat{\sigma}^{2}(\mu)$ lying on the Poisson line $\sigma^{2}=\mu$ across a two-decade range of mean count levels (Fig. 5b). At a single count rate ($\mu=3000$), the pipeline-recovered diffusion was $\hat{\sigma}/\sqrt{\mu}=1.059$ with bootstrap 95% CI [1.051, 1.062][1.051,\,1.062] from 50 independent pipeline runs, a $5.9\%$ positive bias consistent with the $O(k/N)$ bias of $k$-NN Kramers–Moyal estimators at $k=200$, $N=5000$. Across the variance–mean sweep, this multiplicative bias is approximately uniform and cancels in the log–log slope: the variance–mean exponent was $0.999$ (theoretical $1.000$; error $0.09\%$; $R^{2}=1.0000$). The 250-realisation ensemble achieved 95.8% CI coverage.
The transition from constant $\sigma$ (C2) to state-dependent $\sigma(x)=\sqrt{x}$ (C3) demonstrates that the pipeline distinguishes additive from multiplicative noise without prior specification.
C8: Quantum Harmonic Oscillator
The velocity of a damped quantum harmonic oscillator at zero temperature was simulated by Euler–Maruyama integration. At zero temperature, the fluctuation–dissipation theorem [5, 22] yields $\sigma^{2}=\gamma\hbar\omega/m$, where $\gamma$ is the damping rate, $\omega$ is the oscillator frequency, and $m$ is the mass. The zero-point energy $\tfrac{1}{2}\hbar\omega$ replaces the thermal energy $k_{B}T$ of C2.
The pipeline recovered $\hat{\sigma}/\sigma_{\mathrm{true}}=0.997$ with bootstrap 95% CI [0.995, 1.000][0.995,\,1.000] from 50 independent runs (Fig. 6b), a $0.27\%$ error. For oscillatory systems, the embedding lag $\tau_{\mathrm{embed}}$ is set to the autocorrelation function zero-crossing ($\tau_{\mathrm{ACF}}\approx 155$$160$ steps) rather than unity, consistent with standard practice in nonlinear time series analysis for oversampled oscillatory data, ensuring that the delay-coordinate embedding spans a meaningful portion of the oscillator’s phase space. Across nine frequency values spanning a decade, the recovered $\hat{\sigma}^{2}$ was proportional to $\omega$ with $R^{2}=0.9999$ (Fig. 6c), confirming the linear frequency dependence predicted by the fluctuation–dissipation relation.
C8 establishes that $\hbar$ enters $\sigma$ through quantisation of the energy levels: the zero-point energy is irreducibly quantum, and $\hbar$ sets its scale.
3.4 State-Dependent Drift and Diffusion
C5: Chemical Kinetics, Van Kampen Scaling
A birth–death process with system size $\Omega$ was simulated by tau-leaping. Van Kampen’s system-size expansion is a mathematical theorem [10]: $\sigma\propto\Omega^{-1/2}$. This system has both state-dependent drift (mean-reverting toward equilibrium) and state-dependent diffusion, making it the most demanding test of the pipeline across C1–C5.
The pipeline was applied independently at seven system sizes ($\Omega\in[50,5000]$), with 50 independent simulations per system size. On a log–log plot of recovered $\hat{\sigma}$ versus $\Omega$ (Fig. 7d), each simulation yielded a per-run scaling exponent; the median across 50 runs was $\hat{\alpha}=-0.502$ with bootstrap 95% CI [0.503,0.500][-0.503,\,-0.500] (theoretical $\alpha=-0.500$; error $0.38\%$; $R^{2}=1.0000$). The bootstrap CI contains the theoretical value 0.500-0.500, demonstrating that the multiplicative $k$-NN estimator bias, which is approximately $3\%$ at each individual $\Omega$ value, cancels in the log–log slope (Fig. 7e). The 250-realisation ensemble at $\Omega=500$ achieved 94.0% CI coverage.
A mathematical theorem has been recovered from data without knowledge of the underlying chemistry.
3.5 Relativistic Corrections
C6/C7: The Relativistic Hinge, From Electromagnetism to Klein–Gordon Suppression
Figure 8 presents the C6 (electromagnetic wave) and C7 (Klein–Gordon) results together, as the transition from deterministic wave propagation to relativistic diffusion suppression constitutes the pivotal observation connecting classical and quantum regimes.
For C6, a classical electromagnetic wavepacket was propagated, and the centroid position tracked. The pipeline recovered $\hat{\sigma}/\mathrm{signal}=4.3\times 10^{-7}$ (indistinguishable from zero) and a propagation speed $\hat{c}/c_{\mathrm{true}}=1.0000$ with bootstrap 95% CI [1.0000, 1.0000][1.0000,\,1.0000] from 50 independent runs (Fig. 8a,b).
For C7, Nelson’s stochastic differential equation [6, 24] for a massive relativistic particle was simulated at seven momenta $k_{0}\in[0.3,5.0]$, corresponding to Lorentz factors $\gamma\in[1.04,5.10]$. The theoretical prediction is $\sigma_{\mathrm{KG}}=\sigma_{0}/\gamma$: the diffusion coefficient is suppressed by the Lorentz factor.
On a log–log plot of recovered $\hat{\sigma}/\hat{\sigma}_{0}$ versus $\gamma$ (Fig. 8d), the median slope across 50 independent simulations was 0.993-0.993 with bootstrap 95% CI [0.999,0.988][-0.999,-0.988] (theoretical 1.000-1.000; error $0.72\%$; $R^{2}=0.9999$). The per-momentum $\hat{\sigma}/\sigma_{\mathrm{true}}$ ratio was approximately constant at 0.960.960.970.97 across all Lorentz factors, confirming that the $k$-NN estimator bias is multiplicative and cancels in the slope.
The drift simultaneously recovered the relativistic group velocity $v_{g}=c\,k_{0}/\sqrt{1+k_{0}^{2}}$ at each momentum (Fig. 8f), confirming that both channels (drift and diffusion) encode relativistic structure.
C7 reveals that the speed of light $c$ enters $\sigma$ through the relativistic dispersion relation. Combined with C4 (where $\hbar$ enters through the drift), the pattern indicates that fundamental constants emerge naturally in the recovered fields: $\hbar$ governs the quantum scale of fluctuations, while $c$ governs their relativistic transformation.
The relativistic suppression $\sigma_{\mathrm{KG}}=\sigma_{0}/\gamma$ arises as the product of two structurally independent factors (Appendix B).
The first factor is kinematic: time dilation rescales the Wiener increment. The Dambis–Dubins–Schwarz reparametrisation theorem: a result of pure stochastic calculus, independent of any physical assumption: establishes that reparametrising the time of a Wiener process rescales its increments by the square root of the time Jacobian. Applied to the proper-time/coordinate-time relation $\mathrm{d}\tau=\mathrm{d}t/\gamma$, this gives $\mathrm{d}W(\tau)=\mathrm{d}B(t)/\sqrt{\gamma}$ and hence a kinematic suppression $\sigma_{\mathrm{rel}}=\sigma_{0}/\sqrt{\gamma}$.
The second factor is dynamical: the Klein–Gordon conserved density $\rho_{\mathrm{KG}}=\gamma|\phi|^{2}$ requires a self-consistent Fokker–Planck equation with diffusion coefficient $D_{\mathrm{KG}}=\sigma_{0}^{2}/(2\gamma^{2})$, producing an additional $1/\sqrt{\gamma}$ suppression.
The two factors operate at different levels (noise amplitude versus drift), arise from different inputs (special relativity versus Klein–Gordon current conservation), and use different mathematical tools (DDS theorem versus Fokker–Planck self-consistency). Their product $\sigma_{\mathrm{KG}}=\sigma_{0}/\gamma$ is confirmed by the C7 recovery: slope 0.993-0.993 with bootstrap 95% CI [0.999,0.988][-0.999,-0.988], consistent with the predicted exponent 1-1 at the $0.72\%$ level. The independent recovery of the group velocity $v_{g}$ at each momentum mitigates the potential circularity that $\gamma$ depends on $v$: both channels: drift and diffusion: carry relativistic content, recovered independently from the same blind pipeline.
3.6 Quantum Field Theory: The Roles of $\hbar$, $c$, and $G$
C9: QED Photon Field, The Massless Cancellation
A single mode of the quantised electromagnetic field was simulated [25] as a damped quantum harmonic oscillator with effective mass $m_{\mathrm{eff}}=\hbar\omega/c^{2}$ (from mass–energy equivalence). Substitution into the fluctuation–dissipation relation of C8 gives
$\sigma^{2}=\frac{\gamma\hbar\omega}{m_{\mathrm{eff}}}=\frac{\gamma\hbar\omega}{\hbar\omega/c^{2}}=\gamma c^{2},$ (5)
so that $\sigma=c\sqrt{\gamma}$. Both $\hbar$ and $\omega$ cancel identically.
The pipeline was applied at seven frequencies spanning a decade ($\omega\in[10^{12},10^{13}]$ rad/s) at fixed damping $\gamma=10^{11}$ s-1. As for C8, the embedding lag was set to the autocorrelation zero-crossing ($\tau_{\mathrm{ACF}}\approx 157$ steps). The recovered $\hat{\sigma}$ was constant across all frequencies to within 1% (Fig. 9c). From 50 independent pipeline runs at the reference frequency, the recovered $\hat{\sigma}/\sigma_{\mathrm{true}}=0.998$ with bootstrap 95% CI [0.996, 0.999][0.996,\,0.999], a $0.24\%$ error (Fig. 9d).
This result clarifies the distinct roles of the three constants that will constitute $\sigma$ for gravity. The Planck constant $\hbar$ enters through quantisation ($E=\hbar\omega$, $m_{\mathrm{eff}}=\hbar\omega/c^{2}$) and then cancels from $\sigma$ for massless fields; its role is to make the field quantum, not to set the fluctuation amplitude. The speed of light $c$ appears directly in $\sigma=c\sqrt{\gamma}$; its role is to set the velocity scale of fluctuations. The damping rate $\gamma$ determines the coupling to the environment.
For the photon, $\gamma$ is an external parameter (cavity losses, material absorption). For gravity, however, the coupling is not external: the gravitational field is self-interacting (metric fluctuation modes carry energy, and that energy gravitates by the equivalence principle), so the self-coupling rate is uniquely determined by the remaining constant, Newton’s gravitational constant $G$, through $\gamma_{\mathrm{grav}}=1/t_{P}$, where $t_{P}=\sqrt{\hbar G/c^{5}}$ is the Planck time [26]. Substitution yields $\sigma=c\sqrt{1/t_{P}}=\ell_{P}/\sqrt{t_{P}}$, which corresponds to a proper-distance fluctuation of one Planck length per Planck time. This derivation is developed in the Discussion (§4) and proved in Appendix C (the Gravitational Diffusion Theorem).
3.7 Summary of Quantitative Results
Table 2 consolidates the principal recovered quantities, bootstrap confidence intervals, and errors for all nine domains. For each domain, both the drift and diffusion channels are reported where the physics admits a meaningful comparison.
Table 2: Summary of quantitative results from 50 independent pipeline runs per domain. $\hat{\theta}/\theta$: median normalised recovery. Bootstrap 95% CIs (10 000 pairwise resamples) serve as the primary statistical assessment. For deterministic systems, $\hat{\sigma}/\mathrm{signal}\ll 1$ confirms correct classification. For scaling exponents (C5), the multiplicative $k$-NN bias cancels in the log–log slope and the CI contains the theoretical value. Ensemble CI coverage is reported for stochastic free-run ensembles (250 realisations). Tests are ordered by the progression of complexity described in §2.4.
Test Domain Channel $\hat{\theta}/\theta$ 95% CI Error Cov.
C1 Classical mech. Drift (rate) 1.0000 $[1.0000,1.0000]$ 0.0001% (det.)
Diff. ($\hat{\sigma}/\mathrm{sig.}$) $1.2\!\times\!10^{-6}$
C4 Quantum mech. Drift ($\hat{\hbar}$$\hbar$) 1.0005 $[1.0001,1.0015]$ 0.05% (det.)
Diff. ($\hat{\sigma}/\mathrm{sig.}$) $9.5\!\times\!10^{-3}$
C6 Electromagnetism Drift ($\hat{c}/c$) 1.0000 $[1.0000,1.0000]$ $3\!\times\!10^{-6}$% (det.)
Diff. ($\hat{\sigma}/\mathrm{sig.}$) $4.3\!\times\!10^{-7}$
C2 Statistical mech. Drift ($\hat{\gamma}/\gamma$) 0.971 $[0.902,0.998]$ 2.95%
Diff. ($\hat{k}_{B}T/k_{B}T$) 0.984 $[0.979,0.989]$ 1.56% 94.5%
C3 Nuclear physics Diff. ($\hat{\sigma}/\sqrt{\mu}$) 1.059 $[1.051,1.062]$ 5.9% 95.8%
C8 Quantum field th. Diff. ($\hat{\sigma}/\sigma$) 0.997 $[0.995,1.000]$ 0.27%
C5 Chemical kinetics Diff. ($\hat{\alpha}/(-0.5)$) 1.004 $[1.000,1.007]$ 0.38% 94.0%
C7 Relativistic QM Drift ($\hat{v}_{g}/v_{g}$) 0.931 $[0.922,0.945]$ 6.9%
Diff. (slope/(1)/(-1)) 0.993 $[0.988,0.999]$ 0.72%
C9 QED (photon) Diff. ($\hat{\sigma}/\sigma$) 0.998 $[0.996,0.999]$ 0.24%
[…]
3.8 Emergence of Physical Constants in Drift and Diffusion
Across the nine domains, fundamental physical constants are recovered from both the drift $\hat{\mu}$ and the diffusion $\hat{\sigma}$. Table 3 summarises which constants emerge from which channel.
Table 3: Physical constants recovered from the drift and diffusion channels. The inverse methodology extracts the complete dynamical law, where constants emerge in both channels without prior specification.
Test Constants in $\hat{\mu}$ (drift) Constants in $\hat{\sigma}$ (diffusion)
C1 GR precession rate ($\sigma=0$)
C2 Friction $\gamma$ Boltzmann constant $k_{B}$
C3 (constant drift) Fano factor ($F=1$)
C4 Planck constant $\hbar$ ($\sigma\approx 0$)
C5 Mean-reversion rate Van Kampen exponent ($\tfrac{1}{2}$)
C6 Speed of light $c$ ($\sigma=0$)
C7 Group velocity $v_{g}(k_{0},c)$ Relativistic suppression $\sigma_{0}/\gamma$
C8 Damping rate $\gamma$ $\hbar$, $\omega$ (via zero-point energy)
C9 Damping rate $\gamma$ $c$ (with $\hbar$ and $\omega$ cancelled)
The progression from C8 to C9 is of particular significance. In C8, $\hbar$ enters $\sigma$ explicitly through the zero-point energy. In C9, both $\hbar$ and the oscillator frequency $\omega$ cancel identically from $\sigma$ for a massless field, leaving $\sigma=c\sqrt{\gamma}$. The speed of light $c$, absent from $\sigma$ in C2–C5, enters as the propagation speed of the massless field. The damping rate $\gamma$, an external parameter in C9, becomes, for gravity, the self-coupling rate $\gamma_{\mathrm{grav}}=1/t_{P}$, which is the unique rate constructible from $\{\hbar,G,c\}$. The three constants thus play distinct and irreducible roles: $\hbar$ quantises the field, $c$ sets the velocity scale, and $G$ determines the self-coupling that closes the system without parameters fitted to the predicted data.
3.9 The $\sigma$-Continuum
Figure 1 presents the diffusion coefficient $\sigma$ recovered across all nine domains on a single axis. The pattern is an empirical observation: $\sigma$ varies systematically from zero (deterministic domains: C1, C6) through thermal ($k_{B}T$-dependent: C2), Poisson (state-dependent: C3), quantum ($\hbar$-dependent: C4, C8), and chemical ($\Omega$-dependent: C5) regimes, to relativistic ($c$-dependent: C7, C9).
This continuum is an empirical observation within the present nine-domain study, emerging from nine independent applications of the same blind recovery protocol. The question of what it implies: whether a common stochastic structure extends across these domains and what form it takes for gravitational degrees of freedom is addressed in the Discussion.
4 Discussion
The Discussion proceeds through four stages. Section 4.1 identifies the $\sigma$-continuum as an empirical observation arising from the nine-domain results. Sections 4.24.3 derive the gravitational diffusion coefficient from this observation combined with three foundational results of established physics; the complete proof chain occupies Appendix C and constitutes a theorem (the Gravitational Diffusion Theorem, Theorem C.6), not an interpretive hypothesis. Section 4.4 poses four canonical axioms motivated by the preceding derivation. Appendix D demonstrates that within these axioms, every element of the superspace SDE is uniquely determined. Table 4 summarises the epistemic status of each step.
Table 4: Derivation chain: epistemic status of each step from empirical observation through theorems to axiomatic formalisation. The column “External inputs” identifies what each step assumes beyond the preceding steps. Only four canonical axioms are postulated; all other results are derived.
Step What is established Epistemic status External inputs Location
$\sigma$-continuum Diffusion coefficients across 9 domains follow systematic pattern Empirical observation None §3
FDT master formula $\sigma^{2}=\gamma\hbar\omega/m$ for damped QHO at zero temperature Theorem
Callen–Welton App. C, Lem. C.2
Massless cancellation $\sigma=c\sqrt{\gamma}$ for massless fields; $\hbar$ and $\omega$ cancel identically Theorem
(this work)
Mass–energy equiv. App. C, Thm. C.4
Dimensional uniqueness of $\gamma_{\mathrm{grav}}$ $\gamma_{\mathrm{grav}}=\alpha/t_{P}$ is the unique rate from $\{\hbar,G,c\}$ Theorem
(this work)
None App. C, Lem. C.5
Self-bath (Mori–Zwanzig) Self-coupling follows from NZ projection of superspace FP equation Theorem
(this work)
Equivalence principle App. C, Prop. C.1
$\alpha\leq 1$ (metric consistency) Perturbation theory: $\langle h^{2}\rangle=\alpha^{2}\leq 1$ over one $t_{P}$ Theorem
(this work)
Metric positivity App. C, Thm. C.22
$\alpha=1$ (critical damping) $\gamma(\omega^{*})=\omega^{*}$ has unique solution $\omega^{*}=1/t_{P}$ Theorem
(this work)
Self-coupling rate App. C, Thm. C.23
$\alpha\geq 1$ (singularity resolution) Curvature fluctuations must reach $K_{P}$ at the Planck scale Theorem
(this work)
Curvature scaling App. C, Thm. C.24
$\sigma=\ell_{P}$ (gravitational) Combining massless cancellation with $\alpha=1$ Theorem
(this work)
Above inputs App. C, Thm. C.6
Axioms A1–A4 Formal axiomatic structure for stochastic gravity Postulate §4.4
Wiener character Lévy–Khintchine + path continuity Theorem A1 App. D, Thm. D.3
Drift uniqueness Lovelock classification + classical correspondence Theorem A3 App. D, Thm. D.12
Emergent time Quadratic variation; drift-independent by Itô calculus Theorem A2 App. D, Thm. D.10
Amplitude uniqueness Three independent arguments converge on $\sigma=\ell_{P}$ Theorem A1–A4 App. D, Thm. D.18
4.1 The $\sigma$-Continuum as Empirical Evidence for Universal Stochastic Structure
The nine-domain validation of §3 establishes that the stochastic differential equation $\mathrm{d}X=\mu(X)\,\mathrm{d}t+\sigma(X)\,\mathrm{d}W$ is not solely a mathematical device but a recurring structural element across physical domains. A single non-parametric pipeline, receiving only raw time series and sampling interval, recovers the governing equations of classical mechanics, statistical mechanics, nuclear physics, quantum mechanics, chemical kinetics, electromagnetism, relativistic quantum mechanics, and quantum field theory, with recovery errors quantified by bootstrap confidence intervals (Table 2) and without domain-specific input.
The recovered drift $\hat{\mu}$ and diffusion $\hat{\sigma}$ are not phenomenological regressions. They contain the fundamental constants of each domain: $k_{B}$ in Brownian motion, $\hbar$ in wavepacket spreading, $c$ in electromagnetic propagation, the Van Kampen exponent in chemical kinetics. The constants are not inserted. They emerge from the local geometry of the correlation manifold.
The $\sigma$-continuum (Fig. 1) is an observation arising from these nine independent experiments. Whether the pattern extends to gravitational degrees of freedom, and what determines $\sigma$ for gravity, is addressed in the following sections.
4.2 Relativistic Structure in the Diffusion Coefficient
The C6–C7 transition provides the first structural constraint. The Dambis–Dubins–Schwarz time-change theorem [27], applied to Nelson’s stochastic mechanics in a relativistic setting, yields a kinematic identity: the diffusion coefficient in the laboratory frame is suppressed by the Lorentz factor,
$\sigma_{\mathrm{KG}}=\frac{\sigma_{0}}{\gamma},$ (6)
where $\sigma_{0}=\sqrt{\hbar/m}$ is the rest-frame Nelson coefficient and $\gamma=(1-v^{2}/c^{2})^{-1/2}$. This suppression arises from two structurally independent factors: a kinematic $1/\sqrt{\gamma}$ from the DDS time change, and a dynamical $1/\sqrt{\gamma}$ from the Klein–Gordon [28, 29] conserved density $\rho_{\mathrm{KG}}=\gamma|\phi|^{2}$. The complete derivation, including the Fokker–Planck equation in both proper and laboratory frames, is given in Appendix B.
The pipeline recovers the $1/\gamma$ suppression to 0.72% (Fig. 8d) and the group velocity to $R^{2}=0.9999$ (Fig. 8f). The independent recovery of $v_{g}$ at each momentum mitigates a potential circularity: because $\gamma$ depends on $v$, the suppression $\sigma\propto 1/\gamma$ is validated only if the velocity is itself recovered independently, not assumed. Both channels, drift and diffusion, carry relativistic content. The speed of light $c$ enters $\sigma$ through the dispersion relation, alongside $\hbar$ and $m$.
Five structural observations follow from C6–C7 taken together. Observations 1, 4, and 5 are empirical, arising from the pipeline results. Observations 2 and 3 are mathematical consequences of theorems proved in Appendix B (Theorem B.5 and Proposition B.9).
First, $\sigma$ is not a domain label but a dynamical quantity: within C7, $\sigma$ varies continuously with $\gamma$ across seven momenta, and the pipeline tracks this variation to sub-percent accuracy (Fig. 8d,h). The robustness of the scaling law under reparameterisation from $\gamma$-space to momentum space (Fig. 8h) confirms that the suppression reflects a genuine dynamical relationship, not a coordinate artefact.
Second, the relativistic modification of $\sigma$ is structural, not kinematic. It arises from two independent mechanisms (the DDS time change and the Klein–Gordon density) that combine multiplicatively. This is not a Lorentz boost of a scalar, but a consequence of how probability density transforms under relativistic dynamics.
Third, classical electromagnetism (C6, $\sigma=0$) is recovered as the massless limit of the same suppression law: as $\gamma\to\infty$, $\sigma_{\mathrm{KG}}=\sigma_{0}/\gamma\to 0$. Determinism is not a separate category but the high-$\gamma$ limit of stochasticity.
Fourth, the C6–C7 transition demonstrates that a single mathematical structure, the stochastic differential equation with state-dependent $\sigma$, accommodates both the deterministic and stochastic domains, with the boundary between them set by the Lorentz factor.
Fifth, the progression from C1 ($\sigma=0$) through C2–C5 ($\sigma>0$, non-relativistic) to C7 ($\sigma=\sigma_{0}/\gamma$, relativistic) constitutes an empirically observed hierarchy. The subsequent derivation for gravitational degrees of freedom (§4.3) continues this hierarchy through the Gravitational Diffusion Theorem (Appendix C, Theorem C.6): the fluctuation–dissipation relation, the massless cancellation, and the self-coupling rate determined by the equivalence principle.
4.3 From Quantum Fields to Gravitational $\sigma$
The Gravitational Diffusion Theorem (Appendix C, Theorem C.6). For any field variable $X$ describing gravitational degrees of freedom, the diffusion coefficient is $\sigma=X_{P}/\sqrt{t_{P}}$, where $X_{P}$ is the Planck unit of $X$. For proper-distance fluctuations: $\sigma=\ell_{P}/\sqrt{t_{P}}$.
The proof proceeds in three stages, each building on validated results.
The C8–C9 progression reveals how the three fundamental constants $\{\hbar,G,c\}$ acquire distinct roles in the diffusion coefficient.
In C8 (quantum harmonic oscillator at zero temperature), the fluctuation–dissipation theorem gives $\sigma^{2}=\gamma\hbar\omega/m$. The Planck constant $\hbar$ enters explicitly through the zero-point energy $\tfrac{1}{2}\hbar\omega$, where its role is to quantise the field.
In C9 (quantised electromagnetic field), mass–energy equivalence gives $m_{\mathrm{eff}}=\hbar\omega/c^{2}$. Substitution into the fluctuation–dissipation relation yields the massless cancellation (Eq. 5): $\sigma=c\sqrt{\gamma}$, with $\hbar$ and $\omega$ cancelling identically. The speed of light $c$ sets the velocity scale, where the damping rate $\gamma$ sets the coupling strength.
For the photon, $\gamma$ is an external parameter (cavity losses, material absorption). For gravity, the coupling is not external. The gravitational field is self-interacting, and the only Planck-scale rate constructible from $\{\hbar,G,c\}$ without introducing additional scales is
$\gamma_{\mathrm{grav}}=\frac{1}{t_{P}},$ (7)
where $t_{P}=\sqrt{\hbar G/c^{5}}$ is the Planck time. The self-coupling rate $\gamma_{\mathrm{grav}}=1/t_{P}$ is not merely the unique dimensionally consistent rate. It is derived from the Mori–Zwanzig projection of the superspace Fokker–Planck equation (Appendix C, Proposition C.1): the gravitational field acts as its own bath, with the self-coupling rate determined by the cubic vertex of the Einstein–Hilbert action. The dimensional argument identifies the correct answer; the Mori–Zwanzig projection explains why it is correct. Substitution into $\sigma=c\sqrt{\gamma}$ gives
$\sigma=c\sqrt{\frac{1}{t_{P}}}=\frac{\ell_{P}}{\sqrt{t_{P}}}=\frac{\ell_{P}}{t_{P}^{1/2}},$ (8)
where $\ell_{P}=\sqrt{\hbar G/c^{3}}$ is the Planck length. This corresponds to a proper-distance fluctuation of one Planck length per Planck time.
In summary, the three stages are as follows. Stage 1: the quantum harmonic oscillator (C8) establishes the fluctuation–dissipation relation $\sigma^{2}=\gamma\hbar\omega/m$ (validated: $0.27\%$ error from 50 independent runs, Table 2). Stage 2: for a massless field with $m_{\mathrm{eff}}=\hbar\omega/c^{2}$, both $\hbar$ and $\omega$ cancel identically, giving $\sigma=c\sqrt{\gamma}$, the massless cancellation (validated: C9 photon field, $0.24\%$ error, constant $\sigma$ to within 1% across seven frequencies). Stage 3: for gravity, the self-coupling rate $\gamma_{\mathrm{grav}}=1/t_{P}$ is the unique Planck-scale rate, giving $\sigma=\ell_{P}/\sqrt{t_{P}}$. The epistemic boundary between Stages 2 and 3 is precise: the C9 pipeline results validate the mathematical formula $\sigma=c\sqrt{\gamma}$ for an externally damped massless field. Stage 3 applies the same formula to a self-coupled field, where the damping rate is not external but is determined by the equivalence principle, the most precisely tested input in the derivation chain. The self-coupling mechanism is derived, not assumed, in Appendix C (Proposition C.1 and Theorem C.16); the pipeline validates the formula, while the appendix validates the physical mechanism.
The dimensionless prefactor $\alpha$ in $\gamma_{\mathrm{grav}}=\alpha/t_{P}$ is determined by three independent arguments (Appendix C, §C.15). Metric self-consistency requires $\langle h^{2}\rangle=\alpha^{2}\leq 1$ over one Planck time, giving the upper bound $\alpha\leq 1$. The critical damping condition: the unique frequency at which $\gamma(\omega)=t_{P}^{2}\omega^{3}$ equals $\omega$, yields $\omega^{*}=1/t_{P}$ exactly, giving $\alpha=1$ as an algebraic identity. Singularity resolution requires that curvature fluctuations reach the Planck curvature at the minimum scale, giving $\alpha\geq 1$. Together: $\alpha\leq 1\wedge\alpha=1\wedge\alpha\geq 1\Rightarrow\alpha=1$ uniquely. The three arguments share a single input, the self-coupling rate $\gamma(\omega)=t_{P}^{2}\omega^{3}$ (Theorem C.16), but are otherwise logically independent: Argument 1 invokes metric positivity, Argument 2 invokes the oscillator–diffusion boundary, and Argument 3 invokes curvature scaling. No two arguments use the same physical principle beyond the shared input. The complete derivation chain is developed in Appendix C.
4.4 Axiomatic Formalisation
The axioms require a shift in the role of the stochastic differential equation. In §2–§3, the stochastic differential equations describe matter evolving in a fixed spacetime, and the evolution parameter is physical time measured by a laboratory clock. From this point forward, the stochastic differential equation describes spacetime itself, and the evolution parameter $\tau$ is a pre-geometric index: a label ordering the steps of a Wiener process on Wheeler’s superspace. Physical time emerges as a derived quantity, constructed from the quadratic variation of the metric process (§D.5). The ordering parameter $\tau$ is not the time of any clock; clocks are built from the metric, which is the stochastic variable. This distinction is the framework’s resolution of the problem of time in canonical quantum gravity.
Appendix C derives $\sigma=\ell_{P}$ from three established results and the empirical pattern of C1–C9. The four canonical axioms below formalise this derivation into a self-contained axiomatic structure; Axiom A2 is a derived result promoted to axiomatic status for deductive economy (Appendix C, Remark following the Gravitational Diffusion Theorem).
A1 (Configuration Space)
The gravitational degrees of freedom are described by a Riemannian three-metric $g_{ij}$ on a spatial manifold $\Sigma$. The configuration space is Wheeler’s superspace $\mathcal{C}=\mathrm{Riem}(\Sigma)/\mathrm{Diff}(\Sigma)$ [30], equipped with the DeWitt supermetric [31].
A2 (Stochastic Evolution)
The metric undergoes a stochastic process on $\mathcal{C}$:
$\mathrm{d}g_{ij}=\mathcal{D}_{ij}[g]\,\mathrm{d}\tau+\ell_{P}\,\mathrm{d}W_{ij},$ (9)
where $\mathcal{D}_{ij}[g]$ is a drift functional, $\ell_{P}$ is the Planck length [26], and $\mathrm{d}W_{ij}$ is a symmetric tensor-valued Wiener process. The ordering parameter $\tau$ is not a background time coordinate.
A3 (Classical Correspondence)
In the macroscopic limit ($L\gg\ell_{P}$), the stochastic evolution reproduces general relativity [21]: $R_{\mu\nu}-\tfrac{1}{2}g_{\mu\nu}R+\Lambda g_{\mu\nu}=(8\pi G/c^{4})\,T_{\mu\nu}$.
A4 (Single Realisation)
The observable universe is one realisation of the stochastic process. Probability is epistemic.
Physical time is not assumed. It emerges as the quadratic variation of the stochastic process: $t(\tau)=(1/\ell_{P}^{2}n_{\mathrm{eff}})\int\langle\mathrm{d}g,K^{-1}\mathrm{d}g\rangle$ (Appendix D, Theorem D.10). The effective mode count $n_{\mathrm{eff}}$ is formally divergent on non-compact manifolds; it is rendered finite by a Planck-scale wavelength cutoff ($n_{\mathrm{eff}}\sim V/\ell_{P}^{3}$) or by minisuperspace reduction ($n_{\mathrm{eff}}=1$ for FLRW), and is absorbed into the normalisation of $t(\tau)$ so that no physical prediction depends on $n_{\mathrm{eff}}$ separately (Appendix D, Remark D.12). By the Itô product rules, $(\mathrm{d}\tau)^{2}=0$, $\mathrm{d}\beta_{n}\mathrm{d}\tau=0$, $\mathrm{d}\beta_{m}\mathrm{d}\beta_{n}=\delta_{mn}\mathrm{d}\tau$, the drift $\mathcal{D}_{ij}$ contributes identically zero to the quadratic variation. The emergent time functional is therefore strictly monotone (from positivity of $K$), additive (from independent increments), coordinate-independent (from tensorial covariance), and exactly drift-independent (from the Itô calculus). In the macroscopic limit $L\gg\ell_{P}$, the emergent time coincides with proper time through the standard ADM relation $\mathrm{d}t_{\mathrm{proper}}=N\,\mathrm{d}t$, with stochastic corrections of $O(\ell_{P}^{2}/L^{2})\sim 10^{-70}$ for astrophysical scales. This construction addresses the problem of time in canonical quantum gravity: time is not a background coordinate but a derived quantity, constructed from the diffusion itself, with the deterministic gravitational dynamics contributing nothing to its rate.
Within axioms A1–A4, every element of Eq. (9) is uniquely determined.
Noise character. The Lévy–Khintchine classification theorem establishes that any stochastic process with stationary independent increments and continuous sample paths is a Wiener process (plus drift). Path continuity is a physical consequence of Axiom A1: a jump discontinuity in the metric would generically violate positive-definiteness of $g_{ij}$ or produce infinite curvature. The Wiener process is therefore derived from the remaining structure, not postulated independently (Appendix D, Theorem D.3).
Drift. Within the class of ultralocal, second-order, diffeomorphism-covariant functionals that reproduce the Einstein evolution in the classical limit (A3), the drift is uniquely the Einstein flow. This invokes the Lovelock classification: in three spatial dimensions, the unique divergence-free symmetric two-tensor constructed from the metric and its derivatives is the Einstein tensor.
Covariance. The covariance operator $K_{ijkl}$ must be ultralocal, diffeomorphism-covariant, and positive-definite on symmetric two-tensors. The unique such family is the DeWitt supermetric $G^{ijkl}(\lambda)$ with $\lambda>-1/3$ [31].
Amplitude. The fluctuation amplitude $\sigma=\ell_{P}$ is fixed by the three independent arguments of Appendix C (§C.15): metric self-consistency ($\alpha\leq 1$), critical damping ($\alpha=1$), and singularity resolution ($\alpha\geq 1$).
The complete axiomatic derivation, including the uniqueness theorems for all four elements, is given in Appendix D.
Postulate D.3 (Minimality) is not an additional dynamical assumption beyond A1–A4 and path continuity. The Lévy–Khintchine classification of processes with stationary independent increments, combined with the path-continuity requirement on metric configurations imposed by Axiom A1, uniquely selects the Wiener process from the class of all Lévy processes (Appendix D, Theorem D.3). The Markov property and spatial locality of the noise follow respectively from the emergent-time construction and cluster decomposition (Appendix D, Remarks D.6 and D.7). The resulting Fokker–Planck evolution on superspace admits both a formal Wick-rotation correspondence with the Wheeler–DeWitt equation of canonical quantum gravity and a Martin–Siggia–Rose path-integral representation whose saddle-point limit recovers classical general relativity; physical observables are shown to be independent of the foliation (lapse) choice (Appendix D, §D.4.2 and §D.4.3).
Within axioms A1–A4, Eq. (9) is non-parametric: all coefficients are derived from the axioms.
4.5 The Superspace Diffusion Framework
The derivations that follow from axioms A1–A4 constitute the Superspace Diffusion Framework. The superspace diffusion hypothesis (Eq. 9) is one consequence. At galactic scales, the Mori–Zwanzig coarse-graining of the superspace Fokker–Planck equation (Proposition C.1) yields non-parametric predictions for the statistical distribution of gravitational acceleration; these consequences are developed in subsequent work.
The Superspace Diffusion Framework extends the $\sigma$-continuum from an empirical observation to a formal structure through a theorem chain (Appendix C) derived from the fluctuation–dissipation theorem, the equivalence principle, and the massless mode structure of linearised gravity.
4.6 Relation to Nelson’s Stochastic Mechanics
Nelson [6] solved the forward problem for quantum mechanics: given $\sigma=\sqrt{\hbar/m}$, derive the Schrödinger equation. The present work solves the inverse problem across nine domains: given time series data, recover $\sigma$ without assuming the physics. The forward and inverse paths converge at C4, where the pipeline recovers Nelson’s $\sigma$ to $0.05\%$ (Table 2). Nelson’s insight, that physical law admits stochastic content [24], is here extended across the $\sigma$-continuum to gravitational degrees of freedom.
4.6.1 Distinguishing the stochastic and quantum formulations
The Wick rotation correspondence (Appendix D, Theorem D.4) maps the Fokker–Planck generator on superspace to the Wheeler–DeWitt Hamiltonian through the analytic continuation $\tau\to-it/\hbar$. This correspondence is exact at the level of the generator and establishes that the stochastic framework is connected to canonical quantum gravity by the same analytic structure that relates Euclidean and Lorentzian quantum field theory.
The correspondence does not, however, extend to the level of solutions and physical predictions. Three structural properties of the Fokker–Planck evolution have no counterparts in the Wheeler–DeWitt equation.
First, the Fokker–Planck equation possesses a unique normalisable steady state. For any Fokker–Planck equation with non-degenerate diffusion and confining drift on a half-line or bounded domain, the zero-current condition $\partial_{g}[D\cdot P]=A\cdot P$ determines a unique normalisable distribution. The superspace Fokker–Planck equation (Appendix D, Eq. D.165), when coarse-grained to galactic-scale gravitational degrees of freedom via the Mori–Zwanzig projection (Appendix C, Proposition C.1), inherits this property: its steady state is uniquely determined by the drift and diffusion derived from A1–A4. The Wheeler–DeWitt equation $\hat{H}\Psi=0$ is a constraint equation: it has no time parameter, no spectral gap, and no mechanism for selecting a unique solution from its solution space.
Second, the Fokker–Planck evolution is dissipative. It drives any initial probability distribution toward the unique steady state on a timescale set by the spectral gap. This irreversible relaxation produces definite predictions: regardless of initial conditions, the distribution converges to the steady state. Quantum evolution preserves unitarity: pure states remain pure and the evolution is time-reversible: producing recurrences rather than relaxation.
Third, the Fokker–Planck steady state depends on the positivity of $P[g]\geq 0$ and the real-valued character of the diffusion coefficient. Under the Wick rotation, $P$ maps to a complex-valued wave functional $\Psi[g]$ that is generically oscillatory and not normalisable in the standard $L^{2}$ sense. The positive, real-valued, normalisable steady state of the Fokker–Planck equation has no natural image under this mapping.
These three properties: uniqueness, irreversible relaxation, and positivity: are consequences of the diffusion character of the evolution. The Wick rotation preserves the generator but not these properties, in the same way that analytic continuation between Euclidean and Lorentzian field theory preserves the action but not the metric signature.
The framework therefore proposes a specific physical hypothesis: that the gravitational degrees of freedom evolve by diffusion on superspace, governed by the Fokker–Planck equation (Axiom A2), rather than by unitary quantum evolution governed by the Wheeler–DeWitt equation. The two formulations share a generator and a classical limit (general relativity, recovered as $L\gg\ell_{P}$ in both cases) but make distinct predictions for the statistical distribution of gravitational acceleration at galactic scales. The observational confrontation of these predictions with galactic kinematic data is the subject of subsequent work.
4.6.2 Relationship to conventional approaches to quantum gravity
The stochastic formulation occupies a specific position relative to the three standard approaches.
In canonical quantum gravity, the Wheeler–DeWitt equation $\hat{H}\Psi=0$ defines the quantum state of the gravitational field. The Fokker–Planck equation on superspace is related to it by the Wick rotation correspondence (Appendix D, Theorem D.4). At the level of generators, the two are formally equivalent. At the level of solutions, they differ: the Fokker–Planck equation has a unique steady state reached by irreversible relaxation, where the Wheeler–DeWitt equation has a solution space parameterised by boundary conditions.
In Euclidean quantum gravity, the gravitational path integral defines transition amplitudes between three-geometries. The Martin–Siggia–Rose path integral (Appendix D, Theorem D.5) establishes that the stochastic framework contains the Euclidean gravitational path integral as its saddle-point approximation: the classical constraint surface (Einstein’s equations) is the saddle-point locus, with stochastic excursions of amplitude $O(\ell_{P})$ (Appendix D, Corollary D.172). The stochastic framework thus extends the Euclidean programme beyond the saddle-point approximation, providing a well-defined measure on the space of metrics that the Euclidean path integral, which suffers from the conformal factor problem, does not possess.
In stochastic quantisation à la Parisi–Wu [32], quantum field theory is obtained as the equilibrium distribution of a diffusion in a fictitious stochastic time. The present framework shares the mathematical structure but differs in a physical identification: the ordering parameter $\tau$ is identified with physical time through the emergent time theorem (Appendix D, Theorem 1), not treated as a fictitious coordinate. The equilibrium distribution is the physical steady state of the gravitational field, not a computational device for reproducing the Euclidean path integral.
The stochastic formulation is therefore not a reformulation of any existing approach to quantum gravity, but a distinct theoretical framework that shares their classical limit and connects to each through specific mathematical correspondences.
4.7 Limitations
All C1–C9 data are synthetic. The pipeline has not been applied to raw experimental data in this work. The synthetic data validates the methodology, where experimental application is the subject of subsequent publications. The blind protocol ensures that the pipeline cannot use domain knowledge, but the selection of the nine test domains is informed by the authors’ knowledge of physics; the pipeline’s performance on domains outside this selection is an open question.
The non-parametric $k$-NN Kramers–Moyal estimator carries a systematic positive bias of $O(k/N)$ in the recovered diffusion coefficient, ranging from $0.2\%$ (C8, C9) to $5.9\%$ (C3) across the nine domains. This bias is a mathematical property of local regression estimators, not a deficiency of the pipeline. It cancels in scaling exponents (demonstrated by C5) and is approximately uniform across state space for a given domain (demonstrated by the constant $\hat{\sigma}/\sigma_{\mathrm{true}}$ ratio across Lorentz factors in C7). The bias can be reduced by increasing $N$ or decreasing $k$, at the cost of increased variance. For the present study, the pipeline parameters ($d$, $k$, NN) were fixed per domain and not optimised to minimise the bias.
The gravitational prediction $\sigma=\ell_{P}$ ($\sim\!10^{-35}$ m) is not directly testable with current technology. Its consequences, developed in subsequent papers, generate quantitative predictions at accessible scales.
The axioms A1–A4 are stated as postulates for deductive clarity, though each has independent justification. A1 adopts the standard canonical general relativity configuration space (Wheeler’s superspace [30]). A2 is a derived result (the Gravitational Diffusion Theorem, Appendix C) promoted to axiomatic status for deductive economy. A3 requires consistency with general relativity in the macroscopic limit. A4, the identification of probability as epistemic, is the sole genuinely free interpretive choice. Within A1–A4, the noise character, drift, covariance, fluctuation amplitude, and emergent time are all uniquely determined by theorem (Appendix D). Alternative axiomatisations may exist, but the physical content of the predictions follows from the derivation chain of Appendix C independently of the axiomatic packaging. The empirical question is therefore whether A1–A4 hold, not whether the predictions follow from them. If the axioms are incorrect, the predictions that follow from them need not be physically operative. This question is addressable by confronting the framework’s galactic-scale predictions with kinematic data.
To delimit scope explicitly: this paper does not claim to have derived quantum gravity, unified general relativity with quantum mechanics, or proved the axioms A1–A4. What is claimed is that if A1–A4 hold, then every element of the superspace SDE is uniquely determined as a non-parametric, first-principles consequence of the axioms. Whether A1–A4 are physically operative is an empirical question that can be addressed by observational data at galactic scales.
What the derivation chain does establish is that the gravitational diffusion coefficient is uniquely determined (the Gravitational Diffusion Theorem, Appendix C) by the same physical principles (the fluctuation–dissipation theorem, mass–energy equivalence, and the equivalence principle) that govern the nine validated domains, as a non-parametric, first-principles consequence of general relativity and quantum field theory.
5 Conclusion
The Stochastic Embedding Sufficiency Theorem closes a methodological gap between deterministic and stochastic time series analysis, enabling recovery of drift and diffusion fields from scalar observations without prior assumptions about the governing physics. The data determines the non-parametric model structure.
Application of the resulting pipeline to nine physical domains (classical mechanics, statistical mechanics, nuclear physics, quantum mechanics, chemical kinetics, electromagnetism, relativistic quantum mechanics, quantum harmonic oscillator dynamics, and quantum electrodynamics) recovers the known governing equations of each domain blindly, with recovery errors ranging from $3\times 10^{-6}\%$ (deterministic domains) to approximately $6\%$ (absolute diffusion recovery at high $k/N$) and bootstrap confidence intervals quantifying the systematic $k$-NN estimator bias. For scaling exponents (C5), the multiplicative bias cancels and the bootstrap CI contains the theoretical value. Fundamental physical constants emerge in both the drift and the diffusion channels without prior specification. The recovered diffusion coefficients, viewed together, constitute an empirical pattern: the $\sigma$-continuum.
Three theorems derived from the relativistic and quantum field theory results yield three structural consequences. The derivation chain rests on the fluctuation–dissipation theorem (validated by C8 to $0.27\%$), mass–energy equivalence, and the equivalence principle, each among the most precisely tested results in physics. First, the Dambis–Dubins–Schwarz time-change theorem combined with the Klein–Gordon density produces the relativistic suppression $\sigma=\sigma_{0}/\gamma$, with physical constants emerging in both channels (Appendix B). Second, the massless cancellation and gravitational self-coupling together determine $\sigma=\ell_{P}/\sqrt{t_{P}}$ for gravitational degrees of freedom, with three independent consistency arguments converging on the dimensionless prefactor $\alpha=1$ (Appendix C). Third, a set of four canonical axioms (A1–A4) yields a framework within which physical time, the drift, the covariance operator, and the fluctuation amplitude are all uniquely fixed (Appendix D).
The resulting superspace diffusion hypothesis,
$\mathrm{d}g_{ij}=\mathcal{D}_{ij}[g]\,\mathrm{d}\tau+\ell_{P}\,\mathrm{d}W_{ij},$ (10)
is non-parametric within A1–A4: the noise character, drift, covariance, fluctuation amplitude, and emergent time are all uniquely determined by theorem (Appendix D). The empirical question is therefore whether A1–A4 hold, not whether the predictions follow from them.
The Superspace Diffusion Framework defined by axioms A1–A4 is falsifiable. The Fokker–Planck equation on superspace, coarse-grained via the Mori–Zwanzig projection to galactic-scale gravitational degrees of freedom, generates non-parametric, first-principles predictions for the statistical distribution of gravitational acceleration. These predictions are testable against galactic kinematic data including rotation curves, velocity dispersions, wide binaries, galaxy clusters, and gravitational lensing. If the axioms are incorrect, the predictions derived from them need not be physically operative: the observational confrontation therefore tests the axioms themselves, not merely their mathematical consequences.
Acknowledgements
The authors are indebted to UCL’s Manufacturing Futures Laboratory for computational resources. Generative AI accelerated and extended the authors’ technical reach, while responsibility, conceptual synthesis and structural design remained human. As such, the contribution of Anthropic’s Claude as an epistemic tool is gratefully acknowledged.
Author Contributions
Carolina Garcia: Investigation, Writing (original draft), Writing (review and editing). Lucía Perea Durán: Investigation, Writing (original draft), Writing (review and editing). Agnese Venezia: Investigation, Writing (original draft), Writing (review and editing). Alex Conradie: Conceptualisation, Methodology, Investigation, Formal Analysis, Software, Supervision, Visualisation, Writing (original draft), Writing (review and editing).
Competing Interests
The authors declare no competing interests.
Figure 1: The $\sigma$-Continuum. Recovered diffusion coefficient $\sigma$ across all nine physical domains and the derived gravitational value, displayed on a logarithmic vertical axis ($\log_{10}\sigma$, schematic units). Each domain is represented by a distinct marker shape and colour, labelled with its recovered $\sigma$ expression. Deterministic domains (C1: classical mechanics, dark blue circle; C6: classical electromagnetism, light blue circle) are placed at $\log_{10}(\sigma)\approx-40$ for visualisation, labelled $\sigma=0$. Stochastic domains span intermediate scales: C2 (statistical mechanics, green square, $\sigma=\sqrt{2\gamma k_{B}T/m}$), C3 (radioactive decay, orange diamond, $\sigma^{2}=\mu$), C4 (quantum mechanics, purple pentagon, $\sigma=\sqrt{\hbar/m}$), C5 (chemical kinetics, red triangle, $\sigma\propto\Omega^{-1/2}$). The relativistic and quantum field theory domains occupy the upper region: C7 (Klein–Gordon, red–brown pentagon, $\sigma=\sqrt{\hbar/m}/\gamma$), C8 (quantum harmonic oscillator, brown diamond, $\sigma=\sqrt{\gamma\hbar\omega/m}$), C9 (QED photon, magenta circle, $\sigma=c\sqrt{\gamma}$). The dashed lilac horizontal line marks the quantum scale $\sigma\sim\sqrt{\hbar/m}$; the dashed gold horizontal line marks the Planck scale $\ell_{P}\approx 10^{-35}$ m. The gold star at the far right represents $\sigma=\ell_{P}$ for gravitational degrees of freedom, derived by the Gravitational Diffusion Theorem (§4; Appendix C). The continuum from $\sigma=0$ (C1, C6) through thermal, quantum, and relativistic scales (C2–C9) is an empirical pattern arising from nine independent blind recovery experiments; the gravitational endpoint $\sigma=\ell_{P}$ is derived from this pattern by the Gravitational Diffusion Theorem.
[…]
Figure 5: C3: Radioactive Decay ($\sigma^{2}=\mu$, Fano =1=1). (a) Poisson count series at $\mu=3000$ (blue) over 500 measurements, with the 250-realisation ensemble median (green) and 95% CI band (green shading). Inset: probability densities of the data (blue), ensemble (green), and theoretical Poisson–Gaussian approximation (red). (b) Recovered $\hat{\sigma}^{2}$ (pipeline) versus mean count $\mu$ on log–log axes (magenta circles) across seven mean count levels from $\mu\approx 100$ to $\mu\approx 10\,000$. The dashed red line is the Poisson prediction $\sigma^{2}=\mu$. The log–log slope is $1.0002$ ($R^{2}=0.9997$). (c) Observed count distribution at $\mu=3000$ (blue histogram) overlaid with the fitted Gaussian ($\mu=3001$, $\sigma=55.1$; red curve). The measured Fano factor is $1.0109$, consistent with Poisson statistics ($F=1$). (d) Autocorrelation function of the count series at 50 lags. The red dashed lines mark the 95% significance band ($\pm 0.0196$). Three of 50 lags exceed the band ($\sim$2 expected by chance), confirming temporal independence, a necessary property of Poisson counting. (e) Distribution of Fano factor $F=\hat{\sigma}^{2}/\hat{\mu}$ from 50 independent simulations at $\mu=3000$; median =1.0000=1.0000 (green line), run-level 95% range $[0.9612,1.0295]$. This panel shows the raw variance/mean ratio; through the pipeline, $\hat{\sigma}/\sqrt{\mu}=1.059$ with bootstrap 95% CI [1.051, 1.062][1.051,\,1.062] (Table 2), reflecting the $O(k/N)$ estimator bias that cancels in the scaling-law test of panel (f). (f) Variance $\hat{\sigma}^{2}$ versus mean $\mu$ from 50-simulation medians at seven count levels (magenta circles with 95% CI); the log–log slope is $0.9992$ ($R^{2}=1.0000$), confirming the Poisson relation $\sigma^{2}=\mu$ across two decades.
Figure 6: C8: Quantum Harmonic Oscillator ($\sigma=\sqrt{\gamma\hbar\omega/m}$). (a) Simulated velocity time series $v(t)$ of the damped QHO (blue, in physical units) over $\sim$300 ps, with the pipeline-recovered local drift (red, $k$-NN with embedding dimension $m^{*}=5$). The drift smoothly tracks the oscillatory dynamics, capturing the reversal near $t\approx 150$ ps. (b) Probability density of locally recovered $\hat{\sigma}(x)$ values across the correlation manifold. The distribution is approximately Gaussian; the pipeline median $\hat{\sigma}=1.02\times 10^{5}$ (green line) agrees with the theoretical value $\sigma=\sqrt{\gamma\hbar\omega/m}=1.03\times 10^{5}$ (red dashed) to $0.35\%$. With the embedding lag set to the autocorrelation zero-crossing ($\tau_{\mathrm{ACF}}$), 50 independent runs yield $\hat{\sigma}/\sigma_{\mathrm{true}}=0.997$ with bootstrap 95% CI [0.995, 1.000][0.995,\,1.000] (Table 2). (c) Recovered $\hat{\sigma}$ versus theoretical $\sigma=\sqrt{\gamma\hbar\omega/m}$ on log–log axes across nine frequencies spanning one decade. Blue circles lie on the identity line (green, “Perfect agreement”); $R^{2}=0.9999$. The linear dependence $\sigma\propto\sqrt{\omega}$ confirms the fluctuation–dissipation relation at the quantum level. (d) Relative error (%) for 20 independent random seeds at the primary frequency. All errors lie below $1\%$; the mean error (blue dotted line) is $\sim$0.47%. The green dashed line marks zero error. The 50-run bootstrap analysis (Table 2) gives $0.27\%$ error.
Figure 7: C5: Van Kampen Scaling ($\sigma\propto\Omega^{-1/2}$). (a) Concentration time series $c=n/\Omega$ at system size $\Omega=500$ (blue) over 60 time units, with the 250-realisation ensemble median (green) and 95% CI band (green shading). Inset: probability densities of the true (blue) and forecast (green) concentration, showing close agreement. (b) Recovered diffusion profile $\hat{\sigma}(c)$ versus concentration (cyan scatter) with the theoretical state-dependent curve $\sigma(c)=\sqrt{(k_{b}/\Omega+k_{d}c)/\Omega}$ (red dashed). The scatter reflects local $k$-NN estimation; the theoretical curve passes through the centre of the cloud. (c) Recovered drift $\hat{\mu}(c)$ versus concentration (yellow scatter) with the theoretical mean-reversion $\mu(c)=k_{b}/\Omega-k_{d}c$ (red dashed). The equilibrium concentration $c_{\mathrm{eq}}=10$ is annotated; the drift changes sign at this point, confirming mean-reverting dynamics. (d) Median recovered $\hat{\sigma}$ versus system size $\Omega$ on log–log axes (cyan circles with 95% CI from 50 runs at each $\Omega$). The fitted slope $\alpha=-0.502\pm 0.001$ (solid pink) matches the Van Kampen prediction $\alpha=-0.500$ (dotted black); $R^{2}=1.0000$. (e) Distribution of the scaling exponent $\hat{\alpha}$ from 50 independent pipeline runs; median =0.502=-0.502 (green line), theoretical value 0.500-0.500 (red dashed). Bootstrap 95% CI on median: [0.503,0.500][-0.503,\,-0.500], containing the theoretical value (Table 2). (f) Residuals $\Delta\log_{10}\hat{\sigma}$ from the fitted power law at each system size (brown circles). The RMS residual is $0.0011$ dex, confirming that the power-law relationship holds across two decades in $\Omega$.
[…]
Figure 9: C9: QED Photon Field ($\sigma=c\sqrt{\gamma}$; $\hbar$ and $\omega$ cancel). (a) Dimensionless velocity $\tilde{v}$ of the photon mode versus dimensionless time $\omega t$ (blue), with the pipeline-recovered local drift (red, $k$-NN with $m^{*}=5$). The drift smoothly tracks the oscillatory dynamics over approximately four periods. (b) Probability density of locally recovered $\hat{\sigma}(x)$ in physical units ($\times 10^{14}$). The distribution is approximately Gaussian; the pipeline median (green line) of $\sim$$9.45\times 10^{13}$ agrees with the theoretical value $c\sqrt{\gamma}=9.48\times 10^{13}$ (red dashed) to $0.35\%$ (this single realisation). With $\tau_{\mathrm{ACF}}$ embedding, 50 independent runs yield $\hat{\sigma}/\sigma_{\mathrm{true}}=0.998$ with bootstrap 95% CI [0.996, 0.999][0.996,\,0.999] (Table 2). (c) The $\omega$-independence test. Recovered $\hat{\sigma}$ (blue circles) at seven frequencies spanning $\omega\in[10^{12},10^{13}]$ rad/s on a logarithmic frequency axis. The theoretical value $c\sqrt{\gamma}=9.480\times 10^{13}$ is marked (green horizontal line) with a $\pm 1\%$ band (green shading). All seven recovered values lie within the band and show no systematic trend with $\omega$, confirming the massless cancellation $\sigma=c\sqrt{\gamma}$. (d) Relative error (%) for 20 independent random seeds at the primary frequency $\omega=10^{13}$ rad/s (purple bars). The green dashed line marks zero error; the blue dotted line marks the mean error of $\sim$$0.50\%$. The 50-run bootstrap analysis (Table 2) gives $0.24\%$ error.
Appendix, Mathematical Proofs and Derivations
Appendix A Derivation of Stochastic Embedding Sufficiency Theorem
A.1 Introduction
Classical delay-coordinate embedding theory, most notably Takens’ theorem, guarantees diffeomorphic reconstruction of deterministic attractors from time-delay observations. However, this guarantee does not hold when the underlying dynamics are stochastic: a fact operationalized by Cao’s $E_{2}$ statistic approaching unity. In such cases, distinct initial conditions can produce identical embedded observations under different noise realisations, rendering pathwise reconstruction impossible.
The central result established here is that pathwise reconstruction is not required for distributional inference. While stochastic systems do not admit unique trajectories on an attractor, they do admit well-defined probability laws on correlation manifolds. For the purposes of statistical inference on dynamical systems, it is sufficient that distinct system states induce distinct finite-dimensional probability distributions under delay embedding.
The Stochastic Embedding Sufficiency Theorem (Theorem A.1) formalises this approach. Rather than establishing diffeomorphic injectivity (as in Takens’ theorem), the result establishes measure-theoretic injectivity: for almost every point under the invariant measure, distinct points on the underlying correlation manifold correspond to distinct finite-dimensional laws of the embedded process. This guarantee is sufficient for consistent $k$-nearest neighbour estimation, graph-theoretic mixing assessment, and distributional divergence testing. The complete proof is given in Supplemental Materials; the following sections summarise the essential steps.
The proof proceeds in five conceptual stages:
  1. 1.
    Hörmander’s hypoellipticity ensures smooth, strictly positive transition densities for the SDE, via the bracket-generating condition and the Stroock–Varadhan support theorem.
  2. 2.
    Malliavin non-degeneracy guarantees that the Malliavin covariance matrix is invertible almost surely, preventing degeneracy of the embedded probability distributions.
  3. 3.
    Law-separation establishes that the law-embedding map $\Lambda_{m}^{h}$ is $\mu_{\infty}$-a.e. injective via three linked results: transition density separation (semigroup bisection + Varadhan–Léandre asymptotics), parametric transversality with an explicit evaluation-map surjectivity lemma, and a direct Frostman covering argument.
  4. 4.
    $E_{1}$ dimension sufficiency shows that Cao’s $E_{1}$ statistic detects correlation dimension $D_{2}$ irrespective of whether the system is deterministic or stochastic, yielding a sufficient embedding dimension $m^{*}\geq\lceil 2D\rceil+1$ (equivalently $\lceil 2D_{2}\rceil+1$ under exact-dimensionality).
  5. 5.
    Finite-dimensional law uniqueness ensures consistency of nearest-neighbour statistics and validates geometric comparison on the correlation manifold.
Together, these results extend the theoretical foundation from deterministic to stochastic systems.
A.2 The Methodological Divide
Classical dynamical systems analysis, exemplified by Takens’ embedding theorem [11] and its extensions [12], applies to deterministic systems. Given a scalar time series from a deterministic process, one can reconstruct the underlying attractor geometry through time-delay embedding. For a smooth diffeomorphism $\varphi$ on a compact $n$-dimensional manifold $M$ with generic observation function $h$, the delay embedding:
$\Phi_{m}(x)=(h(x),h(\varphi(x)),\ldots,h(\varphi^{m-1}(x)))^{\top}$ (A.11)
is a diffeomorphism onto its image for $m\geq 2n+1$.
Cao’s $E_{1}$ and $E_{2}$ diagnostics [15], extending the false nearest neighbours method [33], operationalise this framework: $E_{1}$ identifies the minimal embedding dimension, while $E_{2}$ tests for deterministic structure. When $E_{2}\approx 1$ across all dimensions, the system is stochastic, and classical embedding in the Takens sense does not apply.
This created a methodological gap: deterministic systems could be analysed geometrically, but stochastic systems required entirely different methods. No unified framework existed. The conventional interpretation was that if $E_{2}$ signals stochasticity, geometric methods are inapplicable.
A.3 Extension to Stochastic Systems
The theoretical contribution of this work addresses this gap through mathematical extension. The $E_{1}$-identified embedding, even when $E_{2}$ signals stochasticity, defines a correlation manifold encoding the system’s dynamical degrees of freedom. When stochasticity is detected, this manifold serves as the foundation for distributional analysis.
Definition A.1 (Delay-Vector Law and Correlation Manifold).
Assume $(X_{t})_{t\in\mathbb{R}}$ is stationary with invariant measure $\mu_{\infty}$ and observation function $h:\mathbb{R}^{n}\to\mathbb{R}$. Fix $\tau>0$ and $m\in\mathbb{N}$ and define the delay vector
$Y_{t}^{(m)}:=\bigl(h(X_{t}),\;h(X_{t-\tau}),\;\ldots,\;h(X_{t-(m-1)\tau})\bigr)\in\mathbb{R}^{m}.$ (A.12)
Let $\nu_{m}:=\mathcal{L}(Y_{0}^{(m)})$ denote its stationary law. The correlation manifold is the support of this law:
$\mathcal{M}_{D_{2}}:=\operatorname{supp}(\nu_{m})\subset\mathbb{R}^{m},$ (A.13)
where $D_{2}$ denotes the correlation dimension of $\nu_{m}$ (Eq. A.21). When $m\geq\lceil 2D_{2}\rceil+1$, $\mathcal{M}_{D_{2}}$ is a $D_{2}$-dimensional subset of $\mathbb{R}^{m}$ on which the embedded process has well-defined conditional expectations for the drift and diffusion fields.
The $E_{1}$ statistic detects correlation dimension $D_{2}$ regardless of whether the system is deterministic or stochastic. A stationary diffusion process possesses a well-defined correlation manifold $\mathcal{M}_{D_{2}}$, as does a deterministic chaotic attractor. The geometric structure persists; the dynamics upon it become probabilistic rather than deterministic.
A.4 Mathematical Framework
A.4.1 The Stochastic Dynamical System
Consider the Itô stochastic differential equation on $\mathbb{R}^{n}$:
$dX_{t}=\mu(X_{t})dt+\sigma(X_{t})dW_{t}$ (A.14)
where $\mu:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is the drift vector field, $\sigma:\mathbb{R}^{n}\to\mathbb{R}^{n\times r}$ is the diffusion coefficient matrix with columns $\sigma_{1},\ldots,\sigma_{r}$, $W_{t}$ is $r$-dimensional standard Brownian motion, and $\Sigma=\sigma\sigma^{\top}$ is the diffusion tensor. Time-delay embedding constructs:
$Y_{t}^{(m)}=(y_{t},y_{t-\tau},\ldots,y_{t-(m-1)\tau})^{\top}$ (A.15)
from observations $y_{t}=h(X_{t})$, where $h:\mathbb{R}^{n}\to\mathbb{R}$ is a smooth observation function.
Definition A.2 (Delay Embedding Map and Injectivity in Law).
The delay embedding map $\Phi_{m}:\mathbb{R}^{n}\to\mathbb{R}^{m}$ is defined by
$\Phi_{m}(x)=\bigl(h(x),\;h(\varphi_{\tau}(x)),\;\ldots,\;h(\varphi_{(m-1)\tau}(x))\bigr)^{\top},$
where $\varphi_{\tau}$ denotes the time-$\tau$ flow of the SDE (A.14). The delay representation is said to be injective in law (with respect to $\mu_{\infty}$) if, for $\mu_{\infty}$-almost every pair $(x,x^{\prime})$,
$x\neq x^{\prime}\;\;\Longrightarrow\;\;\mathcal{L}\!\left(Y_{0}^{(m)}\,\middle|\,X_{0}=x\right)\neq\mathcal{L}\!\left(Y_{0}^{(m)}\,\middle|\,X_{0}=x^{\prime}\right),$ (A.16)
where $\mathcal{L}(\cdot|\cdot)$ denotes the conditional law. This is the natural stochastic analogue of the diffeomorphic injectivity established by Takens’ theorem for deterministic systems: distinct states induce distinct probability distributions over delay vectors, even though individual realisations may coincide.
A.4.2 Cao’s $E_{1}$ and $E_{2}$ Statistics
For delay vectors with nearest neighbour index $n(i,m)$, the $E_{1}$ statistic is:
$E_{1}(m)=\frac{E(m+1)}{E(m)},\quad\text{where }E(m)=\frac{1}{N}\sum_{i=1}^{N}a(i,m)$ (A.17)
and $a(i,m)=\|Y_{i}^{(m+1)}-Y_{n(i,m)}^{(m+1)}\|/\|Y_{i}^{(m)}-Y_{n(i,m)}^{(m)}\|$. The condition $E_{1}(m)\approx 1$ indicates manifold saturation at dimension $m$.
The $E_{2}$ statistic measures future predictability:
$E_{2}(m)=\frac{E^{*}(m+1)/E^{*}(m)}{E(m+1)/E(m)}$ (A.18)
where $E^{*}(m)$ measures divergence of futures from nearby states. Under local Gaussianity, $E_{2}$ quantifies the signal-to-noise ratio: $\text{SNR}\approx(1-E_{2})/(E_{2}\cdot\tau)$. At this stage, classical Takens-style reconstruction fails; the remainder of the proof replaces pathwise injectivity with injectivity in distribution.
A.5 The Proof Architecture
This section establishes the theoretical basis for stochastic embedding. This requires proving that the correlation manifold supports well-defined, uniquely reconstructible dynamics. The proof proceeds through five essential steps.
A.5.1 Step 1: Hörmander Hypoelliptic Regularity
Definition A.3 (Hörmander’s Condition).
The Lie algebra generated by drift and diffusion vector fields spans the tangent space:
$\text{Lie}(\mu,\sigma_{1},\ldots,\sigma_{r})=\text{span}\{\mu,\sigma_{i},[\mu,\sigma_{i}],[\sigma_{i},\sigma_{j}],[[\mu,\sigma_{i}],\sigma_{j}],\ldots\}=T_{x}\mathbb{R}^{n}\quad\forall x\in\mathbb{R}^{n}$ (A.19)
Hörmander’s condition [62] on the SDE coefficients complements the genericity requirement on the observation function $h$: the former ensures smooth transition densities, while the latter avoids degenerate embeddings. Hörmander’s condition ensures the noise structure is sufficiently rich that the stochastic flow explores all directions. This guarantees:
  • Smooth transition densities: $p_{t}(x,y)\in C^{\infty}$ for all $t>0$
  • Strict positivity: $p_{t}(x,y)>0$ for all $x,y\in\mathbb{R}^{n}$ and $t>0$ (by the Stroock–Varadhan support theorem [34] and the Chow–Rashevskii theorem)
  • Non-degeneracy despite rank deficiency: Even if $\sigma$ is not full-rank, iterated Lie brackets ensure noise propagates to all directions through the dynamics
This step ensures that the stochastic dynamics generate smooth probability densities, a prerequisite for establishing uniqueness of finite-dimensional laws under embedding.
A.5.2 Step 2: Malliavin Non-Degeneracy
Malliavin calculus provides the analytical framework [35] for characterising smoothness of stochastic flows (see also [34] for the foundational support theorems). The Malliavin covariance matrix $\gamma_{t}$ measures sensitivity of $X_{t}$ to perturbations in the driving Brownian motion:
$\gamma_{t}=\int_{0}^{t}(D_{s}X_{t})(D_{s}X_{t})^{\top}ds$ (A.20)
where $D_{s}X_{t}$ denotes the Malliavin derivative of $X_{t}$ with respect to the Brownian path at time $s$.
Lemma A.1 (Malliavin Non-Degeneracy).
Under Hörmander’s condition and standard regularity assumptions on $\mu$ and $\sigma$ (Lipschitz continuity, linear growth bound), $\gamma_{t}$ is almost surely invertible for $t>0$.
This non-degeneracy ensures that the law of $X_{t}$ has smooth density with respect to Lebesgue measure, and that, combined with the genericity of $h$, the delay embedding map has full rank in a measure-theoretic sense.
A.5.3 Step 3: Law-Separation and Measure-Theoretic Injectivity
The objective in the stochastic setting is not trajectory reconstruction, but distinguishing system states via their induced probability measures. The upgraded proof (Supplemental Materials, Theorems 7.5–7.7) establishes this through three linked results:
(a) Transition density separation (Theorem 7.5). Under Hörmander’s condition, distinct initial conditions $x\neq x^{\prime}$ produce distinct transition densities: $p_{\tau}(x,\cdot)\neq p_{\tau}(x^{\prime},\cdot)$ for every $\tau>0$. The proof uses semigroup bisection: from $p_{\tau}(x,\cdot)=p_{\tau}(x^{\prime},\cdot)$, injectivity of the semigroup operator $P_{\tau/2}$ (guaranteed by strict positivity of the kernel under Hörmander) yields $p_{\tau/2^{k}}(x,\cdot)=p_{\tau/2^{k}}(x^{\prime},\cdot)$ for all $k$. The Varadhan–Léandre short-time asymptotic then forces $x=x^{\prime}$.
(b) Law-separation via observed delay vectors (Theorem 7.6). For a prevalent observation function $h\in C^{r}$ ($r\geq 2$) and $m\geq\lceil 2D\rceil+1$ (where $D$ is the exact dimension of $\mu_{\infty}$), the law-embedding map $\Lambda_{m}^{h}:x\mapsto\mathrm{Law}(h(X_{0}^{x}),h(X_{\tau}^{x}),\ldots,h(X_{(m-1)\tau}^{x}))$ is $\mu_{\infty}$-a.e. injective. The proof reduces the infinite-dimensional collision condition (equality of densities) to a finite-dimensional evaluation, applies the parametric transversality theorem with an explicit evaluation-map surjectivity lemma, and bounds the collision set using a direct Frostman covering argument (exploiting the exact-dimensional Frostman upper bound $\mu_{\infty}(B(x,r))\leq C_{F}r^{D}$).
(c) Measure-zero geometric collisions (Theorem 7.7). The geometric collision set $\{(x,x^{\prime}):\Phi_{m}(x)=\Phi_{m}(x^{\prime})\}$ is a subset of the law-collision set, hence also has $(\mu_{\infty}\times\mu_{\infty})$-measure zero.
Note that pathwise injectivity (c) is a corollary of law-separation (b), not the primary claim. The law-embedding perspective is operationally relevant because $k$-NN estimators use the empirical distribution of delay vectors, not individual realisations. Bad delays, which require careful treatment in deterministic Takens theory, do not arise: the density separation in (a) holds for every $\tau>0$ under Hörmander’s condition.
A.5.4 Step 4: $E_{1}$ Dimension Sufficiency via Correlation Dimension
For stochastic systems, correlation dimension: not topological dimension: governs the minimal embedding dimension required for consistent geometric inference.
The correlation dimension $D_{2}$ characterises how the invariant measure [36] $\mu_{\infty}$ scales locally:
$C(\varepsilon)=\iint\mathbf{1}_{\{\|Y-Y^{\prime}\|<\varepsilon\}}d\mu_{\infty}(Y)d\mu_{\infty}(Y^{\prime})\sim\varepsilon^{D_{2}}\text{ as }\varepsilon\to 0$ (A.21)
Lemma A.2 ($E_{1}$ Convergence).
$E_{1}(m)\to 1$ precisely when $m$ exceeds $D_{2}$ (see Supplemental Materials for proof).
This follows from the scaling of $k$-NN distances: $\varepsilon_{k}\sim(k/N)^{1/D_{2}}$. When $m<D_{2}$, false neighbors exist (nearby in $\mathbb{R}^{m}$ but distant on the manifold), causing $E_{1}>1$. When $m\geq D_{2}+1$, all neighbors are true neighbors, and $E_{1}\approx 1$.
Correlation dimension remains finite for unbounded stochastic processes. The Ornstein-Uhlenbeck process [4] has topological support $\mathbb{R}$ (unbounded) but $D_{2}=1$. The finite-time marginals of Brownian motion in $\mathbb{R}^{n}$ similarly yield $D_{2}=n$. Consequently, $E_{1}$ succeeds for stochastic systems where topological arguments fail.
A.5.5 Step 5: Finite-Dimensional Law Uniqueness
Given measure-theoretic injectivity (Step 3), the conditional expectations defining the embedded drift and diffusion tensors are single-valued functions on the correlation manifold. The recovery of drift and diffusion coefficients from time-series data via these conditional moments was pioneered by Friedrich and Peinke [37]. For the embedded process, the drift $\tilde{\mu}$ and diffusion tensor $\tilde{\Sigma}$ are defined by:
$\tilde{\mu}(Y)$ $=\lim_{\Delta t\to 0}\frac{1}{\Delta t}\mathbb{E}[Y_{t+\Delta t}-Y_{t}\mid Y_{t}=Y]$ (A.22)
$\tilde{\Sigma}(Y)$ $=\lim_{\Delta t\to 0}\frac{1}{\Delta t}\mathbb{E}[(Y_{t+\Delta t}-Y_{t})(Y_{t+\Delta t}-Y_{t})^{\top}\mid Y_{t}=Y]$ (A.23)
These local conditional moments are the quantities that $k$-NN estimators compute. Because the framework relies on $k$-nearest neighbour graphs rather than trajectory prediction, measure-theoretic injectivity is sufficient for all subsequent divergence calculations. The embedded drift $\tilde{\mu}$ and diffusion $\tilde{\Sigma}$ are thus uniquely recoverable from the observed time series.
A.6 The Stochastic Embedding Sufficiency Theorem
Assumption A.1 (Sampling and Regularity for Nonparametric Reconstruction).
The following conditions are assumed for the estimation result (Part II of Theorem A.3):
  1. (S1)
    The sampled delay vectors $(Y_{k\Delta t}^{(m)})_{k\geq 0}$ are stationary and geometrically ergodic (exponential decay of correlations) under $\mu_{\infty}$.
  2. (S2)
    The conditional drift $\tilde{\mu}$ and diffusion tensor $\tilde{\Sigma}$ on $\mathcal{M}_{D_{2}}$ are $\beta$-Hölder continuous ($\beta>0$).
  3. (S3)
    The delay-vector law $\nu_{m}$ admits a density (with respect to the intrinsic volume on $\mathcal{M}_{D_{2}}$) bounded away from zero on compact subsets.
Theorem A.3 (Stochastic Embedding Sufficiency).
Let $X_{t}$ solve the SDE (A.14) with $C^{\infty}_{b}$ coefficients $\mu$ and $\sigma$ satisfying the uniform Hörmander condition (bracket-generating to depth $k_{0}$), and admitting an ergodic invariant measure $\mu_{\infty}$ that is exact-dimensional with dimension $D$ and satisfies a Frostman upper bound $\mu_{\infty}(B(x,r))\leq C_{F}r^{D}$. Let $h$ be a prevalent $C^{r}$ ($r\geq 2$) observation function (in the sense of Definition A.2). Then the embedding dimension
$m^{*}=\max\{\lceil 2D\rceil+1,m_{E_{1}}\}$ (A.24)
guarantees the following:
Part I, Identifiability.
  1. (i)
    Law-separation ($\mu_{\infty}$-a.e. injectivity): for $\mu_{\infty}$-a.e. $x$, if $x\neq x^{\prime}$ then $\mathrm{Law}(Y_{0}^{(m^{*})}|X_{0}=x)\neq\mathrm{Law}(Y_{0}^{(m^{*})}|X_{0}=x^{\prime})$. The collision set $\{(x,x^{\prime}):x\neq x^{\prime},\;\Lambda_{m^{*}}^{h}(x)=\Lambda_{m^{*}}^{h}(x^{\prime})\}$ has $(\mu_{\infty}\times\mu_{\infty})$-measure zero.
  2. (ii)
    Single-valued drift-diffusion tensors $\tilde{\mu}(Y),\tilde{\Sigma}(Y)$ on $\mathcal{M}_{D_{2}}$.
Part II, Estimation Consistency. Under Assumption A.1:
  1. (iii)
    $k$-NN convergence [38] at rate $O_{P}((k/N)^{\beta/m^{*}})+O(\Delta t)$ as $N\to\infty$, $k\to\infty$, $k/N\to 0$.
Remark A.1 (Correlation-Dimension Version).
When $\mu_{\infty}$ is absolutely continuous on its support (the generic case for Hörmander SDEs on $\mathbb{R}^{n}$), the exact dimension $D$ equals the correlation dimension $D_{2}$, and the threshold becomes $m^{*}\geq\lceil 2D_{2}\rceil+1$. This mirrors Takens’ $2n+1$, replacing state dimension with correlation dimension.
Remark A.2 (Logical Independence of Parts I and II).
Part I is a measure-theoretic result requiring only Hörmander regularity, exact-dimensionality, and the dimension bound $m^{*}\geq\lceil 2D\rceil+1$. Part II additionally requires Assumption A.1 (geometric ergodicity, Hölder smoothness, and non-degenerate sampling density); the rate $\beta/m^{*}$ is the standard nonparametric rate in $m^{*}$-dimensional spaces [38]. The two parts can be verified independently.
The complete proof of Part I is given in the Supplemental Materials (Theorems 7.5–7.7), where transition density separation, law-separation via parametric transversality, and the Frostman covering argument are established in full. The proof architecture: semigroup bisection, evaluation-map surjectivity lemma, and direct Hausdorff pre-measure bound: is summarised in Step 3 above.
A.7 The Unified Framework
The framework unifies deterministic and stochastic dynamics as equivalent descriptions of the same Markov process on the correlation manifold (with the discrete chain arising from time-discretisation):
$\text{Markov chain }p(Y_{t+1}|Y_{t})\longleftrightarrow\text{SDE }dY=\mu dt+\sigma dW\longleftrightarrow\text{Generator }\mathcal{L}=\mu\cdot\nabla+\tfrac{1}{2}\text{tr}(\Sigma\nabla^{2})$ (A.25)
$E_{2}$ positions systems along the spectrum: $E_{2}\to 0$ recovers deterministic Takens ($\mu$ dominates); $E_{2}\to 1$ yields pure diffusion ($\Sigma$ dominates). The correlation manifold identified by $E_{1}$ persists throughout this spectrum.
Appendix B Relativistic Transformation Properties of Nelson’s Stochastic Mechanics
Scope and Logical Position
This appendix derives the relativistic transformation law for the Nelson diffusion coefficient from special relativity and the structure of the Klein–Gordon conserved current. The derivation proceeds in two stages: (i) the kinematic result $\sigma_{\mathrm{rel}}=\sigma_{0}/\sqrt{\gamma}$ is obtained from time dilation alone, via the Dambis–Dubins–Schwarz reparametrisation theorem applied to the proper-time Wiener process; (ii) the full result $\sigma_{\mathrm{KG}}=\sigma_{0}/\gamma$ is obtained by additionally requiring self-consistency of the Fokker–Planck equation with the Klein–Gordon conserved density $\rho_{\mathrm{KG}}=\gamma|\phi|^{2}$. The two factors of $1/\sqrt{\gamma}$ are structurally independent: the first is a theorem of stochastic calculus, the second is forced by relativistic current conservation.
Relationship to the canonical axioms. This derivation provides the theoretical underpinning for Tests C6 (electromagnetic determinism) and C7 (relativistic suppression) of the superspace diffusion framework. Nelson’s SDE (Appendix D of SI-1) postulates $\mathrm{d}X^{i}=b^{i}\,\mathrm{d}t+\sqrt{\hbar/m}\,\mathrm{d}W^{i}$ in the non-relativistic limit where proper time $\tau$ coincides with coordinate time $t$; the present appendix reformulates this in proper time and establishes the unique Lorentz-covariant extension. No additional axioms beyond A1–A4 are introduced; all inputs are either established results of special relativity, the Klein–Gordon equation, or theorems of stochastic calculus. The four canonical axioms are stated in SI-2, §2.1.
B.1 Definitions and Setup
Definition B.1 (Rest-Frame Nelson SDE).
In the co-moving rest frame $\mathcal{S}_{0}$ of a particle of mass $m$, Nelson’s stochastic mechanics [6, 24] postulates the spatial SDE
$\mathrm{d}X^{i}_{0}=b^{i}_{0}(X_{0},\tau)\,\mathrm{d}\tau+\sigma_{0}\,\mathrm{d}W^{i}(\tau),\qquad i=1,2,3,$ (B.26)
where:
  • $\tau$ is proper time along the particle worldline;
  • $b^{i}_{0}$ is the forward drift velocity (determined self-consistently from the quantum potential; cf. Appendix D of SI-1);
  • $\sigma_{0}=\sqrt{\hbar/m}$ is the rest-frame diffusion coefficient;
  • $W^{i}(\tau)$, $i=1,2,3$, are three independent standard Wiener processes satisfying
    $\mathbb{E}[\mathrm{d}W^{i}(\tau)]=0,\qquad\mathbb{E}[\mathrm{d}W^{i}(\tau)\,\mathrm{d}W^{j}(\tau)]=\delta^{ij}\,\mathrm{d}\tau.$ (B.27)
Definition B.2 (Diffusion Coefficient Dimensions).
For a spatial variable $X^{i}$ with dimensions $[X]=L$ evolving in proper time with $\tau]=T$, the diffusion coefficient has dimensions
$\sigma_{0}]=L\cdot T^{-1/2},$ (B.28)
since $\mathrm{d}W]=T^{1/2}$ and $\sigma_{0}\,\mathrm{d}W$ must have dimensions $L$. Explicitly, $\sqrt{\hbar/m}]=\sqrt{ML^{2}T^{-1}/M}=L\,T^{-1/2}$.
Definition B.3 (Worldline Proper-Time Relation).
Along the particle worldline, proper time and coordinate time are related by the deterministic equation
$\mathrm{d}\tau=\frac{\mathrm{d}t}{\gamma},\qquad\gamma=\frac{1}{\sqrt{1-v^{2}/c^{2}}}\,$, (B.29)
where $v$ is the macroscopic (drift) velocity of the particle in the laboratory frame.
Remark B.1 (Why $\mathrm{d}\tau/\mathrm{d}t$ Is Deterministic).
A potential concern is that $\gamma$ should itself fluctuate, since the particle position $X$ is stochastic. However, in Itô calculus the stochastic displacements $\sigma_{0}\,\mathrm{d}W^{i}$ have magnitude $O(\sqrt{\mathrm{d}\tau})$, which contributes to the four-velocity only at order $O(\mathrm{d}\tau^{-1/2})$, handled by the quadratic variation formalism, not by modifying $\gamma$. Along the worldline, $\gamma$ is determined by the macroscopic four-velocity $U^{\mu}=\mathrm{d}x^{\mu}/\mathrm{d}\tau$, which is the drift component of the SDE. The relation $\mathrm{d}\tau=\mathrm{d}t/\gamma$ therefore holds with $\gamma$ evaluated on the drift [39, 40, 41].
For a particle of definite momentum $p=m\gamma v$, $\gamma$ is constant along the worldline. For wavepackets with momentum spread $\Delta p$, $\gamma$ varies over the ensemble but is deterministic for each trajectory at leading order. Stochastic corrections to $\gamma$ enter at $O(\sigma_{0}^{2}/c^{2})=O(\hbar/(mc^{2}))$, negligible in the semiclassical regime relevant to Test C7.
Conventions. Throughout, Greek indices $\mu,\nu,\ldots$ run over $\{0,1,2,3\}$; Latin indices $i,j,\ldots$ run over $\{1,2,3\}$. Summation over repeated indices is understood. We work in SI units except where natural units ($\hbar=c=1$) are adopted for brevity.
B.2 The Foundational Identity: Time Reparametrisation of the Wiener Process
The entire relativistic extension rests on a single result from stochastic calculus.
Lemma B.1 (Dambis–Dubins–Schwarz Reparametrisation).
Let $W(\tau)$ be a standard one-dimensional Wiener process parametrised by $\tau$, and let $t(\tau)$ be a strictly increasing, absolutely continuous function with $t(0)=0$. Let $\tau(t)$ denote the functional inverse of $t(\tau)$, and set
$\widetilde{W}(t)\;\equiv\;W\bigl(\tau(t)\bigr).$ (B.30)
Then $\widetilde{W}(t)$ is a continuous martingale with quadratic variation
$\widetilde{W}]_{t}=\tau(t),$ (B.31)
and the differential relation is
$\mathrm{d}\widetilde{W}(t)=\sqrt{\frac{\mathrm{d}\tau}{\mathrm{d}t}}\;\mathrm{d}B(t),$ (B.32)
where $B(t)$ is a standard Wiener process in the variable $t$.
Proof.
By the Dambis–Dubins–Schwarz theorem [27, 42], any continuous local martingale $M$ with $\langle M\rangle_{\infty}=\infty$ can be written as $M_{t}=\beta_{\langle M\rangle_{t}}$ for some Brownian motion $\beta$. Applied to $\widetilde{W}$:
$\langle\widetilde{W}\rangle_{t}=\int_{0}^{t}\frac{\mathrm{d}\tau}{\mathrm{d}s}\,\mathrm{d}s=\tau(t).$
Setting $\beta=B$ and differentiating gives
$\mathrm{d}\widetilde{W}=\mathrm{d}B_{\tau(t)}=\sqrt{\frac{\mathrm{d}\tau}{\mathrm{d}t}}\;\mathrm{d}B(t)$
by Lévy’s characterisation of Brownian motion (the process $B$ has the correct quadratic variation $\widetilde{W}]_{t}=\tau(t)$). ∎
Corollary B.2 (Brownian Scaling under Constant Lorentz Factor).
When $\gamma$ is constant along the worldline (i.e., $\mathrm{d}\tau/\mathrm{d}t=1/\gamma$ with $\gamma$ independent of $t$), the DDS reparametrisation (Lemma B.1) reduces to the deterministic time-change identity:
$\mathrm{d}W^{i}(\tau)=\frac{1}{\sqrt{\gamma}}\,\mathrm{d}B^{i}(t),\qquad[\widetilde{W}^{i}]_{t}=\frac{t}{\gamma},$ (B.33)
where $B^{i}(t)$ are independent standard Wiener processes in coordinate time. The quadratic variation is linear in $t$ (not merely absolutely continuous), so the reparametrised process is a scaled Brownian motion , not merely a continuous martingale.
Proof.
For constant $\gamma$, $\tau(t)=t/\gamma$ is linear, so $\mathrm{d}\tau/\mathrm{d}t=1/\gamma$ is constant. Lemma B.1 gives $\mathrm{d}\widetilde{W}=(1/\gamma)^{1/2}\,\mathrm{d}B$, and the quadratic variation $\widetilde{W}]_{t}=\tau(t)=t/\gamma$ is linear in $t$. By Lévy’s characterisation, a continuous martingale with linear quadratic variation is a scaled Brownian motion. ∎
Remark B.2 (Itô vs. Stratonovich Convention).
This appendix works exclusively in the Itô convention, consistent with Nelson’s original formulation [6]. The Dambis–Dubins–Schwarz theorem (Lemma B.1) and its specialisation (Corollary B.2) are Itô results. Because the noise in the rest-frame SDE (B.26) is additive ($\sigma_{0}$ is independent of $X$), the Stratonovich correction $\tfrac{1}{2}\sigma_{0}\,\partial\sigma_{0}/\partial X^{i}$ vanishes identically. The Itô and Stratonovich forms of the SDE therefore coincide, and no ambiguity arises from the choice of calculus convention. For multiplicative noise (state-dependent diffusion), a Stratonovich formulation would be required for geometric covariance; the additive structure here avoids this complication.
Remark B.3 (Central Identity).
Equation (B.32) is the foundational identity of this appendix: reparametrising the time of a Wiener process rescales its increments by the square root of the time Jacobian. This is purely a theorem of stochastic calculus [27]. All subsequent results are consequences of this identity applied to the specific Jacobian $\mathrm{d}\tau/\mathrm{d}t=1/\gamma$.
B.3 Nelson’s SDE in Proper Time
Theorem B.3 (Uniqueness within the Additive Isotropic Class).
The SDE (B.26) is the unique additive-noise Itô diffusion in proper time that is compatible with the following requirements:
  1. (R1)
    Isotropy: The diffusion is rotationally invariant in the rest frame: the noise covariance is proportional to $\delta^{ij}$.
  2. (R2)
    Markov property: The process is Markovian with respect to proper time.
  3. (R3)
    Non-relativistic limit: In the limit $v/c\to 0$ (where $\tau\to t$), the SDE reduces to Nelson’s original equation.
  4. (R4)
    Dimensional consistency: The diffusion coefficient has dimensions $\text{length}]\cdot[\text{time}]^{-1/2}$, constructible from $\hbar$ and $m$ alone in the rest frame.
This uniqueness holds within the class of additive-noise, isotropic, Markov diffusions parametrised by proper time, with diffusion coefficient depending only on $\hbar$ and $m$. Relaxing any of conditions (R1)–(R4) admits alternative formulations.
Proof.
Requirements (R1) and (R4) fix the noise covariance to the form $\sigma_{0}^{2}\,\delta^{ij}\,\mathrm{d}\tau$ with $\sigma_{0}=\sqrt{\hbar/m}$, since $\hbar/m$ is the unique combination of $\hbar$ and $m$ with dimensions $L^{2}/T$.
Requirement (R2) determines the SDE structure: an Itô diffusion in proper time.
Requirement (R3) fixes $\tau\to t$ in the non-relativistic limit, recovering Nelson’s original equation.
The SDE is Lorentz-scalar by construction: $\tau$ is a Lorentz scalar, $\mathrm{d}X^{i}_{0}$ are rest-frame spatial components, and $\sigma_{0}$ is a rest-frame constant. ∎
B.4 The Kinematic Result: $\sigma_{\mathrm{rel}}=\sigma_{0}/\sqrt{\gamma}$
We now boost from the rest frame $\mathcal{S}_{0}$ to a laboratory frame $\mathcal{S}$ in which the particle moves with velocity $v$ along the $x^{1}$-axis. In what follows we restrict attention to inertial Lorentz boosts with constant Lorentz factor $\gamma$ (i.e. fixed relative velocity between frames), so the time dilation $\mathrm{d}\tau=\mathrm{d}t/\gamma$ is deterministic. Extension to accelerating worldlines (state-dependent $\gamma$) would require a general DDS time change (Lemma B.1) rather than the constant-$\gamma$ specialisation (Corollary B.2).
Theorem B.4 (Transformation of Wiener Increments under Lorentz Boost).
Under the worldline time reparametrisation $\mathrm{d}\tau=\mathrm{d}t/\gamma$, the proper-time Wiener increments $\mathrm{d}W^{i}(\tau)$ are related to coordinate-time Wiener processes $B^{i}(t)$ by:
$\boxed{\mathrm{d}W^{i}(\tau)=\frac{1}{\sqrt{\gamma}}\,\mathrm{d}B^{i}(t),}$ (B.34)
where $B^{i}(t)$ are independent standard Wiener processes satisfying $\mathbb{E}[\mathrm{d}B^{i}(t)\,\mathrm{d}B^{j}(t)]=\delta^{ij}\,\mathrm{d}t$.
The covariance structure in laboratory time is:
$\mathbb{E}\!\left[\mathrm{d}W^{i}(\tau)\,\mathrm{d}W^{j}(\tau)\right]=\delta^{ij}\,\mathrm{d}\tau=\frac{\delta^{ij}}{\gamma}\,\mathrm{d}t.$ (B.35)
Proof.
Direct application of Lemma B.1 with the time change $\tau(t)=t/\gamma$ (constant $\gamma$). Then $\mathrm{d}\tau/\mathrm{d}t=1/\gamma$, and by (B.32):
$\mathrm{d}W^{i}(\tau)=\sqrt{\frac{\mathrm{d}\tau}{\mathrm{d}t}}\,\mathrm{d}B^{i}(t)=\frac{1}{\sqrt{\gamma}}\,\mathrm{d}B^{i}(t).$
The covariance follows from $\mathbb{E}[\mathrm{d}B^{i}\,\mathrm{d}B^{j}]=\delta^{ij}\,\mathrm{d}t$. ∎
Theorem B.5 (Coordinate-Time SDE and Relativistic Diffusion Coefficient).
In the laboratory frame $\mathcal{S}$, the Nelson SDE for a relativistic particle, expressed in coordinate time $t$, takes the form
$\boxed{\mathrm{d}X^{i}=\mu^{i}(X,t)\,\mathrm{d}t+\frac{\sigma_{0}}{\sqrt{\gamma}}\,\mathrm{d}B^{i}(t),\qquad\sigma_{\mathrm{rel}}=\frac{\sigma_{0}}{\sqrt{\gamma}}=\sqrt{\frac{\hbar}{m\gamma}}\,,}$ (B.36)
where $\mu^{i}$ is the coordinate-time drift (incorporating the group velocity and the boosted quantum potential) and $B^{i}(t)$ are standard Wiener processes in coordinate time.
Proof.
Substituting the time reparametrisation $\mathrm{d}\tau=\mathrm{d}t/\gamma$ and the Wiener transformation (B.34) into the rest-frame SDE (B.26):
$\mathrm{d}X^{i}_{0}$ $=b^{i}_{0}\,\mathrm{d}\tau+\sigma_{0}\,\mathrm{d}W^{i}(\tau)$
$=b^{i}_{0}\cdot\frac{\mathrm{d}t}{\gamma}+\sigma_{0}\cdot\frac{1}{\sqrt{\gamma}}\,\mathrm{d}B^{i}(t)$
$=\frac{b^{i}_{0}}{\gamma}\,\mathrm{d}t+\frac{\sigma_{0}}{\sqrt{\gamma}}\,\mathrm{d}B^{i}(t).$ (B.37)
For the transverse directions ($i=2,3$), the laboratory coordinates coincide with the rest-frame coordinates: $X^{a}=X^{a}_{0}$, so $\mathrm{d}X^{a}$ takes the form (B.37) directly.
For the longitudinal direction ($i=1$), the Lorentz boost gives $X^{1}=\gamma(X^{1}_{0}+v\tau)$, so:
$\mathrm{d}X^{1}=\gamma\,\mathrm{d}X^{1}_{0}+\gamma v\,\mathrm{d}\tau=\gamma\!\left(\frac{b^{1}_{0}}{\gamma}\,\mathrm{d}t+\frac{\sigma_{0}}{\sqrt{\gamma}}\,\mathrm{d}B^{1}\right)+v\,\mathrm{d}t.$ (B.38)
The diffusion coefficient is determined by the quadratic variation:
$\bigl[\mathrm{d}X^{1},\mathrm{d}X^{1}\bigr]$ $=\gamma^{2}\cdot\sigma_{0}^{2}\cdot[\mathrm{d}W^{1},\mathrm{d}W^{1}]=\gamma^{2}\cdot\sigma_{0}^{2}\cdot\mathrm{d}\tau=\gamma^{2}\cdot\frac{\sigma_{0}^{2}}{\gamma}\,\mathrm{d}t=\gamma\,\sigma_{0}^{2}\,\mathrm{d}t.$ (B.39)
This gives a longitudinal coordinate-space diffusion coefficient $\sigma_{\parallel}^{\mathrm{(coord)}}=\sigma_{0}\sqrt{\gamma}$. This apparent enhancement is a coordinate artefact arising from the Lorentz contraction of the spatial coordinate; it does not represent enhanced physical diffusion.
The operationally meaningful diffusion coefficient is the invariant (proper-time, rest-frame) quadratic variation:
$\bigl[\mathrm{d}X^{i}_{0},\mathrm{d}X^{i}_{0}\bigr]=\sigma_{0}^{2}\,\mathrm{d}\tau=\frac{\sigma_{0}^{2}}{\gamma}\,\mathrm{d}t,\qquad\text{(no sum on $i$)}.$ (B.40)
This is a Lorentz scalar and gives $\sigma_{\mathrm{rel}}=\sigma_{0}/\sqrt{\gamma}$ unambiguously. ∎
Definition B.4 (Operational Diffusion Coefficient).
The operational diffusion coefficient $\sigma_{\mathrm{rel}}$ is defined as the quantity appearing in the coordinate-time SDE
$\mathrm{d}X^{i}=\mu^{i}\,\mathrm{d}t+\sigma_{\mathrm{rel}}\,\mathrm{d}B^{i}(t)$ (B.41)
when the SDE is expressed in rest-frame spatial coordinates and laboratory time. This is the quantity directly reconstructed by the Stochastic Embedding pipeline from the time series $X(t)$.
Remark B.4 (Why Rest-Frame Spatial Coordinates).
The choice of rest-frame spatial coordinates is not arbitrary. It is the unique frame in which: (a) the diffusion is isotropic ($D^{ij}\propto\delta^{ij}$); (b) the probability density equals $|\psi|^{2}$ (the Born rule); (c) the Fokker–Planck equation has the standard Nelson form. In any other spatial coordinates, the diffusion tensor acquires frame-dependent distortions (cf. Remark B.5).
B.5 The Laboratory-Frame Fokker–Planck Equation
Theorem B.6 (Fokker–Planck Equation in Coordinate Time).
The probability density $\rho_{0}(x,t)=|\psi(x,t)|^{2}$ (the Born density; Formulation A of §B.9.3) for the coordinate-time SDE (B.36) satisfies the Fokker–Planck equation
$\frac{\partial\rho_{0}}{\partial t}=-\nabla\cdot(\mu\,\rho_{0})+\frac{\sigma_{0}^{2}}{2\gamma}\,\nabla^{2}\rho_{0},$ (B.42)
with diffusion coefficient
$D_{\mathrm{lab}}=\frac{\sigma_{\mathrm{rel}}^{2}}{2}=\frac{\sigma_{0}^{2}}{2\gamma}=\frac{\hbar}{2m\gamma}\,.$ (B.43)
Proof.
This is the standard Fokker–Planck equation associated with the Itô SDE (B.36) [43, 44]. For a general SDE $\mathrm{d}X=\mu\,\mathrm{d}t+\sigma\,\mathrm{d}B$, the Fokker–Planck equation is $\partial_{t}\rho=-\nabla\cdot(\mu\rho)+(\sigma^{2}/2)\nabla^{2}\rho$. Substituting $\sigma=\sigma_{0}/\sqrt{\gamma}$ gives (B.42). ∎
Proposition B.7 (Consistency with the Non-Relativistic Limit).
In the limit $v\to 0$ (hence $\gamma\to 1$), the Fokker–Planck equation (B.42) reduces to the standard Nelson Fokker–Planck equation with $D=\hbar/(2m)$, recovering the Schrödinger equation via the Madelung decomposition.
Proof.
Setting $\gamma=1$ in (B.43) gives $D_{\mathrm{lab}}=\hbar/(2m)$, the non-relativistic Nelson diffusion coefficient. The Fokker–Planck equation becomes $\partial_{t}\rho_{0}=-\nabla\cdot(\mu\rho_{0})+[\hbar/(2m)]\nabla^{2}\rho_{0}$, which is precisely the equation whose Madelung decomposition [45] yields the Schrödinger equation (Appendix D of SI-1). ∎
B.6 Transverse Components and Isotropy
Proposition B.8 (Isotropy in the Rest Frame; Anisotropy in the Lab Frame).
For a particle boosted along $x^{1}$:
  1. (a)
    In the rest-frame spatial coordinates, the diffusion is isotropic with coefficient $\sigma_{0}/\sqrt{\gamma}$ per unit coordinate time in all three directions.
  2. (b)
    In the laboratory spatial coordinates, the diffusion tensor is anisotropic:
    $D^{ij}_{\mathrm{lab}}=\frac{\sigma_{0}^{2}}{2\gamma}\,\mathrm{diag}\bigl(\gamma^{2},\;1,\;1\bigr).$ (B.44)
  3. (c)
    In the one-dimensional projection onto any fixed direction (as in Test C7), the reconstructed diffusion coefficient is $\sigma_{\mathrm{rel}}=\sigma_{0}/\sqrt{\gamma}$.
Proof.
(a) follows directly from the SDE (B.37), which gives $\sigma_{0}/\sqrt{\gamma}$ for all three rest-frame components.
(b) The Lorentz boost maps $\mathrm{d}X^{1}_{0}\mapsto\gamma\,\mathrm{d}X^{1}_{0}$ (plus drift terms), while $\mathrm{d}X^{a}=\mathrm{d}X^{a}_{0}$ for $a=2,3$. The quadratic variations are:
$\mathrm{d}X^{1},\mathrm{d}X^{1}]=\gamma^{2}\cdot\frac{\sigma_{0}^{2}}{\gamma}\,\mathrm{d}t=\gamma\sigma_{0}^{2}\,\mathrm{d}t,\qquad[\mathrm{d}X^{a},\mathrm{d}X^{a}]=\frac{\sigma_{0}^{2}}{\gamma}\,\mathrm{d}t.$
This gives the diagonal entries $(D_{\parallel},D_{\perp},D_{\perp})=(\gamma\sigma_{0}^{2}/2,\;\sigma_{0}^{2}/(2\gamma),\;\sigma_{0}^{2}/(2\gamma))$, which is (B.44).
(c) The pipeline reconstructs the diffusion coefficient from the rest-frame SDE (B.37), which gives $\sigma_{\mathrm{rel}}=\sigma_{0}/\sqrt{\gamma}$. The anisotropy (B.44) is a coordinate artefact that does not affect the reconstructed value. ∎
Remark B.5 (Physical Origin of the Anisotropy).
The anisotropy (B.44) is entirely analogous to the anisotropy of the electromagnetic field under a Lorentz boost: a Coulomb field that is isotropic at rest appears anisotropic in a boosted frame. The underlying physics (isotropic rest-frame diffusion) is Lorentz-invariant; the apparent anisotropy arises from expressing this invariant content in non-rest-frame coordinates.
B.7 Photon Limit and Connection to Test C6
Proposition B.9 (Deterministic Limit as $\gamma\to\infty$).
In the ultrarelativistic limit $v\to c$ (equivalently $\gamma\to\infty$), the diffusion coefficient vanishes:
$\sigma_{\mathrm{rel}}=\frac{\sigma_{0}}{\sqrt{\gamma}}\to 0\qquad\text{as}\quad\gamma\to\infty.$ (B.45)
This is consistent with Test C6 (electromagnetism), where the pipeline recovers $\hat{\sigma}=3.5\times 10^{-14}\approx 0$ from Maxwell’s equations: photons propagate deterministically because their effective $\gamma$ is infinite.
Proof.
$\sigma_{0}/\sqrt{\gamma}\to 0$ as $\gamma\to\infty$, since $\sigma_{0}$ is finite. Physically: in the photon limit, proper time ceases to advance ($\mathrm{d}\tau=\mathrm{d}t/\gamma\to 0$), so no stochastic increments accumulate per unit coordinate time. (The full result $\sigma_{\mathrm{KG}}=\sigma_{0}/\gamma$, derived in §B.9, vanishes even faster and strengthens this conclusion.) ∎
B.8 Summary of the Kinematic Result
Quantity Rest frame $\mathcal{S}_{0}$ Lab frame $\mathcal{S}$
Time parameter $\tau$ (proper time) $t=\gamma\tau$ (coordinate time)
Wiener increment $\mathrm{d}W^{i}(\tau)$ $\gamma^{-1/2}\,\mathrm{d}B^{i}(t)$
Diffusion coefficient $\bm{\sigma_{0}=\sqrt{\hbar/m}}$ $\bm{\sigma_{0}/\sqrt{\gamma}}$
Fokker–Planck diffusion $D_{0}=\hbar/(2m)$ $D_{\mathrm{lab}}=\hbar/(2m\gamma)$
Quadratic variation per $\mathrm{d}t$ : $\sigma_{0}^{2}/\gamma$
Limit $\gamma\to 1$ $\sigma_{0}$ $\sigma_{0}$ (non-relativistic)
Limit $\gamma\to\infty$ : 0 (deterministic)
Table 5: Transformation properties of stochastic mechanical quantities under a Lorentz boost. The operational diffusion coefficient (bold row) transforms as $\sigma_{0}/\sqrt{\gamma}$. This is the quantity directly reconstructed by the Stochastic Embedding pipeline from coordinate-time trajectories.
B.9 The Klein–Gordon Density and the Full $1/\gamma$ Suppression
The result $\sigma_{\mathrm{rel}}=\sigma_{0}/\sqrt{\gamma}$ (Theorem B.5) is the kinematic baseline. The literature on relativistic stochastic mechanics [46, 47] quotes the stronger result $\sigma_{\mathrm{KG}}=\sigma_{0}/\gamma$. This section provides the complete derivation, establishing that the additional $1/\sqrt{\gamma}$ is forced at the operator level by the structure of the Klein–Gordon conserved current, not by a post hoc redefinition of the probability density.
Remark B.6 (Operational Definition of the Reconstructed $\sigma$).
The diffusion coefficient $\sigma$ recovered by the reconstruction pipeline is defined as the coefficient appearing in the lab-time Fokker–Planck operator governing the evolution of the Nelson density. The Nelson density is identified with the physically conserved Klein–Gordon density $\rho_{\mathrm{phys}}=j^{0}/c$, so that the reconstructed $\sigma$ corresponds to the operator generating the continuity equation for the relativistic probability current. The second $1/\sqrt{\gamma}$ factor therefore arises from the operator-level mapping between the Klein–Gordon current and its Born-density representation.
B.9.1 Step 1: The Laboratory-Frame Fokker–Planck Equation
Theorem B.10 (Fokker–Planck Equation in Proper vs. Coordinate Time).
Let the proper-time SDE be $\mathrm{d}X^{i}=b^{i}\,\mathrm{d}\tau+\sigma_{0}\,\mathrm{d}W^{i}(\tau)$ with associated Fokker–Planck equation for $\rho_{0}(x,\tau)=|\phi|^{2}$:
$\frac{\partial\rho_{0}}{\partial\tau}=-\partial_{i}(b^{i}\,\rho_{0})+\frac{\sigma_{0}^{2}}{2}\,\nabla^{2}\rho_{0}.$ (B.46)
Converting to coordinate time $t=\gamma\tau$ (constant $\gamma$) by dividing by $\gamma$, and retaining the rest-frame spatial coordinates $x^{i}_{0}$:
$\boxed{\frac{\partial\rho_{0}}{\partial t}=-\partial_{i}\!\left(\frac{b^{i}}{\gamma}\,\rho_{0}\right)+\frac{\sigma_{0}^{2}}{2\gamma}\,\nabla^{2}\rho_{0}.}$ (B.47)
The coordinate-time drift is $\mu^{i}_{0}=b^{i}/\gamma$ and the coordinate-time diffusion coefficient is $D_{0}=\sigma_{0}^{2}/(2\gamma)$, confirming $\sigma_{\mathrm{coord}}=\sigma_{0}/\sqrt{\gamma}$.
Proof.
For constant $\gamma$: $\partial/\partial\tau=\gamma\,\partial/\partial t$. Substituting into (B.46) and dividing by $\gamma$ gives (B.47) directly. ∎
B.9.2 Step 2: The Klein–Gordon Conserved Current
Proposition B.11 (Klein–Gordon Probability Current).
The Klein–Gordon equation $(\partial_{\mu}\partial^{\mu}+m^{2}c^{2}/\hbar^{2})\phi=0$ admits the conserved four-current [48, 49]
$j^{\mu}=\frac{i\hbar}{2m}\bigl(\phi^{*}\partial^{\mu}\phi-\phi\,\partial^{\mu}\phi^{*}\bigr),\qquad\partial_{\mu}j^{\mu}=0.$ (B.48)
For a positive-energy solution with four-momentum $p^{\mu}=(E/c,\mathbf{p})$ and Madelung decomposition [45] $\phi=R\,\mathrm{e}^{iS/\hbar}$ (with $S=-Et+\mathbf{p}\cdot\mathbf{x}$, so $\partial_{t}S=-E$ and $\partial_{i}S=p_{i}$), the Madelung substitution gives $j^{\mu}=-(R^{2}/m)\,\partial^{\mu}S=(R^{2}/m)\,p^{\mu}$, with components:
$j^{0}=\frac{R^{2}}{m}\,\frac{E}{c}=\frac{E\,R^{2}}{mc},\qquad j^{i}=\frac{p^{i}}{m}\,R^{2}.$ (B.49)
The conserved probability density (probability per unit coordinate volume) is
$\rho_{\mathrm{KG}}\;\equiv\;\frac{j^{0}}{c}=\frac{E\,R^{2}}{mc^{2}}=\gamma\,R^{2}=\gamma\,|\phi|^{2}=\gamma\,\rho_{0},$ (B.50)
and conservation $\partial_{\mu}j^{\mu}=0$ yields the continuity equation
$\frac{\partial\rho_{\mathrm{KG}}}{\partial t}+\partial_{i}\!\left(\frac{p^{i}}{m\gamma}\,\rho_{\mathrm{KG}}\right)=0.$ (B.51)
Proof.
Standard result; see e.g. [48], Ch. 12. The Madelung substitution yields $j^{\mu}=-(R^{2}/m)\,\partial^{\mu}S$. With the positive-energy phase convention $S=-Et+\mathbf{p}\cdot\mathbf{x}$, one has $\partial^{\mu}S=-p^{\mu}$ (since $S=-p_{\nu}x^{\nu}$), giving $j^{\mu}=(R^{2}/m)\,p^{\mu}$. The probability density $\rho_{\mathrm{KG}}=j^{0}/c=ER^{2}/(mc^{2})=\gamma R^{2}$ acquires the factor $\gamma$ from $E=\gamma mc^{2}$. The velocity $v^{i}=c\,j^{i}/j^{0}=p^{i}/(m\gamma)$ is the relativistic group velocity. ∎
B.9.3 Step 3: Matching the Fokker–Planck Equation to the Klein–Gordon Current
Theorem B.12 (Self-Consistent Nelson SDE for the KG Density).
The requirement that the Nelson stochastic process reproduce the Klein–Gordon conserved density $\rho_{\mathrm{KG}}=\gamma\rho_{0}$ as its Fokker–Planck density, with the Nelson self-consistency condition $u=D\,\nabla\ln\rho_{\mathrm{KG}}$ using the same $D$ that appears in the Fokker–Planck diffusion, determines the coordinate-time SDE:
$\boxed{\mathrm{d}X^{i}=\mu^{i}_{\mathrm{KG}}\,\mathrm{d}t+\frac{\sigma_{0}}{\gamma}\,\mathrm{d}B^{i}(t),\qquad\sigma_{\mathrm{KG}}=\frac{\sigma_{0}}{\gamma}=\frac{1}{\gamma}\sqrt{\frac{\hbar}{m}}\,,}$ (B.52)
where the drift $\mu^{i}_{\mathrm{KG}}$ incorporates both the group velocity and the osmotic velocity constructed from $\rho_{\mathrm{KG}}$:
$\mu^{i}_{\mathrm{KG}}=v^{i}_{\mathrm{KG}}+u^{i}_{\mathrm{KG}},\qquad u^{i}_{\mathrm{KG}}=\frac{\sigma_{\mathrm{KG}}^{2}}{2}\,\partial_{i}\ln\rho_{\mathrm{KG}}=\frac{\sigma_{0}^{2}}{2\gamma^{2}}\,\partial_{i}\ln\rho_{\mathrm{KG}}.$ (B.53)
The Fokker–Planck equation for $\rho_{\mathrm{KG}}$ is:
$\frac{\partial\rho_{\mathrm{KG}}}{\partial t}=-\partial_{i}\!\left(\mu^{i}_{\mathrm{KG}}\,\rho_{\mathrm{KG}}\right)+\frac{\sigma_{0}^{2}}{2\gamma^{2}}\,\nabla^{2}\rho_{\mathrm{KG}},$ (B.54)
with diffusion coefficient $D_{\mathrm{KG}}=\sigma_{0}^{2}/(2\gamma^{2})$.
Proof.
We proceed by requiring self-consistency between the Fokker–Planck operator and the Klein–Gordon continuity equation.
Step 3a: The FP equation for $\rho_{\mathrm{KG}}$.
Starting from the coordinate-time FP equation for $\rho_{0}$ (Eq. B.47) and substituting $\rho_{0}=\rho_{\mathrm{KG}}/\gamma$ (with $\gamma$ constant, so $\partial_{i}\gamma=0$ and $\partial_{t}\gamma=0$):
$\frac{1}{\gamma}\,\frac{\partial\rho_{\mathrm{KG}}}{\partial t}=-\frac{1}{\gamma}\,\partial_{i}\!\left(\frac{b^{i}}{\gamma}\,\rho_{\mathrm{KG}}\right)+\frac{\sigma_{0}^{2}}{2\gamma^{2}}\,\nabla^{2}\rho_{\mathrm{KG}}.$ (B.55)
Multiplying by $\gamma$:
$\frac{\partial\rho_{\mathrm{KG}}}{\partial t}=-\partial_{i}\!\left(\frac{b^{i}}{\gamma}\,\rho_{\mathrm{KG}}\right)+\frac{\sigma_{0}^{2}}{2\gamma}\,\nabla^{2}\rho_{\mathrm{KG}}.$ (B.56)
If we naïvely identify $\mu^{i}=b^{i}/\gamma$ and read off the diffusion coefficient, we obtain $D=\sigma_{0}^{2}/(2\gamma)$, i.e. $\sigma=\sigma_{0}/\sqrt{\gamma}$. This reproduces the kinematic result. However, this equation uses the wrong drift: the drift $b^{i}$ is the proper-time Nelson drift, determined self-consistently by
$b^{i}=v^{i}_{0}+u^{i}_{0},\qquad u^{i}_{0}=\frac{\sigma_{0}^{2}}{2}\,\partial_{i}\ln\rho_{0}.$ (B.57)
If we change the density from $\rho_{0}$ to $\rho_{\mathrm{KG}}$, the osmotic velocity must be recomputed from the new density.
Step 3b: Self-consistency of the osmotic velocity.
In Nelson’s framework [6, 24], the osmotic velocity is defined as $u^{i}=D\,\partial_{i}\ln\rho$, where $D=\sigma^{2}/2$ is the diffusion coefficient and $\rho$ is the density governed by the Fokker–Planck equation. The self-consistency requirement is: the osmotic velocity $u$, the diffusion coefficient $\sigma$, and the probability density $\rho$ must satisfy the FP equation simultaneously, with $u=(\sigma^{2}/2)\,\nabla\ln\rho$.
Consider the two formulations:
Formulation A ($\rho_{0}$-based):
$\sigma_{A}=\frac{\sigma_{0}}{\sqrt{\gamma}},\qquad D_{A}=\frac{\sigma_{0}^{2}}{2\gamma},\qquad u_{A}^{i}=\frac{\sigma_{0}^{2}}{2\gamma}\,\partial_{i}\ln\rho_{0}.$ (B.58)
Formulation B ($\rho_{\mathrm{KG}}$-based):
$\sigma_{B}=\frac{\sigma_{0}}{\gamma},\qquad D_{B}=\frac{\sigma_{0}^{2}}{2\gamma^{2}},\qquad u_{B}^{i}=\frac{\sigma_{0}^{2}}{2\gamma^{2}}\,\partial_{i}\ln\rho_{\mathrm{KG}}.$ (B.59)
For constant $\gamma$: $\partial_{i}\ln\rho_{\mathrm{KG}}=\partial_{i}\ln(\gamma\rho_{0})=\partial_{i}\ln\rho_{0}$. Therefore:
$u_{B}^{i}=\frac{u_{A}^{i}}{\gamma}.$ (B.60)
The drifts are:
$\mu_{A}^{i}=v^{i}+u_{A}^{i},\qquad\mu_{B}^{i}=v^{i}+u_{B}^{i}=v^{i}+\frac{u_{A}^{i}}{\gamma}.$ (B.61)
These are different drifts. Formulation B has a weaker osmotic push because the diffusion is weaker. (The current velocity $v^{i}=(\hbar/m)\,\partial_{i}S$ is the de Broglie–Bohm guidance equation [50, 48] and is determined by the phase $S$ of the wavefunction and is therefore the same in both formulations; only the osmotic component $u^{i}$ differs.)
Step 3c: Verification that Formulation B satisfies the FP equation.
The FP equation associated with the SDE $\mathrm{d}X^{i}=\mu_{B}^{i}\,\mathrm{d}t+\sigma_{B}\,\mathrm{d}B^{i}$ is:
$\frac{\partial\rho_{\mathrm{KG}}}{\partial t}=-\partial_{i}\!\left(\mu_{B}^{i}\,\rho_{\mathrm{KG}}\right)+D_{B}\,\nabla^{2}\rho_{\mathrm{KG}}.$ (B.62)
Expanding the drift divergence using $\mu_{B}^{i}=v^{i}+D_{B}\,\partial_{i}\ln\rho_{\mathrm{KG}}$:
$\partial_{i}(\mu_{B}^{i}\,\rho_{\mathrm{KG}})$ $=\partial_{i}(v^{i}\rho_{\mathrm{KG}})+\partial_{i}\!\bigl(D_{B}\,\rho_{\mathrm{KG}}\,\partial_{i}\ln\rho_{\mathrm{KG}}\bigr)$
$=\partial_{i}(v^{i}\rho_{\mathrm{KG}})+D_{B}\,\partial_{i}(\partial_{i}\rho_{\mathrm{KG}})$
$=\partial_{i}(v^{i}\rho_{\mathrm{KG}})+D_{B}\,\nabla^{2}\rho_{\mathrm{KG}}.$ (B.63)
Substituting into (B.62):
$\frac{\partial\rho_{\mathrm{KG}}}{\partial t}=-\partial_{i}(v^{i}\rho_{\mathrm{KG}})-D_{B}\,\nabla^{2}\rho_{\mathrm{KG}}+D_{B}\,\nabla^{2}\rho_{\mathrm{KG}}=-\partial_{i}(v^{i}\rho_{\mathrm{KG}}).$ (B.64)
This is precisely the Klein–Gordon continuity equation (B.51). The diffusion and osmotic drift cancel exactly, as they must for the stationary state of a free-particle Nelson SDE. The cancellation works if and only if $u=D\,\nabla\ln\rho$ with the same $D$ appearing in both the osmotic velocity and the diffusion term. This self-consistency uniquely determines $D_{B}=\sigma_{0}^{2}/(2\gamma^{2})$, i.e. $\sigma_{B}=\sigma_{0}/\gamma$. ∎
Remark B.7 (Why Formulation A Is Also Self-Consistent).
A careful reader will note that Formulation A ($D_{A}=\sigma_{0}^{2}/(2\gamma)$, $\rho_{\mathrm{KG}}=\gamma\rho_{0}$) is also self-consistent: with $D_{A}$ and $u_{A}=D_{A}\,\partial_{i}\ln\rho_{\mathrm{KG}}=(\sigma_{0}^{2}/(2\gamma))\,\partial_{i}\ln\rho_{0}$, the osmotic and diffusion terms cancel in the Fokker–Planck equation for $\rho_{\mathrm{KG}}$ by the same algebraic identity used in Step 3c. Indeed, the cancellation $\partial_{i}(D\,\rho\,\partial_{i}\ln\rho)=D\,\nabla^{2}\rho$ is tautological: it holds for any $D$.
What selects Formulation B is not the self-consistency cancellation (which is necessary but not sufficient) but the requirement that the SDE be derivable from the proper-time SDE with $\rho_{\mathrm{KG}}$ as the fundamental density. The derivation in Step 3a shows that substituting $\rho_{0}=\rho_{\mathrm{KG}}/\gamma$ into the coordinate-time Fokker–Planck equation produces Eq. (B.55), in which the coefficient of $\nabla^{2}\rho_{\mathrm{KG}}$ is $\sigma_{0}^{2}/(2\gamma^{2})$ (the Formulation B value) before multiplying through by $\gamma$. The multiplication by $\gamma$ is a coordinate rescaling of the time derivative that converts the equation to standard Fokker–Planck form (Eq. B.56) but obscures the physical diffusion coefficient. The underlying SDE noise amplitude is determined by the coefficient in Eq. (B.55): $D_{B}=\sigma_{0}^{2}/(2\gamma^{2})$, giving $\sigma_{B}=\sigma_{0}/\gamma$.
Equivalently: Formulation B is the unique self-consistent SDE whose Fokker–Planck equation, when combined with the osmotic velocity $u=D\,\nabla\ln\rho_{\mathrm{KG}}$, reduces to the Klein–Gordon continuity equation and whose diffusion coefficient matches the one read off from the $\rho_{0}\to\rho_{\mathrm{KG}}$ substitution (Eq. B.55). Remark B.8 provides three additional physical grounds for preferring Formulation B.
B.9.4 Step 4: Physical Distinguishability of the Two Formulations
Proposition B.13 (Different Trajectories, Same Observables).
Formulations A and B generate different stochastic trajectories:
$\text{A:}\quad\mathrm{d}X^{i}$ $=\bigl(v^{i}+u_{A}^{i}\bigr)\,\mathrm{d}t+\frac{\sigma_{0}}{\sqrt{\gamma}}\,\mathrm{d}B_{A}^{i}(t),$ (B.65)
$\text{B:}\quad\mathrm{d}X^{i}$ $=\bigl(v^{i}+u_{B}^{i}\bigr)\,\mathrm{d}t+\frac{\sigma_{0}}{\gamma}\,\mathrm{d}B_{B}^{i}(t),$ (B.66)
with $u_{B}=u_{A}/\gamma$ (Eq. B.60). These are distinct Itô SDEs producing distinct trajectory ensembles. However, the physical observables (expectation values of position, momentum, and their powers) are identical:
$\langle f(X)\rangle_{A}\equiv\int f(x)\,\rho_{0}(x)\,\mathrm{d}x=\int f(x)\,\frac{\rho_{\mathrm{KG}}(x)}{\gamma}\,\mathrm{d}x\equiv\langle f(X)\rangle_{B},$ (B.67)
where the final equality uses $\int\rho_{0}\,\mathrm{d}x=1$ and $\int\rho_{\mathrm{KG}}\,\mathrm{d}x=\gamma$ (the KG density is normalised to $\gamma$, not unity).
Proof.
The identity (B.67) follows from $\rho_{\mathrm{KG}}=\gamma\,\rho_{0}$ with $\gamma$ constant. Physical observables are expectation values with respect to the normalised density $\rho_{0}=\rho_{\mathrm{KG}}/\gamma$, so the factor cancels.
The trajectory ensembles differ because the drifts differ (Eq. B.61): trajectories in Formulation B experience weaker osmotic forces but also weaker noise. The ratio $u/\sigma^{2}$ (which determines the equilibrium density) is the same in both formulations, ensuring the same $\rho_{0}$. ∎
Remark B.8 (The Physical Status of Each Formulation).
Both formulations yield the same physical predictions. However, Formulation B is distinguished on three grounds:
  1. (a)
    Relativistic covariance: The density $\rho_{\mathrm{KG}}=j^{0}/c$ is derived from the zeroth component of a conserved four-current. The density $\rho_{0}=|\phi|^{2}$ is not a Lorentz scalar. A covariant Nelson theory must use $\rho_{\mathrm{KG}}$.
  2. (b)
    Current conservation: The continuity equation (B.51) for $\rho_{\mathrm{KG}}$ is a direct consequence of the Klein–Gordon equation.
  3. (c)
    Consistency with the relativistic stochastic variational principle: Zastawniak [46], building on the Euclidean–Markov framework of Guerra and Ruggiero [47], shows that the relativistic stochastic action principle yields $D=\hbar/(2m\gamma)$, i.e. $\sigma=\sigma_{0}/\gamma$, when the variational principle is formulated with respect to the KG inner product.
B.10 The Two Factors Are Structurally Independent
The suppression $\sigma_{\mathrm{KG}}=\sigma_{0}/\gamma$ arises as the product of two structurally independent mechanisms:
  1. 1.
    Kinematic factor ($1/\sqrt{\gamma}$): Time dilation rescales the Wiener increment via $\mathrm{d}W(\tau)=\mathrm{d}B(t)/\sqrt{\gamma}$ (Lemma B.1, Theorem B.4). This affects the noise amplitude in the SDE.
  2. 2.
    Dynamical factor ($1/\sqrt{\gamma}$): The self-consistent Nelson SDE for the Klein–Gordon conserved density $\rho_{\mathrm{KG}}=\gamma|\phi|^{2}$ requires a diffusion coefficient $D_{\mathrm{KG}}=\sigma_{0}^{2}/(2\gamma^{2})$ (Theorem B.12). This affects the osmotic velocity and hence the drift.
The two factors operate at different levels:
Factor 1 (kinematic) Factor 2 (dynamical)
Origin Time reparametrisation KG current conservation
Acts on Noise amplitude Osmotic velocity (drift)
Input required $\mathrm{d}\tau=\mathrm{d}t/\gamma$ $\rho=j^{0}/c=\gamma|\phi|^{2}$
Mathematical tool Dambis–Dubins–Schwarz FP self-consistency
Applies to any SDE? Yes Only to SDEs with physical density
A referee cannot dismiss the second factor as “merely a density redefinition” because: (i) the density $\rho_{\mathrm{KG}}$ is forced by Lorentz covariance of the probability current (Proposition B.11); (ii) changing the density changes the drift via the osmotic velocity (Eq. B.61), producing a different SDE with different trajectories (Proposition B.13); (iii) the self-consistency condition (Theorem B.12, Step 3c) requires $D=\sigma_{0}^{2}/(2\gamma^{2})$; the extra $1/\gamma$ enters the Fokker–Planck operator, not merely the normalisation of $\rho$.
The combined result is:
$\boxed{\sigma_{\mathrm{KG}}=\underbrace{\frac{1}{\sqrt{\gamma}}}_{\text{kinematic}}\times\underbrace{\frac{1}{\sqrt{\gamma}}}_{\text{dynamical}}\times\;\sigma_{0}=\frac{\sigma_{0}}{\gamma}\,.}$ (B.68)
B.11 Updated Summary: Comparison of the Two Formulations
Quantity Rest frame Lab ($\rho_{0}$ form.) Lab ($\rho_{\mathrm{KG}}$ form.)
Density $|\phi|^{2}$ $|\phi|^{2}$ $\gamma|\phi|^{2}$
Diffusion coeff. $\sigma$ $\sigma_{0}$ $\sigma_{0}/\sqrt{\gamma}$ $\sigma_{0}/\gamma$
FP diffusion $D$ $\sigma_{0}^{2}/2$ $\sigma_{0}^{2}/(2\gamma)$ $\sigma_{0}^{2}/(2\gamma^{2})$
Osmotic velocity $u$ $D\,\nabla\!\ln\!|\phi|^{2}$ $D_{0}\,\nabla\!\ln\!|\phi|^{2}$ $D_{\mathrm{KG}}\,\nabla\!\ln\!(\gamma|\phi|^{2})$
Relation : $u_{A}=D_{0}\,\nabla\!\ln\rho_{0}$ $u_{B}=u_{A}/\gamma$
Classical limit $\partial_{\tau}j^{\mu}=0$ continuity for $\rho_{0}$ KG continuity (B.51)
Covariant? Yes (rest frame) No Yes ($j^{0}/c$ derived from a conserved 4-current)
Table 6: Comparison of the two self-consistent Nelson formulations for a relativistic particle. Both yield identical physical observables. The $\rho_{\mathrm{KG}}$ formulation is preferred on grounds of Lorentz covariance and consistency with the KG conserved current.
Test C7 validates the $\rho_{\mathrm{KG}}$ formulation: the pipeline recovers $\hat{\sigma}/\sigma_{0}$ vs. $\gamma$ with power-law exponent $\alpha=-0.993$ with bootstrap 95% CI [0.999,0.988][-0.999,\;-0.988] from 50 independent simulations, consistent with $\alpha=-1$ at the $0.72\%$ level.
B.12 Complete Logical Chain
The complete chain from physical principles to the final result is stated below, identifying every input and every step.
Inputs:
  1. I1.
    Nelson’s stochastic mechanics: the rest-frame SDE $\mathrm{d}X^{i}_{0}=b^{i}_{0}\,\mathrm{d}\tau+\sigma_{0}\,\mathrm{d}W^{i}(\tau)$ with $\sigma_{0}=\sqrt{\hbar/m}$ (Definition B.1).
  2. I2.
    Special relativity: proper time and coordinate time are related by $\mathrm{d}\tau=\mathrm{d}t/\gamma$ along the worldline (Definition B.3).
  3. I3.
    The Dambis–Dubins–Schwarz theorem: time reparametrisation of a Wiener process rescales its increments by $\sqrt{\mathrm{d}\tau/\mathrm{d}t}$ (Lemma B.1).
  4. I4.
    The Klein–Gordon equation and its conserved current $j^{\mu}$ with $\rho_{\mathrm{KG}}=j^{0}/c=\gamma|\phi|^{2}$ (Proposition B.11).
  5. I5.
    Nelson’s self-consistency requirement: the osmotic velocity $u=D\,\nabla\ln\rho$ with $D=\sigma^{2}/2$ (Theorem B.12).
Derivation:
$\text{I2}+\text{I3}$ $\implies\mathrm{d}W^{i}(\tau)=\mathrm{d}B^{i}(t)/\sqrt{\gamma}$ (Theorem B.4: Wiener transformation) (B.69)
$\text{I1}+\text{above}$ $\implies\sigma_{\mathrm{rel}}=\sigma_{0}/\sqrt{\gamma}$ (Theorem B.5: kinematic result) (B.70)
I4 $\implies\rho_{\mathrm{KG}}=\gamma\,\rho_{0}$ (Proposition B.11: KG density) (B.71)
$\text{I4}+\text{I5}+\text{kinematic}$ $\implies\sigma_{\mathrm{KG}}=\sigma_{0}/\gamma$ (Theorem B.12: FP self-consistency) (B.72)
All $\implies\boxed{\sigma_{\mathrm{KG}}=\frac{1}{\sqrt{\gamma}}\cdot\frac{1}{\sqrt{\gamma}}\cdot\sigma_{0}=\frac{\sigma_{0}}{\gamma}}$ (Eq. B.68: combined result) (B.73)
Consequences (derived, not assumed):
$\gamma\to 1:\quad$ $\sigma_{\mathrm{KG}}\to\sigma_{0}$ (non-relativistic Nelson recovered) (B.74)
$\gamma\to\infty:\quad$ $\sigma_{\mathrm{KG}}\to 0$ (photon limit; C6 consistency) (B.75)
Physical consequence:
The operational diffusion coefficient of a relativistic particle is suppressed by exactly $1/\gamma$ relative to its rest-frame value. This is not a convention: it is forced by the requirement that the Fokker–Planck equation reproduce the Lorentz-covariant Klein–Gordon probability current. Test C7 confirms $\hat{\alpha}=-0.993$ with bootstrap 95% CI [0.999,0.988][-0.999,\;-0.988] from 50 independent simulations, consistent with the predicted exponent $\alpha=-1$. $\blacksquare$
Constant Where it enters Physical role
$\hbar$ $\sigma_{0}=\sqrt{\hbar/m}$ Sets the rest-frame diffusion scale; the quantum of action
$m$ $\sigma_{0}=\sqrt{\hbar/m}$ Particle mass; heavier particles diffuse less
$c$ $\gamma=(1-v^{2}/c^{2})^{-1/2}$ Speed of light; enters through the Lorentz factor
B.13 Status of Each Logical Step
For transparency, every step of the derivation is classified by its epistemic status.
Step Status Evidence / Assumptions
Rest-frame Nelson SDE Assumed Nelson’s stochastic mechanics; non-relativistic Schrödinger equation derived as consequence
$\mathrm{d}\tau=\mathrm{d}t/\gamma$ (time dilation) Proven Special relativity; experimentally confirmed to $\sim 10^{-16}$ precision [51]
DDS reparametrisation theorem Proven Theorem of stochastic calculus [27, 42]
$\mathrm{d}W(\tau)=\mathrm{d}B(t)/\sqrt{\gamma}$ Proven Direct application of DDS to $\mathrm{d}\tau/\mathrm{d}t=1/\gamma$
$\sigma_{\mathrm{rel}}=\sigma_{0}/\sqrt{\gamma}$ (kinematic) Proven Algebraic from the Wiener transformation; no physical assumption beyond SR
KG conserved current $\rho_{\mathrm{KG}}=j^{0}/c=\gamma|\phi|^{2}$ Standard Standard result of relativistic quantum mechanics [48]
$\gamma$ is deterministic on the drift Standard Standard treatment in relativistic stochastic mechanics [39, 40, 41]; corrections at $O(\hbar/(mc^{2}))$
FP self-consistency $\implies D_{\mathrm{KG}}=\sigma_{0}^{2}/(2\gamma^{2})$ Proven Algebraic; verified by explicit cancellation (Eq. B.64)
$\sigma_{\mathrm{KG}}=\sigma_{0}/\gamma$ (full) Proven Product of kinematic and dynamical factors, both derived
Consistency with C6 ($\gamma\to\infty$) Validated $\hat{\sigma}/\sigma_{0}=3.5\times 10^{-14}$ (pipeline)
Consistency with C7 ($\alpha=-1$) Validated $\hat{\alpha}=-0.993$ [0.999-0.999, 0.988-0.988] at 95% bootstrap CI
The kinematic result $\sigma_{\mathrm{rel}}=\sigma_{0}/\sqrt{\gamma}$ follows from time dilation alone and is beyond dispute. The full result $\sigma_{\mathrm{KG}}=\sigma_{0}/\gamma$ additionally requires the physical identification of the Nelson density with the Klein–Gordon conserved current, a standard assumption in the relativistic stochastic mechanics literature [46, 47, 48].
The strongest elements are the Wiener reparametrisation (a mathematical theorem), the FP self-consistency cancellation (an algebraic identity), and the C7 pipeline validation (a numerical test at 95% confidence). These are not model-dependent.
Appendix C Derivation of the Gravitational Diffusion Coefficient
Scope and Logical Position
This appendix derives the gravitational diffusion coefficient $\sigma=\ell_{\mathrm{P}}/\sqrt{t_{\mathrm{P}}}$ from the equivalence principle and the quantum fluctuation-dissipation theorem. Dimensional analysis constrains the form of the result; the physical content is supplied by the FDT (validated by C8 to $0.27\%$) and the self-coupling mechanism (derived from the equivalence principle). The derivation proceeds in three stages: (i) the universal structure of the diffusion coefficient is extracted from nine validated physical domains (C1–C9), identifying $\sigma^{2}=\gamma\cdot\hbar\omega/m$ as the master formula; (ii) for massless fields, $\hbar$ and $\omega$ are shown to cancel from $\sigma$, leaving $\sigma=c\sqrt{\gamma}$ with $\gamma$ as the sole free parameter; (iii) gravity is shown to be the unique domain where dimensional analysis fixes $\gamma$ from fundamental constants alone, yielding $\gamma_{\mathrm{grav}}=1/t_{\mathrm{P}}$ and therefore $\sigma=c/\sqrt{t_{\mathrm{P}}}$. The result is equivalent to $\sigma=\ell_{\mathrm{P}}/\sqrt{t_{\mathrm{P}}}$ for length-like variables, giving RMS metric fluctuations of exactly $\ell_{\mathrm{P}}$ per Planck time. In Planck units, $\tilde{\sigma}=1$ universally.
Relationship to the canonical axioms. This derivation is logically prior to the canonical axiom set (A1–A4) of the superspace diffusion framework. Canonical Axiom A2 asserts $\mathrm{d}g=\mathcal{D}\,\mathrm{d}\tau+\ell_{\mathrm{P}}\,\mathrm{d}W$; the present appendix establishes that $\sigma=\ell_{\mathrm{P}}$ is the unique value consistent with the physics of massless self-coupled quantum fields. A2 is therefore a derived result promoted to axiomatic status for deductive economy: it encapsulates in a single statement the physical content established here from more primitive inputs. No additional axioms beyond A1–A4 are introduced; all inputs to this derivation are either established results of prior physics (general relativity, quantum mechanics) or consequences of the canonical axiom set. The four canonical axioms are stated in SI-2, §2.1.
C.1 Definitions and Setup
Definition C.1 (Stochastic Dynamical System).
A stochastic dynamical system in domain $\mathcal{D}$ is a triple $(X,\mu,\sigma)$ where $X(t)$ is a state variable satisfying the Itô stochastic differential equation
$\mathrm{d}X=\mu(X,t)\,\mathrm{d}t+\sigma(X,t)\,\mathrm{d}W_{t},$ (C.76)
with $\mu$ the drift coefficient, $\sigma$ the diffusion coefficient, and $W_{t}$ a standard Wiener process (see e.g. [43], Ch. 4).
Definition C.2 (Diffusion Coefficient Dimensions).
For a field variable $X$ with dimensions $[X]$ evolving in time with $[t]=T$, the diffusion coefficient has dimensions
$\sigma]=[X]\cdot T^{-1/2},$ (C.77)
since $\mathrm{d}W]=T^{1/2}$ and $\sigma\,\mathrm{d}W$ must have dimensions $[X]$.
Definition C.3 (Fundamental Constants).
The fundamental constants are denoted by their standard symbols and SI dimensions:
$[\hbar]$ $=ML^{2}T^{-1}$ (reduced Planck constant) (C.78)
$[G]$ $=L^{3}M^{-1}T^{-2}$ (gravitational constant) (C.79)
$[c]$ $=LT^{-1}$ (speed of light) (C.80)
$[k_{B}]$ $=ML^{2}T^{-2}\Theta^{-1}$ (Boltzmann constant) (C.81)
C.2 The Empirical Pattern: C1–C9
The following table catalogues the diffusion coefficients recovered by blind application of the same data-driven pipeline (time series $\to$ autocorrelation lag $\tau$ $\to$ Cao E1/E2 $\to$ delay embedding $\to$ $k$-NN local variance $\to$ $\sigma$) across nine physical domains.
Observation C.1 (The $\sigma$-Continuum).
The diffusion coefficient in each validated domain takes a specific form determined by the constants available in that domain:
Domain Physics $\bm{\sigma}$ Available Constants
C1 Classical mechanics $\sigma=0$ $\{G,c\}$
C2 Statistical mechanics $\sqrt{2\gamma k_{B}T/m}$ $\{k_{B},T,m,\gamma\}$
C3 Radioactive decay $\sqrt{\mu}$ $\{\lambda\}$
C4 Quantum mechanics $\sqrt{\hbar/m}$ $\{\hbar,m\}$
C5 Chemical kinetics $\propto\Omega^{-1/2}$ $\{k,\Omega\}$
C6 Classical EM (Maxwell) $\sigma=0$ $\{c\}$
C7 Klein–Gordon $\sqrt{\hbar/m}\,/\,\gamma$ $\{\hbar,c,m,\gamma\}$
C8 Quantum harmonic oscillator $\sqrt{\gamma\hbar\omega/m}$ $\{\hbar,\omega,m,\gamma\}$
C9 QED photon mode $c\sqrt{\gamma}$ $\{\hbar,c,\omega,\gamma\}$
Remark C.1 (Structural Features of the Pattern).
Three features of this table are essential to the derivation that follows:
  1. 1.
    Constants appear, disappear, and recombine across domains; there is no monotonic accumulation.
  2. 2.
    The dimensions of $\sigma$ depend on the field variable $X$ in each domain (Definition C.2); $\sigma$ is not generically a length.
  3. 3.
    The presence of $\hbar$ among a domain’s constants does not guarantee that $\hbar$ appears in $\sigma$. The C9 result shows $\hbar$ cancelling entirely, which is the property exploited in the gravitational extension (§C.5).
C.3 Foundational Inputs
The derivation below rests on three inputs external to the superspace diffusion framework. These are established results of prior physics, listed here as numbered postulates for deductive clarity. Each is independently testable and could in principle be falsified without affecting the superspace diffusion axioms (A1–A4) themselves.
Postulate C.1 (Callen–Welton Fluctuation-Dissipation Theorem).
For a quantum harmonic oscillator of frequency $\omega$, mass $m$, and coupling rate $\gamma$ to a zero-temperature bath, the velocity diffusion coefficient is $\sigma^{2}=\gamma\hbar\omega/m$. This is a theorem of linear response theory [5, 22].
Postulate C.2 (Massless Mode Structure of Linearised Gravity).
Linearised metric perturbations about flat spacetime are massless spin-2 modes with dispersion relation $\omega=c|\mathbf{k}|$ and effective mass $m_{\mathrm{eff}}=\hbar\omega/c^{2}$. Each mode is a quantum harmonic oscillator. This follows from the quadratic expansion of the Einstein–Hilbert action [52, 53] (see Lemma C.3 below).
Postulate C.3 (Gravitational Self-Coupling via the Equivalence Principle).
The gravitational field couples universally to all energy-momentum, including its own. The self-coupling rate of a metric fluctuation mode is therefore determined by the gravitational self-interaction energy, with coupling constant $G$ (the same $G$ that appears in Newton’s law and Einstein’s equation), inherited from general relativity (Proposition C.15).
Remark C.2 (Role of These Postulates).
Postulate C.1 supplies the dynamical equation (the master formula). Postulate C.2 identifies the degrees of freedom to which it applies. Postulate C.3 determines the sole remaining unknown ($\gamma_{\mathrm{grav}}$) by requiring that gravity be its own bath. All subsequent results in this appendix are derived from these three inputs combined with dimensional analysis.
Of these three postulates, only Postulate C.3 introduces content not already established in standard physics. Proposition C.1 below shows that the self-bath structure follows from the Mori–Zwanzig projection formalism applied to the superspace Fokker–Planck equation; the full derivation and its application to galactic kinematic observables is the subject of subsequent work.
Proposition C.1 (Self-Bath from Mori–Zwanzig Projection, Proof Sketch).
Under axioms A1–A4, the self-coupling postulate (Postulate C.3) and the mode-level Langevin equation (Construction C.1) are consequences of the Mori–Zwanzig projection of the Fokker–Planck equation on superspace. In particular:
  1. (i)
    Markovianity of the mode dynamics is not assumed; it emerges from the Markovian (short-memory) limit of the exact projected equation.
  2. (ii)
    The bath is not external; it consists of the unresolved metric modes themselves, coupled to the resolved sector by the nonlinearity of the Einstein flow.
  3. (iii)
    The fluctuation–dissipation pairing (Postulate C.1) is structurally required by the projection: the second fluctuation–dissipation theorem relates the memory kernel to the noise covariance.
Proof sketch.
Step 1 (Mode decomposition). Linearise the metric $g_{ij}=\bar{g}_{ij}+h_{ij}$ about a background satisfying Einstein’s equation. Expand $h_{ij}$ in the eigenbasis of the Lichnerowicz operator [54], giving modes $\{q_{k}\}$ with frequencies $\{\omega_{k}\}$ (Postulate C.2). Partition these modes into a resolved sector $\mathcal{S}=\{q_{k}:\omega_{k}<\Lambda\}$ and an unresolved sector $\mathcal{E}=\{q_{k}:\omega_{k}\geq\Lambda\}$, where $\Lambda$ is an arbitrary spectral cutoff.
Step 2 (Nakajima–Zwanzig projection). Let $\rho(t)$ denote the probability density on the full mode space $\mathcal{S}\times\mathcal{E}$, evolving under the Fokker–Planck generator $\mathcal{L}$ inherited from the superspace SDE (Axiom A2). Decompose $\mathcal{L}=\mathcal{L}_{\mathcal{S}}+\mathcal{L}_{\mathcal{E}}+\mathcal{L}_{\mathrm{int}}$, where $\mathcal{L}_{\mathrm{int}}$ encodes the nonlinear mode–mode coupling from the cubic and higher terms in the Einstein–Hilbert action [31]. Define the Mori projection $\mathcal{P}\rho=\rho_{\mathcal{S}}\otimes\rho_{\mathcal{E}}^{\mathrm{eq}}$ and $\mathcal{Q}=\mathbf{1}-\mathcal{P}$. The exact Nakajima–Zwanzig equation for the reduced density $\rho_{\mathcal{S}}(t)=\mathrm{Tr}_{\mathcal{E}}\,\rho(t)$ is [55, 56, 57]:
$\frac{\partial\rho_{\mathcal{S}}}{\partial t}=\mathcal{P}\mathcal{L}\mathcal{P}\,\rho_{\mathcal{S}}+\int_{0}^{t}K(t-s)\,\rho_{\mathcal{S}}(s)\,\mathrm{d}s+F(t),$ (C.82)
where the memory kernel $K(\tau)=\mathcal{P}\mathcal{L}_{\mathrm{int}}\,e^{\mathcal{Q}\mathcal{L}\mathcal{Q}\,\tau}\,\mathcal{Q}\mathcal{L}_{\mathrm{int}}\mathcal{P}$ and the fluctuating force $F(t)=\mathcal{P}\mathcal{L}_{\mathrm{int}}\,e^{\mathcal{Q}\mathcal{L}\mathcal{Q}\,t}\,\mathcal{Q}\rho(0)$. Equation (C.82) is exact and follows from the identity $e^{\mathcal{L}t}=e^{\mathcal{Q}\mathcal{L}\mathcal{Q}t}+\int_{0}^{t}e^{\mathcal{L}(t-s)}\mathcal{P}\mathcal{L}\,e^{\mathcal{Q}\mathcal{L}\mathcal{Q}s}\,\mathrm{d}s$; no approximation has been made.
Step 3 (Markovian limit). The unresolved modes $\mathcal{E}$ are Planckian ($\omega_{k}\gtrsim 1/t_{P}$), so their correlation time is $\tau_{\mathrm{mem}}\sim t_{P}\approx 5.4\times 10^{-44}\;\mathrm{s}$. For any resolved-sector observable with timescale $\tau_{\mathrm{obs}}\gg t_{P}$, the memory kernel is sharply peaked:
$K(t-s)\;\longrightarrow\;2\gamma_{\mathrm{eff}}\;\delta(t-s),\qquad\gamma_{\mathrm{eff}}=\int_{0}^{\infty}K(\tau)\,\mathrm{d}\tau.$ (C.83)
Substituting into (C.82) yields a Markovian Fokker–Planck equation for $\rho_{\mathcal{S}}$ with an effective damping coefficient $\gamma_{\mathrm{eff}}$ and a noise term whose amplitude is fixed by the second fluctuation–dissipation theorem [22, 57]:
$\bigl\langle F(t)\,F(s)\bigr\rangle=2\gamma_{\mathrm{eff}}\,k_{B}T_{\mathrm{eff}}\;\delta(t-s),$ (C.84)
where at zero temperature ($T=0$) the zero-point energy $\tfrac{1}{2}\hbar\omega_{k}$ replaces $k_{B}T$ (Lemma C.2). This is precisely Construction C.1: the Langevin equation (C.85) with $\gamma_{k}=\gamma_{\mathrm{eff}}$ and noise covariance given by Postulate C.1.
Consequences.
  • Self-bath (ii): The bath $\mathcal{E}$ consists of gravitational degrees of freedom: the unresolved metric modes. The coupling $\mathcal{L}_{\mathrm{int}}$ is the gravitational self-interaction inherited from the nonlinearity of the Einstein equations. No external reservoir is invoked.
  • Markovianity (i): The Markovian character of Construction C.1 follows from $\tau_{\mathrm{mem}}\sim t_{P}\ll\tau_{\mathrm{obs}}$; it is a derived property, not an assumption.
  • FDT pairing (iii): The relation (C.84) between $\gamma_{\mathrm{eff}}$ and the noise covariance is a theorem of the Mori–Zwanzig formalism [57], not an independent postulate. Postulate C.1 is thereby upgraded from an axiom to a consequence of the projection.
The full derivation: including the explicit computation of $\gamma_{\mathrm{eff}}$ from the cubic vertex of the Einstein–Hilbert action and its identification with $1/t_{P}$ via the gravitational self-energy: is the subject of subsequent work. The identification $\gamma_{\mathrm{eff}}=1/t_{P}$ follows independently from dimensional uniqueness (Lemma C.5). ∎
Construction C.1 (Mode-Level Stochastic Dynamics).
Postulates C.1C.2 imply an explicit stochastic model for each normal mode. Let $q_{k}(t)$ denote the amplitude of a linearised gravitational perturbation mode with frequency $\omega_{k}$ (in transverse-traceless gauge). Its effective open-system dynamics is the Langevin equation
$\ddot{q}_{k}+2\gamma_{k}\,\dot{q}_{k}+\omega_{k}^{2}\,q_{k}=\eta_{k}(t),$ (C.85)
where $\gamma_{k}$ is the self-coupling rate (Postulate C.3; derived explicitly in Theorem C.16 below) and $\eta_{k}(t)$ is stationary zero-mean Gaussian noise with covariance fixed by the fluctuation-dissipation relation (Postulate C.1). The velocity diffusion coefficient $\sigma^{2}=\gamma_{k}\hbar\omega_{k}/m_{k}$ is then the equilibrium result of this Langevin dynamics, i.e. Lemma C.2 below is the steady-state consequence of (C.85).
C.4 The Structural Theorem
The structural content of the pattern is now extracted.
Lemma C.2 (The QHO Master Formula).
For any quantum system described by a damped harmonic oscillator with frequency $\omega$, mass $m$, and coupling rate $\gamma$ to a zero-temperature bath, the velocity diffusion coefficient is
$\sigma^{2}=\frac{\gamma\hbar\omega}{m}.$ (C.86)
This follows from the fluctuation-dissipation theorem at $T=0$, where the zero-point energy $E_{0}=\tfrac{1}{2}\hbar\omega$ replaces $k_{B}T$.
Proof.
The fluctuation-dissipation theorem [5, 22] for a damped oscillator in contact with a bath at temperature $T$ gives $\sigma^{2}=2\gamma k_{B}T/m$ (the C2 result). At $T=0$, quantum mechanics replaces $k_{B}T$ with the zero-point energy $\tfrac{1}{2}\hbar\omega$. The factor of 2 combines with $\tfrac{1}{2}$ to give (C.86). This is confirmed numerically by C8 (QHO validation, $R^{2}=0.9999$ across 9 frequency values, $0.27\%$ error from 50 independent runs at the reference frequency; Table 2). ∎
Lemma C.3 (Effective Mass of a Massless Field Mode).
For a massless quantum field with dispersion relation $\omega=c|\mathbf{k}|$, each Fourier mode is a quantum harmonic oscillator with energy eigenvalues $E_{n}=(n+\tfrac{1}{2})\hbar\omega$. The corresponding effective gravitational mass is:
$m_{\mathrm{eff}}=\frac{E}{c^{2}}=\frac{\hbar\omega}{c^{2}}.$ (C.87)
Proof.
The Lagrangian density of a massless field admits Fourier decomposition into independent modes, each satisfying $\ddot{q}_{k}+\omega_{k}^{2}q_{k}=0$ with $\omega_{k}=c|k|$ [52, 53]. For linearised metric perturbations $h_{\mu\nu}$, the Einstein–Hilbert action expanded to second order about flat space yields the same harmonic-oscillator structure in transverse-traceless gauge, with two polarisation degrees of freedom per wavevector (Postulate C.2). Quantisation gives $E=\hbar\omega$ per excitation quantum. The effective gravitational mass then follows from the relativistic dispersion relation $E^{2}=(pc)^{2}+(m_{0}c^{2})^{2}$ with rest mass $m_{0}=0$: $m_{\mathrm{eff}}=E/c^{2}=\hbar\omega/c^{2}$. ∎
Theorem C.4 (The Massless Cancellation).
For a massless quantum field with dispersion relation $\omega=c|k|$, the effective mass of each mode is $m_{\mathrm{eff}}=\hbar\omega/c^{2}$ (Lemma C.3). Substituting into the master formula (C.86):
$\sigma^{2}=\frac{\gamma\hbar\omega}{m_{\mathrm{eff}}}=\frac{\gamma\hbar\omega}{\hbar\omega/c^{2}}=\gamma c^{2}.$ (C.88)
Therefore:
$\boxed{\sigma=c\sqrt{\gamma}}$ (C.89)
Both $\hbar$ and $\omega$ cancel identically. The diffusion coefficient of a massless quantum field depends only on the speed of propagation $c$ and the coupling rate $\gamma$.
Proof.
Algebraic. Confirmed numerically by C9 (QED photon validation): at fixed $\gamma=10^{11}\;\mathrm{s}^{-1}$, $\sigma$ was recovered to $1\%$ error across frequencies spanning $\omega\in[10^{12},10^{13}]\;\mathrm{rad/s}$ (a factor of 10), with no frequency dependence observed. ∎
Remark C.3 (Role of Each Constant in C9).
Theorem C.4 reveals the precise role of each constant:
  • $\hbar$: enters through quantisation ($E=\hbar\omega$, $m_{\mathrm{eff}}=\hbar\omega/c^{2}$), then cancels. It governs the quantum nature of the field but does not determine $\sigma$.
  • $c$: appears directly in $\sigma=c\sqrt{\gamma}$. It determines the propagation speed of fluctuations.
  • $\gamma$: the coupling rate to the environment. In C9 (QED), this is external (cavity losses, etc.) and is a free parameter.
C.5 Application to Gravity
Linearised metric perturbations are massless spin-2 fluctuations, with dispersion relation $\omega=c|k|$. By Theorem C.4, the diffusion coefficient for a metric fluctuation mode is therefore:
$\sigma_{\mathrm{grav}}=c\sqrt{\gamma_{\mathrm{grav}}},$ (C.90)
where $\gamma_{\mathrm{grav}}$ is the gravitational self-coupling rate.
The remaining unknown is $\gamma_{\mathrm{grav}}$, whose determination is addressed below.
Observation C.2 (The Uniqueness of Gravity).
In every domain C2–C9, $\gamma$ is an external parameter set by the system’s coupling to its environment: cavity loss rate (C9), medium viscosity (C2), nuclear decay probability (C3), etc. Gravity is structurally different: the gravitational field couples to all energy-momentum, including its own. There is no “external bath”; gravity is its own bath. The coupling rate $\gamma_{\mathrm{grav}}$ must therefore be determined by the fundamental constants of gravity alone.
Lemma C.5 (Unique Rate from Fundamental Constants).
The only combination of $\{\hbar,G,c\}$ with dimensions of a rate $T^{-1}]$ is:
$\gamma_{\mathrm{grav}}=\alpha\,\frac{1}{t_{\mathrm{P}}}=\alpha\,\sqrt{\frac{c^{5}}{\hbar G}},$ (C.91)
where $\alpha>0$ is a dimensionless constant and $t_{\mathrm{P}}=\sqrt{\hbar G/c^{5}}\approx 5.391\times 10^{-44}\;\mathrm{s}$ is the Planck time.
Proof.
Seek $\hbar^{a}G^{b}c^{d}$ with dimensions $T^{-1}$, i.e., $M^{0}L^{0}T^{-1}$:
$M:$ $\quad a-b=0,$ (C.92)
$L:$ $\quad 2a+3b+d=0,$ (C.93)
$T:$ $\quad-a-2b-d=-1.$ (C.94)
From the first equation, $a=b$. Substituting into the second: $5a+d=0$, so d=5ad=-5a. Substituting into the third: a2a+5a=1-a-2a+5a=-1, giving 2a=12a=-1, hence a=1/2a=-1/2. The unique solution is a=b=1/2a=b=-1/2, $d=5/2$:
$\hbar^{-1/2}G^{-1/2}c^{5/2}=\sqrt{\frac{c^{5}}{\hbar G}}=\frac{1}{t_{\mathrm{P}}}.$ (C.95)
Since three equations in three unknowns have a unique solution, no other independent combination exists. ∎
Remark C.4 (Why $\hbar$ Reappears).
A role reversal occurs. In the QHO master formula (Lemma C.2), $\hbar$ enters through quantisation and cancels from $\sigma$ for massless fields (Theorem C.4). But $\hbar$ re-enters through the self-coupling rate $\gamma_{\mathrm{grav}}=1/t_{\mathrm{P}}$, which requires $\hbar$ to set the energy scale at which gravitational self-coupling becomes non-perturbative.
The three constants thus play distinct physical roles:
Constant Enters through Physical role
$c$ $\sigma=c\sqrt{\gamma}$ Speed of fluctuation propagation
$G$ $\gamma_{\mathrm{grav}}\propto 1/\sqrt{G}$ Strength of gravitational self-coupling
$\hbar$ $\gamma_{\mathrm{grav}}\propto 1/\sqrt{\hbar}$ Scale where coupling becomes non-perturbative
C.6 The Main Theorem
Theorem C.6 (The Gravitational Diffusion Theorem).
Let $X$ be any field variable with dimensions $X]=L^{p}T^{q}M^{r}$ describing gravitational degrees of freedom, and let $\sigma$ be its diffusion coefficient in the sense of Definition C.1. Given only $\{\hbar,G,c\}$, then:
$\boxed{\sigma=\frac{X_{\mathrm{P}}}{\sqrt{t_{\mathrm{P}}}}}$ (C.96)
where $X_{\mathrm{P}}$ is the Planck unit of $X$, the unique combination of $\{\hbar,G,c\}$ with dimensions $[X]$.
In Planck units ($\hbar=G=c=1$):
$\tilde{\sigma}=1.$ (C.97)
Proof.
Combining Theorem C.4 and Lemma C.5 with $\alpha=1$ (established by three independent arguments in §C.15):
Step 1. Since metric perturbations are massless, $\sigma=c\sqrt{\gamma_{\mathrm{grav}}}$ (Theorem C.4).
Step 2. The self-coupling rate is $\gamma_{\mathrm{grav}}=1/t_{\mathrm{P}}$ (Lemma C.5 with $\alpha=1$).
Step 3 (velocity variable). For a velocity-like variable ($X]=LT^{-1}$):
$\sigma_{v}=c\cdot\frac{1}{\sqrt{t_{\mathrm{P}}}}=\frac{c}{\sqrt{t_{\mathrm{P}}}}=\frac{X_{\mathrm{P}}^{(\mathrm{vel})}}{\sqrt{t_{\mathrm{P}}}},$ (C.98)
since the Planck velocity is $c$.
Step 4 (arbitrary variable). By Definition C.2, $\sigma]=[X]\cdot T^{-1/2}$. Seek $\hbar^{a}G^{b}c^{d}$ with these dimensions. This is a system of three linear equations in three unknowns $(a,b,d)$:
$M:$ $\quad a-b=r,$ (C.99)
$L:$ $\quad 2a+3b+d=p,$ (C.100)
$T:$ $\quad-a-2b-d=q-\tfrac{1}{2}.$ (C.101)
This system has a unique solution for any $(p,q,r)$, since the coefficient matrix
$A=\begin{pmatrix}1&-1&0\\ 2&3&1\\ -1&-2&-1\end{pmatrix}$ (C.102)
has $\det A=-2\neq 0$ (verified by cofactor expansion along the first row), hence a unique solution exists for any $(p,q,r)$.
This is confirmed by direct construction. The Planck unit $X_{\mathrm{P}}$ satisfies (C.99)–(C.101) with the right-hand side $(r,p,q)$ instead of (r,p,q1/2)(r,p,q-1/2). Dividing by $\sqrt{t_{\mathrm{P}}}$ (which has dimensions $T^{1/2}$) shifts only the $T$-exponent by 1/2-1/2. Therefore $X_{\mathrm{P}}/\sqrt{t_{\mathrm{P}}}$ has exactly dimensions $X]\cdot T^{-1/2}=[\sigma]$.
Since both $X_{\mathrm{P}}$ and $\sqrt{t_{\mathrm{P}}}$ are uniquely determined by $\{\hbar,G,c\}$, the ratio is unique.
Step 5 (Planck units). Setting $\hbar=G=c=1$, all Planck units equal 1 and $t_{\mathrm{P}}=1$, giving $\tilde{\sigma}=1/\sqrt{1}=1$. ∎
Corollary C.7 (Specific Cases).
$\text{Metric perturbation}\;h_{\mu\nu}\;([X]=1):\quad$ $\sigma_{h}=\frac{1}{\sqrt{t_{\mathrm{P}}}}\approx 4.31\times 10^{21}\;\mathrm{s}^{-1/2}$ (C.103)
$\text{Proper distance}\;\delta L\;([X]=L):\quad$ $\sigma_{L}=\frac{\ell_{\mathrm{P}}}{\sqrt{t_{\mathrm{P}}}}\approx 6.96\times 10^{-14}\;\mathrm{m\cdot s}^{-1/2}$ (C.104)
$\text{Velocity}\;v\;([X]=LT^{-1}):\quad$ $\sigma_{v}=\frac{c}{\sqrt{t_{\mathrm{P}}}}\approx 1.29\times 10^{30}\;\mathrm{m\cdot s}^{-3/2}$ (C.105)
C.7 Physical Content
Theorem C.8 (Planck-Scale Fluctuations).
The RMS fluctuation of proper distance accumulated over one Planck time is exactly one Planck length:
$\sqrt{\langle(\delta L)^{2}\rangle}\bigg|_{t=t_{\mathrm{P}}}=\sigma_{L}\cdot\sqrt{t_{\mathrm{P}}}=\ell_{\mathrm{P}}.$ (C.106)
Proof.
For a Wiener process with diffusion coefficient $\sigma_{L}$, $\langle(\delta L)^{2}\rangle=\sigma_{L}^{2}\cdot t$. At $t=t_{\mathrm{P}}$:
$\sigma_{L}^{2}\cdot t_{\mathrm{P}}=\frac{\ell_{\mathrm{P}}^{2}}{t_{\mathrm{P}}}\cdot t_{\mathrm{P}}=\ell_{\mathrm{P}}^{2}.$ (C.107)
Algebraic verification: $\sigma_{L}=(\hbar G/c)^{1/4}$, $t_{\mathrm{P}}=(\hbar G/c^{5})^{1/2}$, so
$\sigma_{L}\cdot\sqrt{t_{\mathrm{P}}}=(\hbar G/c)^{1/4}\cdot(\hbar G/c^{5})^{1/4}=\left(\frac{(\hbar G)^{2}}{c^{6}}\right)^{1/4}=\left(\frac{\hbar G}{c^{3}}\right)^{1/2}=\ell_{\mathrm{P}}.\qed$ (C.108)
Theorem C.9 (Uniqueness of Gravity).
Gravity is the only fundamental interaction for which the diffusion coefficient $\sigma$ is a pure constant of nature without parameters fitted to the predicted data.
Proof.
By Theorem C.4, $\sigma=c\sqrt{\gamma}$ for any massless gauge field. The determining factor is whether $\gamma$ is fixed by fundamental constants alone.
Electromagnetism (QED): The coupling constant is the fine structure constant $\alpha=e^{2}/(4\pi\epsilon_{0}\hbar c)\approx 1/137$, which is dimensionless. A dimensionless constant cannot fix a rate $\gamma]=T^{-1}$ without an additional dimensional scale (particle mass, frequency, or environment parameter). Therefore $\gamma_{\mathrm{QED}}$ is always system-dependent.
Strong force (QCD): The coupling constant $\alpha_{s}$ is likewise dimensionless (at any given scale), with the same consequence.
Weak force: The Fermi constant $G_{F}\approx 1.166\times 10^{-5}\;\mathrm{GeV}^{-2}$ has dimensions, but the weak interaction is massive (mediated by $W^{\pm}$, $Z^{0}$), so the massless cancellation theorem does not apply.
Gravity: The coupling constant $G$ has dimensions $G]=L^{3}M^{-1}T^{-2}$. Combined with $\hbar$ and $c$, this uniquely determines a rate: $\gamma_{\mathrm{grav}}=1/t_{\mathrm{P}}$ (Lemma C.5). No external parameter is needed.
Therefore, within the assumptions of the massless cancellation theorem, gravity is the only interaction whose self-coupling rate: and hence $\sigma$, is determined without external input. ∎
C.8 Consistency Check: Dimensional Analysis
As an independent verification, pure dimensional analysis (without the C8/C9 structural pattern) gives the same result.
Proposition C.10 (Dimensional Uniqueness).
For any field variable $X$ with $X]=L^{p}T^{q}M^{r}$, the unique combination of $\{\hbar,G,c\}$ with dimensions $\sigma]=[X]\cdot T^{-1/2}$ is $\sigma=X_{\mathrm{P}}/\sqrt{t_{\mathrm{P}}}$.
Proof.
This was established as Step 4 of Theorem C.6. The system (C.99)–(C.101) has three equations and three unknowns with a non-singular coefficient matrix (verified: $\det A=-2\neq 0$; the matrix AA maps $\{\hbar,G,c\}$ exponents to $\{M,L,T\}$ powers), hence a unique solution exists for every choice of $(p,q,r)$. ∎
Remark C.5.
The dimensional argument and the structural argument (C8 $\to$ C9 $\to$ QG) converge to the same result. This is not tautological: the dimensional argument requires only the assumption that $\sigma$ is built from $\{\hbar,G,c\}$, while the structural argument derives why these three constants, and no others, determine $\sigma$, because gravity is a massless self-coupled quantum field.
C.9 Four Independent Arguments for $\sigma=\ell_{\mathrm{P}}/\sqrt{t_{\mathrm{P}}}$
The result $\sigma=X_{\mathrm{P}}/\sqrt{t_{\mathrm{P}}}$ is supported by four independent lines of reasoning.
  1. 1.
    Structural Argument (from C8/C9): The validated pattern $\sigma=c\sqrt{\gamma}$ for massless fields, combined with gravitational self-coupling $\gamma_{\mathrm{grav}}=1/t_{\mathrm{P}}$, gives $\sigma_{v}=c/\sqrt{t_{\mathrm{P}}}$. For proper distance: $\sigma_{L}=\ell_{\mathrm{P}}/\sqrt{t_{\mathrm{P}}}$. RMS fluctuation over $t_{\mathrm{P}}$: exactly $\ell_{\mathrm{P}}$.
  2. 2.
    Dimensional Argument: The unique combination of $\{\hbar,G,c\}$ with dimensions $\sigma]=[X]\cdot T^{-1/2}$ is $X_{\mathrm{P}}/\sqrt{t_{\mathrm{P}}}$, regardless of the field variable’s dimensions. In Planck units, $\tilde{\sigma}=1$.
  3. 3.
    Equivalence Principle Argument: If gravity is geometry, then quantum gravity is quantum geometry. Metric fluctuations must exist, and their amplitude must be set by the constants of quantum gravity. The structural and dimensional arguments then fix the result.
  4. 4.
    Observational Argument: All known quantum gravitational effects (Hawking temperature $T_{H}\propto\hbar c^{3}/(GMk_{B})$ [58], Bekenstein entropy $S=A/(4\ell_{\mathrm{P}}^{2})$ [59], holographic bound) involve $\ell_{\mathrm{P}}$ as the fundamental scale, consistent with $\sigma=\ell_{\mathrm{P}}/\sqrt{t_{\mathrm{P}}}$ setting the characteristic fluctuation scale.
C.10 On the Dimensionless Constant $\alpha$
Remark C.6 (The $\alpha=1$ Question).
Lemma C.5 determines $\gamma_{\mathrm{grav}}$ up to a dimensionless constant $\alpha$. The rigorous proof that $\alpha=1$ is given in §C.15; the three independent arguments are:
  1. 1.
    Metric self-consistency: $\langle h^{2}\rangle=\alpha^{2}$ over one Planck time. For $\alpha>1$, the perturbation exceeds the background metric and the theory is mathematically inconsistent. Hard bound: $\alpha\leq 1$.
  2. 2.
    Critical damping: The metric self-coupling rate $\gamma(\omega)=t_{\mathrm{P}}^{2}\omega^{3}$ equals the oscillation frequency at $\omega_{*}=1/t_{\mathrm{P}}$, giving $\alpha=1$ exactly. This is the boundary between coherent propagation and quantum foam.
  3. 3.
    Singularity resolution: The curvature fluctuation must reach the Planck curvature at the minimum scale for singularity regularisation. This requires $\alpha\geq 1$.
Together: $\alpha\leq 1$ (hard) and $\alpha=1$ (critical damping) and $\alpha\geq 1$ (physical) give $\alpha=1$ uniquely.
The Hawking temperature and Bekenstein–Hawking entropy are $\alpha$-independent (§C.15, Theorem C.20) and are therefore consequences of the theory, not inputs to the determination of $\alpha$.
C.11 Summary of the Chain
The logical chain from C1–C9 to quantum gravity is:
$\underbrace{\text{C8: }\sigma^{2}=\frac{\gamma\hbar\omega}{m}}_{\text{QHO master formula}}\xrightarrow{m\to\hbar\omega/c^{2}}\underbrace{\text{C9: }\sigma=c\sqrt{\gamma}}_{\text{massless cancellation}}\xrightarrow{\gamma\to 1/t_{\mathrm{P}}}\underbrace{\text{QG: }\sigma=c/\sqrt{t_{\mathrm{P}}}}_{\text{gravitational self-coupling}}$
Equivalently, in Planck units: $\tilde{\sigma}=1$.
The derivation rests on three ingredients, each validated:
  1. 1.
    The fluctuation-dissipation theorem at $T=0$ (Lemma C.2, validated by C8).
  2. 2.
    The cancellation of $\hbar$ and $\omega$ for massless fields (Theorem C.4, validated by C9).
  3. 3.
    The uniqueness of the gravitational self-coupling rate from $\{\hbar,G,c\}$ (Lemma C.5, dimensional analysis).
No assumption is made about the dimensions of $\sigma$. The result follows from the physics of massless self-coupled quantum fields.
Rigorous Extensions
The preceding sections established the main result. The following sections provide rigorous foundations for each step of the derivation.
C.12 The Quantum Fluctuation-Dissipation Theorem
The master formula (Lemma C.2) is now placed on rigorous footing.
Theorem C.11 (Callen–Welton Theorem, Specialised to a Single Mode).
Consider a quantum harmonic oscillator of frequency $\omega$ and mass $m$, linearly coupled to a heat bath at temperature $T$ with coupling rate $\gamma$. The spectral density of the fluctuating force on the oscillator is:
$S_{F}(\omega^{\prime})=2m\gamma\cdot\hbar\omega^{\prime}\coth\!\left(\frac{\hbar\omega^{\prime}}{2k_{B}T}\right),$ (C.109)
which in the zero-temperature limit $T\to 0$ reduces to:
$S_{F}(\omega^{\prime})\big|_{T=0}=2m\gamma\hbar|\omega^{\prime}|.$ (C.110)
Proof.
This is a standard result of quantum statistical mechanics [5]; see also the general formalism of Kubo [22]. The derivation proceeds from the Kubo formula relating the response function $\chi(\omega)$ to the equilibrium fluctuation spectrum via:
$S_{F}(\omega)=2\hbar\,\mathrm{Im}\,\chi(\omega)\cdot\left[\bar{n}(\omega)+1\right],$
where $\bar{n}(\omega)=(e^{\hbar\omega/k_{B}T}-1)^{-1}$ is the Bose–Einstein distribution. For a damped oscillator, $\mathrm{Im}\,\chi(\omega)=m\gamma\omega$, and $\bar{n}(\omega)+\tfrac{1}{2}=\tfrac{1}{2}\coth(\hbar\omega/2k_{B}T)$, yielding the stated formula. At $T=0$, $\bar{n}\to 0$ and only the vacuum term survives: $S_{F}=2m\gamma\hbar|\omega^{\prime}|$. ∎
Corollary C.12 (Single-Mode Velocity Diffusion at $T=0$).
For a single mode at its resonant frequency $\omega$, the velocity power spectral density at $T=0$ is:
$S_{v}(\omega)=\frac{S_{F}(\omega)}{m^{2}|\chi(\omega)|^{-2}}\bigg|_{T=0},$ (C.111)
where $\chi(\omega)$ is the oscillator response function. At resonance, the dominant contribution to the integrated velocity variance gives an effective diffusion coefficient:
$\sigma_{v}^{2}=\frac{\gamma\hbar\omega}{m},$ (C.112)
reproducing Lemma C.2 from the microscopic quantum theory. The factor $\gamma$ appears because the oscillator accumulates fluctuation energy from the bath at a rate proportional to the coupling strength.
Remark C.7 (On the Factor of $\gamma$).
The manner in which $\gamma$ enters requires care. In the Caldeira–Leggett model [60], $\gamma$ appears both in the dissipation and in the fluctuation. For a single oscillator mode weakly coupled to a bath, the effective Langevin equation is:
$m\ddot{x}+m\gamma\dot{x}+m\omega^{2}x=\xi(t),\qquad\langle\xi(t)\xi(t^{\prime})\rangle=2m\gamma\hbar\omega\,\delta(t-t^{\prime}).$ (C.113)
Writing this in velocity form $\mathrm{d}v=(-\omega^{2}x-\gamma v)\mathrm{d}t+\sigma\mathrm{d}W$ and matching the noise correlator $\langle\xi\xi^{\prime}\rangle=2m\gamma\hbar\omega\,\delta(t-t^{\prime})$ with the Itô convention $\langle\sigma^{2}\rangle=S_{F}/(2m^{2}\gamma)$ at the resonant frequency yields $\sigma^{2}=\gamma\hbar\omega/m$. C8 validates this result numerically to $0.27\%$ (Table 2).
C.13 Validity of the Metric Perturbation–Photon Analogy
Proposition C.13 (Linearised Metric Perturbations as QHO Modes).
In linearised gravity about a background metric $\bar{g}_{\mu\nu}$, the metric perturbation $h_{\mu\nu}=g_{\mu\nu}-\bar{g}_{\mu\nu}$ satisfies:
$\Box\,\bar{h}_{\mu\nu}=0\qquad\text{(in de Donder gauge)},$ (C.114)
where $\bar{h}_{\mu\nu}=h_{\mu\nu}-\tfrac{1}{2}\bar{g}_{\mu\nu}h$ is the trace-reversed perturbation and $\Box=\bar{g}^{\alpha\beta}\nabla_{\alpha}\nabla_{\beta}$ is the d’Alembertian.
Expanding in Fourier modes:
$\bar{h}_{\mu\nu}(x)=\sum_{\lambda=1}^{2}\int\frac{\mathrm{d}^{3}k}{(2\pi)^{3/2}}\frac{1}{\sqrt{2\omega_{k}}}\left[a_{k\lambda}\,\epsilon_{\mu\nu}^{(\lambda)}(k)\,e^{ikx}+\mathrm{h.c.}\right],$ (C.115)
where $\omega_{k}=c|k|$, $\lambda\in\{+,\times\}$ labels the two physical polarisations, and $\epsilon_{\mu\nu}^{(\lambda)}$ are the transverse-traceless polarisation tensors.
Each mode $(k,\lambda)$ is an independent quantum harmonic oscillator with effective mass:
$m_{\mathrm{eff}}(k)=\frac{\hbar\omega_{k}}{c^{2}}$ (C.116)
and zero-point energy $E_{0}=\tfrac{1}{2}\hbar\omega_{k}$.
Proof.
This is the standard result of linearised quantum gravity, see e.g. Weinberg [52], §10.2, or Wald [53], §4.4. The perturbation field has spin 2, but each polarisation mode individually satisfies the massless Klein–Gordon equation and is quantised as an independent oscillator. The effective mass follows from $E=\hbar\omega=m_{\mathrm{eff}}c^{2}$. ∎
Corollary C.14 (Applicability of Theorem C.4 to Metric Modes).
Since each linearised metric perturbation mode is a massless QHO with $m_{\mathrm{eff}}=\hbar\omega/c^{2}$, Theorem C.4 applies directly: the diffusion coefficient for each mode is $\sigma=c\sqrt{\gamma}$, independent of $\omega$ and $\hbar$.
Remark C.8 (Limitations).
The linearised approximation breaks down when $h_{\mu\nu}\sim O(1)$, i.e. at the Planck scale. This is precisely the regime where $\sigma$ matters. Three observations mitigate this concern:
  1. 1.
    The derivation uses the linearised theory, but the result $\sigma=c\sqrt{\gamma_{\mathrm{grav}}}$ involves only fundamental constants. By analytic continuation, the formula applies beyond the linearised regime.
  2. 2.
    The result $\tilde{\sigma}=1$ in Planck units is a statement about dimensional analysis, not about perturbation theory. It holds whether or not the perturbation expansion converges.
  3. 3.
    This is analogous to Hawking’s derivation of black hole temperature [58]: the calculation uses the semiclassical approximation, but the result $T_{H}=\hbar c^{3}/(8\pi GMk_{B})$ is expected to be exact on grounds of dimensional analysis and thermodynamic consistency.
C.14 The Metric Self-Coupling Rate
The self-coupling rate of metric fluctuation modes is now derived from first principles; the prefactor is shown to be exactly unity.
Remark C.9 (Logical Ordering: Modes Before Particles).
The following logical ordering is essential. The stochastic equation $\mathrm{d}g=G\,\mathrm{d}t+\sigma\,\mathrm{d}W$ introduces metric fluctuations as Fourier modes of the Wiener process $\mathrm{d}W$. These are not gravitons: they are simply oscillatory components of stochastic noise. No assumption is made that gravitons exist as particles.
The logical chain is:
  1. 1.
    Metric fluctuation modes carry energy (from the Planck relation applied to mode energy).
  2. 2.
    This energy gravitates (equivalence principle).
  3. 3.
    The resulting self-interaction fixes $\gamma(\omega)=t_{\mathrm{P}}^{2}\omega^{3}$.
  4. 4.
    Critical damping fixes $\sigma=\ell_{\mathrm{P}}/\sqrt{t_{\mathrm{P}}}$.
  5. 5.
    The Fokker–Planck operator on configuration space then has a discrete spectrum, among whose eigenmodes is a massless spin-2 excitation.
  6. 6.
    That eigenmode is the graviton.
Steps 1–4 require no particle concept. The graviton emerges at step 6 as an output, not an input. This avoids circularity with the particle-emergence programme developed in Supplementary Information 6 (SI-6: Particle Spectrum), where all particles arise as Fokker–Planck eigenmodes on the configuration space of metrics.
C.14.1 The Equivalence Principle and Gravitational Universality
The following property of general relativity is inherited by the superspace diffusion framework via canonical Axiom A3 (Classical Correspondence), which requires the macroscopic limit to satisfy the Einstein field equations.
Proposition C.15 (Equivalence Principle, Inherited from General Relativity).
Gravity couples universally to the stress-energy tensor $T_{\mu\nu}$. Everything with energy and momentum gravitates, including the energy stored in the gravitational field itself. The coupling is characterised by a single constant $G$.
Remark C.10 (Three Roles of $G$).
The constant $G$ appears in three contexts:
  1. 1.
    Newton’s law: $V=Gm_{1}m_{2}/r$  (defines $G$ operationally).
  2. 2.
    Einstein’s equation: $G_{\mu\nu}=(8\pi G/c^{4})\,T_{\mu\nu}$  (generalises Newton to all of spacetime geometry; see e.g. [52], Ch. 7, or [53], Ch. 4).
  3. 3.
    Metric self-coupling: the gravitational field’s own energy density sources curvature, with the same constant $G$.
The equivalence principle guarantees these are the same $G$. The factors of $8\pi$ and $16\pi$ from the Einstein–Hilbert action are geometric factors arising from the trace structure of $T_{\mu\nu}$ and the angular integration in the Newtonian limit; they combine to give exactly $V=Gm_{1}m_{2}/r$ with no residual prefactors.
Remark C.11 (Why Gravity Uniquely Fixes $\gamma$).
In the physical domains C1–C9 (Observation C.1), the coupling rate $\gamma$ is a phenomenological parameter set by the environment: for example, the damping rate of a mechanical oscillator depends on the medium. In the gravitational setting this freedom disappears: the equivalence principle (Proposition C.15) requires gravity to couple universally to all energy-momentum, including its own, while dimensional closure (only $G$, $c$, $\hbar$ are available) forces $\gamma$ to be a universal function of $\omega$ constructible from Planck units alone. The derivation below shows that these two constraints determine $\gamma(\omega)=t_{\mathrm{P}}^{2}\omega^{3}$ uniquely.
C.14.2 Derivation of $\gamma(\omega)$ from First Principles
Theorem C.16 (Metric Fluctuation Self-Coupling Rate).
A metric fluctuation mode of frequency $\omega$ has a self-coupling rate:
$\boxed{\gamma(\omega)=\frac{G\hbar}{c^{5}}\,\omega^{3}=t_{\mathrm{P}}^{2}\,\omega^{3}}$ (C.117)
with prefactor exactly unity. This is not a perturbative estimate; it follows from three exact principles applied to metric fluctuation modes, without assuming gravitons exist as particles.
Proof.
The derivation uses three inputs, each exact:
Step 1: Mode energy (Planck relation, exact).
A metric fluctuation mode of angular frequency $\omega$ carries energy:
$E=\hbar\omega.$ (C.118)
This follows from treating $\mathrm{d}W_{\mu\nu}$ as a superposition of oscillatory modes, each of which is a quantum harmonic oscillator with zero-point energy $\hbar\omega/2$ and excitation quantum $\hbar\omega$.
Step 2: Gravitational self-interaction energy (equivalence principle, exact).
By Lemma C.3, the effective gravitational mass of a mode is $m_{\mathrm{eff}}=E/c^{2}=\hbar\omega/c^{2}$. The gravitational self-energy of a mode localised over its own reduced wavelength $\bar{\lambda}$ is:
$V=\frac{Gm_{\mathrm{eff}}^{2}}{\bar{\lambda}}=\frac{G(\hbar\omega)^{2}}{c^{4}\,\bar{\lambda}}.$ (C.119)
This is Newton’s law applied via the equivalence principle: gravity couples to all energy–momentum, including the energy stored in metric fluctuations. The $G$ here is the same $G$ as in Newton’s law (Remark C.10).
No particle interpretation is required. The calculation concerns the gravitational self-energy of a classical field mode that carries energy $\hbar\omega$.
Step 3: Natural self-interaction scale (reduced wavelength).
The self-coupling of a mode with itself occurs over its own reduced wavelength:
$r=\bar{\lambda}=\frac{c}{\omega}.$ (C.120)
For a massless field, $\bar{\lambda}=\hbar/(m_{\mathrm{eff}}c)=c/\omega$. The choice of $c/\omega$ (not $2\pi c/\omega$) follows the universal quantum-mechanical convention: $\hbar$ (not $h$) serves as the quantum of action, and correspondingly $\bar{\lambda}$ (not $\lambda$) as the natural length. This is the same convention that defines $\ell_{\mathrm{P}}=\sqrt{\hbar G/c^{3}}$ rather than $\sqrt{hG/c^{3}}$.
Combining:
$V=\frac{G(\hbar\omega)^{2}}{c^{4}\cdot c/\omega}=\frac{G\hbar^{2}\omega^{3}}{c^{5}}.$ (C.121)
The gravitational self-energy of a metric fluctuation mode scales as $\omega^{3}$: higher-frequency modes are more strongly self-coupled.
Step 4: Self-coupling rate (quantum energy–time relation).
The natural rate associated with an interaction energy $V$ is:
$\gamma=\frac{V}{\hbar}=\frac{G\hbar\omega^{3}}{c^{5}}=t_{\mathrm{P}}^{2}\,\omega^{3},$ (C.122)
where the identity $G\hbar/c^{5}=t_{\mathrm{P}}^{2}$ has been applied. This is the quantum energy–time relation $\Delta t\sim\hbar/\Delta E$ inverted: an interaction of strength $V$ induces transitions on a timescale $\hbar/V$, corresponding to a rate $V/\hbar$. This is not the Fermi golden rule (which gives $\gamma\propto|V|^{2}/\hbar$ and applies to weak coupling); it is the linear energy–rate correspondence that holds at all coupling strengths. Every step uses an exact physical law. No step assumes the existence of gravitons as particles. ∎
Remark C.12 (Why This Is Not Perturbative).
A perturbative (one-loop) calculation of the graviton self-energy yields the same functional form $\gamma\propto\omega^{3}$, but with a prefactor involving factors from the Einstein–Hilbert action, spin degeneracies, symmetry factors, and $(2\pi)^{n}$ from momentum integrals. However, perturbation theory breaks down at $\omega\sim\omega_{P}$, where the coupling becomes $O(1)$.
The derivation above avoids perturbation theory entirely. It uses:
  • Newton’s law ($V=Gm_{1}m_{2}/r$), which defines $G$;
  • the Planck relation ($E=\hbar\omega$), applied to field modes;
  • mass–energy equivalence ($m=E/c^{2}$), which is exact;
  • the equivalence principle (gravity couples universally to energy), which guarantees the same $G$.
The factors of $16\pi$ in the Einstein–Hilbert action are absorbed into the definition of $G$ via Newton’s law. When one computes the potential between two point sources from the full field equation, all geometric factors combine to give $V=Gm_{1}m_{2}/r$ with the Cavendish $G$, no residual prefactors.
Corollary C.17 (Verification at the Planck Scale).
At $\omega=\omega_{P}=1/t_{\mathrm{P}}$:
  • Mode energy: $E=\hbar\omega_{P}=E_{\mathrm{P}}$ (the Planck energy).
  • Effective mass: $m_{\mathrm{eff}}=E_{\mathrm{P}}/c^{2}=m_{\mathrm{P}}$ (the Planck mass).
  • Self-interaction distance: $r=c/\omega_{P}=ct_{\mathrm{P}}=\ell_{\mathrm{P}}$ (the Planck length).
  • Potential: $V=Gm_{\mathrm{P}}^{2}/\ell_{\mathrm{P}}=E_{\mathrm{P}}$ (gravitational self-energy equals mode energy).
  • Rate: $\gamma(\omega_{P})=V/\hbar=E_{\mathrm{P}}/\hbar=1/t_{\mathrm{P}}$.
At the Planck frequency, a metric fluctuation mode’s self-interaction energy equals its own energy, and the self-coupling rate equals $1/t_{\mathrm{P}}$. This is the physical definition of the Planck scale: the scale at which gravitational self-coupling becomes $O(1)$.
Theorem C.18 (The Prefactor $\beta$ and the Equivalence Principle).
Let $\gamma(\omega)=\beta\cdot t_{\mathrm{P}}^{2}\omega^{3}$ with $\beta$ a dimensionless prefactor. Assuming the Planck relation $E=\hbar\omega$ and the reduced wavelength convention $\bar{\lambda}=c/\omega$, then $\beta=1$ if and only if the equivalence principle holds for the gravitational field.
Proof.
$(\Rightarrow)$: If $\beta=1$, the self-coupling derivation (Theorem C.16) recovers $V=Gm_{\mathrm{eff}}^{2}/r$ with the same $G$ as Newton’s law. This is a consequence of the equivalence principle applied to the gravitational field itself.
$(\Leftarrow)$: If the equivalence principle holds, the gravitational self-interaction is governed by the same $G$ as all other gravitational interactions. The derivation of Theorem C.16 then gives $\gamma=(G\hbar/c^{5})\omega^{3}=t_{\mathrm{P}}^{2}\omega^{3}$, hence $\beta=1$.
Conversely, if $\beta\neq 1$, then either:
  1. 1.
    $G_{\mathrm{self}}\neq G_{\mathrm{Newton}}$ (the gravitational field couples to itself with a different strength than to matter), violating the equivalence principle; or
  2. 2.
    $E\neq\hbar\omega$ for metric fluctuation modes (violating the Planck relation); or
  3. 3.
    the natural self-interaction scale is not $c/\omega$ (violating the $\hbar$-convention for the reduced wavelength).
Options (1) and (2) contradict established physics. Option (3) contradicts the universal convention $\ell_{\mathrm{P}}=\sqrt{\hbar G/c^{3}}$. ∎
C.14.3 Self-Consistency and the Graviton as Emergent
Theorem C.19 (Frequency-Dependent Diffusion and the Onset of Foam).
The metric self-coupling rate $\gamma(\omega)=t_{\mathrm{P}}^{2}\omega^{3}$ combined with the massless cancellation $\sigma=c\sqrt{\gamma}$ gives a frequency-dependent velocity-variable diffusion coefficient:
$\sigma_{v}(\omega)=c\sqrt{t_{\mathrm{P}}^{2}\omega^{3}}=ct_{\mathrm{P}}\omega^{3/2}.$ (C.123)
For the dimensionless metric perturbation $h_{\mu\nu}$, the corresponding diffusion coefficient is $\sigma_{h}(\omega)=\sigma_{v}(\omega)/c=t_{\mathrm{P}}\omega^{3/2}$.
At $\omega=\omega_{P}=1/t_{\mathrm{P}}$: $\sigma_{h}(\omega_{P})=t_{\mathrm{P}}\cdot t_{\mathrm{P}}^{-3/2}=1/\sqrt{t_{\mathrm{P}}}$, reproducing Theorem C.6.
The RMS metric fluctuation over one oscillation period $\Delta t=2\pi/\omega$ is:
$\sqrt{\langle h^{2}\rangle}=\sigma_{h}\cdot\sqrt{\Delta t}=t_{\mathrm{P}}\omega^{3/2}\cdot\sqrt{\frac{2\pi}{\omega}}=t_{\mathrm{P}}\sqrt{2\pi}\,\omega\sim\frac{\omega}{\omega_{P}}\quad\text{(since }t_{\mathrm{P}}\omega_{P}=1\text{)}.$ (C.124)
The metric fluctuation becomes $O(1)$ precisely at $\omega=\omega_{P}$. Below the Planck frequency, $\langle h^{2}\rangle\sim(\omega/\omega_{P})^{2}\ll 1$: spacetime is smooth.
Remark C.13 (The Graviton Emerges).
With $\sigma=\ell_{\mathrm{P}}/\sqrt{t_{\mathrm{P}}}$ now determined, the Fokker–Planck equation on the configuration space of metrics (superspace) has a well-defined eigenvalue problem. Among its eigenmodes is a massless, transverse-traceless, spin-2 excitation of the metric. This is the graviton: an emergent output of the theory, not an assumed input.
The self-consistency check is that the emergent graviton’s coupling strength is exactly $G$ and its frequency-dependent damping rate is exactly $\gamma=t_{\mathrm{P}}^{2}\omega^{3}$, the same values used in the derivation. This is not circular: the derivation used only the energy content of metric modes and the equivalence principle, both of which are properties of the classical gravitational field. The particle interpretation is a consequence, not a premise.
C.15 Fixing $\alpha=1$: Three Independent Arguments
The following three arguments establish that the dimensionless constant $\alpha$ in $\gamma_{\mathrm{grav}}=\alpha/t_{\mathrm{P}}$ is exactly unity, using arguments that do not invoke the Bekenstein–Hawking entropy.
C.15.1 Why Bekenstein–Hawking Cannot Fix $\alpha$
Theorem C.20 ($\alpha$-Independence of the Hawking Temperature).
The Hawking temperature [58] derived from $\sigma=\alpha\ell_{\mathrm{P}}$ is independent of $\alpha$:
$T_{H}(\alpha)=T_{H}=\frac{\hbar c^{3}}{8\pi GMk_{B}}\qquad\text{for all }\alpha>0.$ (C.125)
Proof.
The stochastic derivation of $T_{H}$ proceeds in three steps:
Step 1. Local fluctuation energy at the horizon. The position uncertainty is $\delta r\sim\sigma=\alpha\ell_{\mathrm{P}}$. By Heisenberg:
$E_{\mathrm{local}}=\frac{\hbar c}{\delta r}=\frac{\hbar c}{\alpha\ell_{\mathrm{P}}}=\frac{E_{P}}{\alpha}.$ (C.126)
Step 2. Gravitational redshift. The energy ratio from proper distance $\sim\delta r$ above the horizon to infinity is, in the near-horizon Schwarzschild geometry:
$\frac{E_{\infty}}{E_{\mathrm{local}}}\sim\frac{\delta r}{r_{s}},$ (C.127)
where the proportionality constant depends on the precise definition of $\delta r$ (coordinate vs. proper distance) but is $\alpha$-independent.
Step 3. Energy at infinity:
$E_{\infty}\sim E_{\mathrm{local}}\times\frac{\delta r}{r_{s}}=\frac{E_{P}}{\alpha}\times\frac{\alpha\ell_{\mathrm{P}}}{r_{s}}=\frac{E_{P}\ell_{\mathrm{P}}}{r_{s}}.$ (C.128)
The factors of $\alpha$ cancel identically: $E_{\mathrm{local}}\propto 1/\alpha$ and $\delta r\propto\alpha$. The thermal identification $T\sim E_{\infty}/k_{B}$ then gives $T_{H}=\hbar c^{3}/(8\pi GMk_{B})$, where the numerical prefactor $1/(8\pi)$ is fixed by thermodynamic consistency with the first law, independent of $\alpha$. ∎
Corollary C.21 ($\alpha$-Independence of Bekenstein–Hawking Entropy).
The Bekenstein–Hawking entropy [59, 58] $S_{BH}=k_{B}A/(4\ell_{\mathrm{P}}^{2})$ is also independent of $\alpha$, since it follows from the first law of black hole mechanics [61] $\mathrm{d}S=\mathrm{d}M/T_{H}$ via integration, and $T_{H}$ is $\alpha$-independent.
Remark C.14.
The $\alpha$-independence of both $T_{H}$ and $S_{BH}$ means that black hole thermodynamics cannot determine $\alpha$; an independent argument is required. The three arguments below provide exactly this.
C.15.2 Argument 1: Metric Self-Consistency (Upper Bound)
Theorem C.22 (Hard Upper Bound: $\alpha\leq 1$).
For the stochastic metric theory to be internally consistent, $\alpha\leq 1$.
Proof.
Consider the dimensionless metric perturbation $h_{\mu\nu}$. With $\sigma_{h}=\alpha/\sqrt{t_{\mathrm{P}}}$, the RMS fluctuation accumulated over one Planck time is:
$\langle h^{2}\rangle\big|_{t=t_{\mathrm{P}}}=\sigma_{h}^{2}\cdot t_{\mathrm{P}}=\frac{\alpha^{2}}{t_{\mathrm{P}}}\cdot t_{\mathrm{P}}=\alpha^{2}.$ (C.129)
The metric is $g_{\mu\nu}=\bar{g}_{\mu\nu}+h_{\mu\nu}$, where $\bar{g}$ is the background. For the perturbative decomposition to be meaningful, we require $|h|<|\bar{g}|$, i.e., $\langle h^{2}\rangle\lesssim O(1)$.
For $\alpha>1$: $\langle h^{2}\rangle=\alpha^{2}>1$. The perturbation exceeds the background metric over a single Planck time. The linearised framework destroys its own geometrical foundation. The theory is inconsistent.
Therefore $\alpha\leq 1$. This is a hard bound: not an approximation or physical assumption, but a mathematical requirement for internal consistency. ∎
C.15.3 Argument 2: The Critical Damping Theorem (Exact Value)
Theorem C.23 (Critical Damping at the Planck Scale).
The metric self-coupling rate $\gamma(\omega)=t_{\mathrm{P}}^{2}\omega^{3}$ (Theorem C.16) equals the oscillation frequency $\omega$ at exactly one frequency:
$\gamma(\omega_{*})=\omega_{*}\quad\Longleftrightarrow\quad\omega_{*}=\frac{1}{t_{\mathrm{P}}}=\omega_{P}.$ (C.130)
This is the critical damping condition. The corresponding diffusion coefficient is:
$\sigma(\omega_{P})=c\sqrt{\gamma(\omega_{P})}=c\sqrt{1/t_{\mathrm{P}}}=\frac{c}{\sqrt{t_{\mathrm{P}}}},$ (C.131)
which corresponds to $\alpha=1$.
Proof.
The critical damping condition $\gamma(\omega_{*})=\omega_{*}$ gives:
$t_{\mathrm{P}}^{2}\omega_{*}^{3}=\omega_{*}\qquad\implies\qquad\omega_{*}^{2}=\frac{1}{t_{\mathrm{P}}^{2}}\qquad\implies\qquad\omega_{*}=\frac{1}{t_{\mathrm{P}}}=\omega_{P}.$ (C.132)
This is exact (algebraic identity, no approximation). At this frequency, $\gamma(\omega_{P})=1/t_{\mathrm{P}}$, and Theorem C.4 gives $\sigma=c/\sqrt{t_{\mathrm{P}}}$, i.e., $\alpha=1$. ∎
Remark C.15 (Physical Interpretation).
The critical damping condition divides metric fluctuation modes into three regimes:
Regime Condition Physics
Underdamped $\omega<\omega_{P}$ Modes propagate coherently
Critically damped $\omega=\omega_{P}$ Boundary: onset of quantum foam
Overdamped $\omega>\omega_{P}$ Modes cannot propagate; pure diffusion
The Planck frequency is the unique frequency at which a metric mode’s self-coupling damping rate exactly matches its oscillation rate. The selection of this boundary as the physically relevant scale is motivated by the transition from coherent propagation to incoherent diffusion, but it is not logically mandatory: Arguments 1 and 3 independently force $\alpha\leq 1$ and $\alpha\geq 1$, so the bracket closes at $\alpha=1$ regardless of whether the critical damping interpretation is accepted. Argument 2 provides an independent physical route to the same conclusion and identifies $\omega_{P}$ as the natural boundary scale.
C.15.4 Argument 3: Singularity Resolution (Lower Bound)
Theorem C.24 (Lower Bound: $\alpha\geq 1$).
For metric fluctuations to resolve classical singularities, $\alpha\geq 1$.
Proof.
The Kretschmann curvature scalar $K=R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}$ has dimensions $K]=L^{-4}$. A dimensionless metric perturbation $h$ varying over length scale $L$ induces curvature fluctuations:
$\delta K\sim\frac{h^{2}}{L^{4}}.$ (C.133)
The Planck curvature is $K_{P}=1/\ell_{\mathrm{P}}^{4}$, the scale at which quantum gravity becomes dominant.
The RMS dimensionless metric perturbation accumulated over one Planck time is $h_{\mathrm{rms}}=\sigma_{h}\sqrt{t_{\mathrm{P}}}=\alpha$ (from $\sigma_{h}=\alpha/\sqrt{t_{\mathrm{P}}}$). At the Planck scale $L=\ell_{\mathrm{P}}$:
$\delta K\big|_{L=\ell_{\mathrm{P}}}=\frac{\alpha^{2}}{\ell_{\mathrm{P}}^{4}}=\alpha^{2}K_{P}.$ (C.134)
For singularity resolution, the fluctuation-induced curvature at $L=\ell_{\mathrm{P}}$ must reach the Planck curvature: $\delta K\geq K_{P}$. This requires $\alpha^{2}\geq 1$, hence $\alpha\geq 1$.
Together with Theorem C.22 ($\alpha\leq 1$): the only value satisfying both $\alpha\geq 1$ and $\alpha\leq 1$ is $\alpha=1$. ∎
C.15.5 Convergence of the Three Arguments
Corollary C.25 ($\alpha=1$: Three Independent Routes).
The three arguments converge:
  1. 1.
    Metric self-consistency: $\alpha\leq 1$ (hard bound; Theorem C.22).
  2. 2.
    Critical damping: $\alpha=1$ (exact; Theorem C.23).
  3. 3.
    Singularity resolution: $\alpha\geq 1$ (physical; Theorem C.24).
None invokes the Bekenstein–Hawking entropy or Hawking temperature. The result $\alpha=1$ is derived from metric self-coupling dynamics alone.
The Hawking temperature and Bekenstein–Hawking entropy can now be derived as consequences of $\sigma=\ell_{\mathrm{P}}/\sqrt{t_{\mathrm{P}}}$, rather than assumed as inputs.
Remark C.16 (Robustness of $\gamma(\omega)=t_{P}^{2}\omega^{3}$ to the microscopic prescription).
The derivation of the self-coupling rate $\gamma(\omega)=t_{P}^{2}\omega^{3}$ (Theorem C.14) uses the linear energy-rate relation $\gamma=V/\hbar$. The Fermi golden rule (FGR) gives $\gamma_{\mathrm{FGR}}=2\pi|V|^{2}/(\hbar^{2}\rho_{f})$, which might appear to give a different result. However, both prescriptions produce the same $\omega^{3}$ scaling.
Point 1: identical functional form. The self-energy is $V\propto\omega^{3}$ and the graviton density of states scales as $\rho_{f}\propto\omega^{3}$. The FGR therefore gives $\gamma_{\mathrm{FGR}}\propto(\omega^{3})^{2}/\omega^{3}=\omega^{3}$, identical to the linear result. The two prescriptions differ only in an $O(1)$ prefactor.
Point 2: the linear relation is exact for self-coupling. The FGR applies to a system weakly coupled to a dense bath of final states. Gravitational self-coupling is structurally a Rabi-type problem: a single mode at frequency $\omega$ couples to its own gravitational field at the same frequency, with coupling $V=Gm_{\mathrm{eff}}^{2}/\bar{\lambda}$. For a two-level system, the exact transition rate is $V/\hbar$, not $V^{2}/(\hbar^{2}\rho_{f})$.
Point 3: independent confirmation from the perturbative mode sum. A perturbative computation of the Mori–Zwanzig memory kernel using the standard $|V|^{2}$ prescription yields $\gamma_{\mathrm{eff}}\propto\omega^{3}$ mode by mode, confirming that the functional form is not an artefact of the linear prescription.
Point 4: the prefactor $\alpha=1$ is fixed independently. The three convergent arguments above (metric self-consistency, critical damping, singularity resolution) operate at the level of the effective rate $\gamma(\omega_{P})$ and its macroscopic consequences: not at the level of the microscopic scattering calculation. They yield $\alpha=1$ for any prescription producing $\gamma\propto t_{P}^{2}\omega^{3}$.
The choice between $V/\hbar$ and $2\pi|V|^{2}/(\hbar^{2}\rho_{f})$ is therefore a question about an $O(1)$ prefactor within a scaling relation that both prescriptions produce identically. No physical prediction of the framework depends on this choice.
Remark C.17 (Independence of the three $\alpha=1$ arguments).
All three arguments assume Axiom A2 (the fundamental equation is valid at the Planck scale); this is by construction. However, they do not uniformly assume perturbative validity. The metric self-consistency bound ($\alpha\leq 1$) defines the boundary of perturbative validity: $\alpha>1$ means the perturbation exceeds the background, a mathematical inconsistency in the decomposition $g=\bar{g}+h$ regardless of whether perturbation theory is trusted. The critical damping argument uses $\gamma(\omega)=t_{P}^{2}\omega^{3}$, whose $\omega^{3}$ scaling is confirmed by both perturbative and non-perturbative routes (Remark C.16). The singularity resolution argument ($\alpha\geq 1$) is a physical requirement on any theory of quantum gravity: that Planck-scale effects resolve classical singularities: not an assumption about perturbative validity. The three arguments share A2 but differ in their additional inputs: mathematical consistency, self-coupling dynamics, and physical adequacy, respectively.
C.16 Dimensional Consistency
The following constraint is a consequence of requiring $\sigma$ to be constructed from $\{\hbar,G,c\}$ alone, as encoded in canonical Axiom A2 (Stochastic Evolution), which specifies $\sigma=\ell_{\mathrm{P}}=\sqrt{\hbar G/c^{3}}$.
Proposition C.26 (Dimensional Consistency).
The fluctuation amplitude $\sigma$ is constructed solely from the fundamental constants of quantum gravity $\{\hbar,G,c\}$. Its dimensions are determined by the choice of field variable via $\sigma]=[X]\cdot T^{-1/2}$.
Remark C.18.
When $X$ is a proper-distance variable, $\sigma]=L\cdot T^{-1/2}$ and the RMS fluctuation over $t_{\mathrm{P}}$ is $\sigma\sqrt{t_{\mathrm{P}}}=\ell_{\mathrm{P}}$. The Planck length is not assumed; it emerges as a consequence of Theorems C.4 and C.5.
C.17 Complete Logical Chain
The complete chain from physical principles to the final result is stated below, identifying every input and every step.
Inputs:
  1. I1.
    Gravity is a quantum field (standard expectation; no consistent coupling of quantum matter to classical gravity exists).
  2. I2.
    Linearised metric perturbations are massless spin-2 with $\omega=c|k|$ (from linearised GR).
  3. I3.
    The fluctuation-dissipation theorem holds at $T=0$ (Callen–Welton, Theorem C.11).
  4. I4.
    The equivalence principle: gravity couples universally to energy–momentum with coupling constant $G$ (Proposition C.15).
  5. I5.
    The metric perturbation must satisfy $\langle h^{2}\rangle\leq 1$ at the fluctuation scale (mathematical consistency).
Derivation:
I3 $\implies\sigma^{2}=\gamma\hbar\omega/m$ (Lemma C.2: QHO master formula) (C.135)
$\text{I2}+\text{above}$ $\implies\sigma=c\sqrt{\gamma}$ (Theorem C.4: massless cancellation) (C.136)
I4 $\implies\gamma(\omega)=t_{\mathrm{P}}^{2}\omega^{3}\;\text{(exact, }\beta=1\text{)}$ (Theorem C.16: equivalence principle) (C.137)
[…]
I5 $\implies\alpha\leq 1;\;\text{crit.\ damp.}\implies\alpha=1$ (Theorems C.22C.23) (C.139)
[…]
Consequences (derived, not assumed):
$T_{H}$ $=\frac{\hbar c^{3}}{8\pi GMk_{B}}$ $\text{(exact; }\alpha\text{-independent)}$ (C.141)
$S_{BH}$ $=\frac{k_{B}A}{4\ell_{\mathrm{P}}^{2}}$ $\text{(from first law }\mathrm{d}S=\mathrm{d}M/T_{H}\text{)}$ (C.142)
Physical consequence:
$\Delta L_{\mathrm{rms}}=\sigma_{L}\sqrt{t_{\mathrm{P}}}=\frac{\ell_{\mathrm{P}}}{\sqrt{t_{\mathrm{P}}}}\cdot\sqrt{t_{\mathrm{P}}}=\ell_{\mathrm{P}}.$ (C.143)
Spacetime fluctuates by one Planck length per Planck time. In Planck units, $\tilde{\sigma}=1$. $\blacksquare$
Constant Where it enters Physical role
$c$ $\sigma=c\sqrt{\gamma}$ Speed of fluctuation propagation; enters through the massless dispersion relation
$G$ $\gamma_{\mathrm{grav}}\propto G^{-1/2}$ Strength of gravitational self-coupling; the only dimensionful gauge coupling in the Standard Model + gravity
$\hbar$ $\gamma_{\mathrm{grav}}\propto\hbar^{-1/2}$ Sets the energy scale at which self-coupling becomes non-perturbative; cancels from single-mode physics but re-enters through the self-coupling rate
Canonical Axiom A2 of the superspace diffusion framework encapsulates this result: $\mathrm{d}g_{ij}=\mathcal{D}_{ij}\,\mathrm{d}\tau+\ell_{\mathrm{P}}\,\mathrm{d}W_{ij}$. The value $\sigma=\ell_{\mathrm{P}}$ is not an independent postulate but the unique output of the derivation chain above.
C.18 Status of Each Logical Step
For transparency, every step of the derivation is classified by its epistemic status.
Step Status Evidence / Assumptions
$\sigma^{2}=\gamma\hbar\omega/m$ (master formula) Proven Postulate C.1; validated numerically by C8 ($R^{2}=0.9999$)
$\sigma=c\sqrt{\gamma}$ (massless cancellation) Proven Algebraic from Lemma C.3; validated by C9 (7 frequencies, all $1\%$)
Metric perturbations are massless QHO modes Standard Postulate C.2; see Lemma C.3
$\gamma(\omega)=t_{\mathrm{P}}^{2}\omega^{3}$ (self-coupling) Derived From Postulate C.3; prefactor $\beta=1$ exact (Theorem C.16)
$\alpha\leq 1$ (upper bound) Proven $\langle h^{2}\rangle=\alpha^{2}$; exceeding 1 destroys metric structure (mathematical consistency)
$\alpha=1$ (critical damping) Proven $\gamma(\omega_{*})=\omega_{*}$ solves to $\omega_{*}=1/t_{\mathrm{P}}$ exactly (algebraic identity)
$\alpha\geq 1$ (lower bound) Physical Singularity resolution requires $\delta K\sim K_{P}$ at the minimum scale
$\alpha$ cancels from $T_{H}$ Proven $E_{\mathrm{local}}\propto 1/\alpha$ and $\delta r\propto\alpha$ cancel identically
$\tilde{\sigma}=1$ in Planck units Proven Follows from dimensional analysis alone, independent of $\alpha$
$\Delta L_{\mathrm{rms}}=\ell_{\mathrm{P}}$ per $t_{\mathrm{P}}$ Proven Algebraic identity $\sigma_{L}\sqrt{t_{\mathrm{P}}}=\ell_{\mathrm{P}}$ (with $\alpha=1$)
The $\alpha=1$ determination rests on three independent arguments: the hard upper bound from metric self-consistency ($\alpha\leq 1$), the critical damping theorem ($\alpha=1$ exactly), and the singularity resolution lower bound ($\alpha\geq 1$). The Bekenstein–Hawking entropy is derived as a consequence.
The strongest elements are the massless cancellation (proven algebraically and validated numerically), the dimensional uniqueness (a theorem of linear algebra), and the critical damping theorem (an algebraic identity). These are not model-dependent.
Appendix D Uniqueness of $\sigma=\ell_{\mathrm{P}}$
Scope and Logical Position
This appendix provides a fully self-contained proof that, within axioms A1–A4, the metric fluctuation amplitude $\sigma$ is uniquely determined to be the Planck length $\ell_{\mathrm{P}}=\sqrt{\hbar G/c^{3}}$. The proof is situated within the conceptual framework of emergent time: the quadratic variation of the stochastic process on superspace. The mathematical characterisation of the noise (§D.4) establishes the uniqueness of the Wiener structure from the Lévy–Khintchine theorem, and connects the Fokker–Planck evolution to both the Wheeler–DeWitt equation and the gravitational path integral.
The logical chain proceeds from the four canonical axioms (A1–A4) through four stages. First, the Lévy–Khintchine theorem establishes that the Wiener process is the unique noise structure consistent with A1 (§D.4). Second, the emergent time theorem establishes that physical time is constructed from the diffusion (not assumed as a background coordinate), with the drift contributing exactly zero by the Itô product rule. Third, the uniqueness theorem proves via three independent arguments (curvature self-consistency ($\alpha\leq 1$), critical damping ($\alpha=1$ exactly), and singularity resolution ($\alpha\geq 1$)) that $\sigma=\ell_{\mathrm{P}}$ is the only consistent fluctuation amplitude. Fourth, the classical limit theorem verifies that general relativity emerges for $L\gg\ell_{\mathrm{P}}$, and the MSR path-integral representation establishes connections to canonical quantum gravity and the Euclidean gravitational path integral (§D.4.2–§D.4.3).
Together, these results establish that the fundamental equation $\mathrm{d}g_{ij}=\mathcal{D}_{ij}\,\mathrm{d}\tau+\ell_{\mathrm{P}}\,\mathrm{d}W_{ij}$ contains non-parametric, first-principles predictions: the amplitude is uniquely determined, the drift and covariance are fixed by symmetry, and time itself is a derived quantity.
All definitions and the complete proof chain are included. The sole external input is the gravitational self-coupling rate $\gamma(\omega)=t_{\mathrm{P}}^{2}\omega^{3}$, derived from the equivalence principle in Appendix C.
Relationship to the canonical axioms. This proof unpacks canonical Axiom A2 of the superspace diffusion framework, which asserts $\mathrm{d}g_{ij}=\mathcal{D}_{ij}\,\mathrm{d}\tau+\ell_{\mathrm{P}}\,\mathrm{d}W_{ij}$. The value $\sigma=\ell_{\mathrm{P}}$ is shown to be the unique output of five physical postulates (P1–P5) that constitute the justification for A2. The four canonical axioms are stated in §D.1.
D.1 The Four Canonical Axioms
The framework rests on four canonical axioms from which all subsequent results are derived. Three (A1–A3) are physical hypotheses about the gravitational degrees of freedom; the fourth (A4) is an interpretive commitment about the relationship between the stochastic formalism and observational reality. Together they introduce no new dimensional parameters beyond $\hbar$, $G$, and $c$.
Axiom D.1 (Configuration Space).
The gravitational degrees of freedom are described by a Riemannian three-metric $g_{ij}(x)$ on a spatial manifold $\Sigma$. The space of all such metrics modulo spatial diffeomorphisms constitutes the gravitational configuration space (Wheeler superspace [30, 31]):
$\mathcal{C}\;=\;\frac{\mathrm{Riem}(\Sigma)}{\mathrm{Diff}(\Sigma)}$ (D.144)
The tangent space $T_{g}\mathcal{C}$ is equipped with the DeWitt supermetric $\mathcal{G}^{ijkl}(g)$ [31].
Axiom D.2 (Stochastic Evolution).
The spatial metric configuration undergoes a stochastic process on $\mathcal{C}$:
$\mathrm{d}g_{ij}(x)=\mathcal{D}_{ij}[g](x)\,\mathrm{d}\tau+\ell_{\mathrm{P}}\,\mathrm{d}W_{ij}(x)$ (D.145)
where:
  • $\mathcal{D}_{ij}[g]$ is a drift functional encoding gravitational dynamics,
  • $\ell_{\mathrm{P}}=\sqrt{\hbar G/c^{3}}$ is the Planck length,
  • $\mathrm{d}W_{ij}(x)$ is a symmetric tensor-valued Wiener process satisfying
    $\mathbb{E}[\mathrm{d}W_{ij}(x)]$ $=0,$ (D.146)
    $\mathbb{E}[\mathrm{d}W_{ij}(x)\,\mathrm{d}W_{kl}(y)]$ $=K_{ijkl}[g](x,y)\,\mathrm{d}\tau,$ (D.147)
    where $K_{ijkl}$ is a positive (semi-)definite covariance operator on symmetric 2-tensors.
The ordering parameter $\tau$ is not a background spacetime coordinate. It labels the filtration $\{\mathcal{F}_{\tau}\}$ generated by the Wiener process. Physical time emerges as a monotone functional of the stochastic evolution (§D.5).
Remark D.1 (Positive-definiteness of $K$).
Axiom D.2 permits $K_{ijkl}$ to be positive semi-definite in its statement, but three independent requirements force strict positive-definiteness: (i) the Wiener process must have non-degenerate increments in every tensorial direction, (ii) the emergent time functional (§D.5) must be strictly monotone, and (iii) the Fokker-Planck generator must be hypoelliptic [62], guaranteeing smooth probability densities for $\tau>0$. Theorem D.12 establishes that $K_{ijkl}$ is a member of the DeWitt family $\mathcal{G}^{ijkl}(\lambda)$ with $\lambda>-1/3$, which is positive-definite on traceless symmetric 2-tensors and on trace modes separately. This fixes the covariance operator uniquely (up to the overall scale $\ell_{\mathrm{P}}^{2}$) and ensures that all three conditions are satisfied.
Axiom D.3 (Classical Correspondence).
In the macroscopic limit ($L\gg\ell_{\mathrm{P}}$), there exists an emergent four-dimensional Lorentzian geometry $(M,g^{(4)})$ obtained by coarse-graining the $\tau$-indexed stochastic evolution on $\mathcal{C}$ such that $g^{(4)}$ satisfies the Einstein field equations [21] to observed accuracy:
$R_{\mu\nu}-\tfrac{1}{2}g_{\mu\nu}R+\Lambda g_{\mu\nu}\;=\;\frac{8\pi G}{c^{4}}\,T_{\mu\nu}$ (D.148)
The Lorentzian signature (,+,+,+)(-,+,+,+) is a property of the emergent classical geometry, not an input at the fundamental level.
Axiom D.4 (Single Realisation).
The observable universe is one realisation of the stochastic process (D.145). Probability distributions $P(g,\tau)$ describe epistemic uncertainty about the metric configuration, not ontological multiplicity of worlds.
D.1.1 Summary
# Axiom Content Status
A1 Configuration Space 3-metrics on $\Sigma$; superspace $\mathcal{C}$ Canonical GR basis
A2 Stochastic Evolution $\mathrm{d}g=\mathcal{D}\,\mathrm{d}\tau+\ell_{\mathrm{P}}\,\mathrm{d}W$ Central hypothesis
A3 Classical Correspondence Emergent GR in macroscopic limit Empirical consistency
A4 Single Realisation Probability is epistemic Interpretation
Table 7: The four canonical axioms of the superspace diffusion framework.
D.2 Mathematical Preliminaries
The mathematical machinery required for the subsequent development is now established.
D.2.1 Stochastic Calculus
Definition D.1 (Probability Space).
A probability space is a triple $(\Omega,\mathcal{F},\mathbb{P})$ where:
  1. (a)
    $\Omega$ is a sample space (set of outcomes)
  2. (b)
    $\mathcal{F}$ is a $\sigma$-algebra on $\Omega$ (events)
  3. (c)
    $\mathbb{P}:\mathcal{F}\to[0,1]$ is a probability measure with $\mathbb{P}(\Omega)=1$
Definition D.2 (Filtration).
A filtration $\{\mathcal{F}_{t}\}_{t\geq 0}$ is an increasing family of $\sigma$-algebras:
$s\leq t\implies\mathcal{F}_{s}\subseteq\mathcal{F}_{t}\subseteq\mathcal{F}$
representing information available at time $t$.
Definition D.3 (Wiener Process).
A Wiener process (or Brownian motion) $W=\{W_{t}\}_{t\geq 0}$ on $(\Omega,\mathcal{F},\mathbb{P})$ adapted to $\{\mathcal{F}_{t}\}$ is a stochastic process satisfying:
  1. (W1)
    $W_{0}=0$ almost surely
  2. (W2)
    $W_{t}-W_{s}\sim\mathcal{N}(0,t-s)$ for $0\leq s<t$ (Gaussian increments)
  3. (W3)
    For $0\leq t_{1}<t_{2}<\cdots<t_{n}$, the increments $W_{t_{2}}-W_{t_{1}},\ldots,W_{t_{n}}-W_{t_{n-1}}$ are independent
  4. (W4)
    $t\mapsto W_{t}(\omega)$ is continuous for almost all $\omega\in\Omega$
Proposition D.1 (Properties of Wiener Process).
Let $W$ be a Wiener process. Then:
  1. (a)
    $\mathbb{E}[W_{t}]=0$ for all $t\geq 0$
  2. (b)
    $\mathbb{E}[W_{t}^{2}]=t$ for all $t\geq 0$
  3. (c)
    $\mathrm{Var}[W_{t}]=t$ for all $t\geq 0$
  4. (d)
    $W_{t}$ is almost surely nowhere differentiable
  5. (e)
    The quadratic variation is $W]_{t}=t$
Proof.
(a) follows from $W_{t}-W_{0}\sim\mathcal{N}(0,t)$ and $W_{0}=0$. (b) and (c) follow from the variance of the normal distribution. (d) is a classical result of Paley, Wiener, and Zygmund [63]. (e) follows from the definition of quadratic variation and the properties of Gaussian increments. ∎
Definition D.4 (Itô Stochastic Differential Equation).
An Itô stochastic differential equation [1] (SDE) is an equation of the form:
$\mathrm{d}X_{t}=\mu(X_{t},t)\,\mathrm{d}t+\sigma(X_{t},t)\,\mathrm{d}W_{t}$
where $\mu:\mathbb{R}^{n}\times\mathbb{R}_{\geq 0}\to\mathbb{R}^{n}$ is the drift coefficient and $\sigma:\mathbb{R}^{n}\times\mathbb{R}_{\geq 0}\to\mathbb{R}^{n\times m}$ is the diffusion coefficient.
Remark D.2 (Itô product rules).
In Itô calculus, the following product rules hold for the ordering parameter $\tau$ and independent standard Brownian motions $\beta_{m}$, $\beta_{n}$:
$(\mathrm{d}\tau)^{2}=0,\qquad\mathrm{d}\beta_{n}\,\mathrm{d}\tau=0,\qquad\mathrm{d}\beta_{m}\,\mathrm{d}\beta_{n}=\delta_{mn}\,\mathrm{d}\tau.$ (D.149)
These are identities of the calculus, not approximations. The first two express that $\tau$ has zero quadratic variation; the third is the fundamental property that distinguishes stochastic from ordinary calculus.
D.2.2 Differential Geometry
Definition D.5 (Lorentzian Manifold).
A Lorentzian manifold is a pair $(\mathcal{M},g)$ where $\mathcal{M}$ is a smooth manifold and $g$ is a metric tensor of signature (,+,+,+)(-,+,+,+).
Definition D.6 (Space of Metrics).
Let $\mathcal{M}$ be a smooth manifold. The space of Lorentzian metrics on $\mathcal{M}$ is:
$\mathrm{Met}(\mathcal{M})=\{g\in\Gamma(S^{2}T^{*}\mathcal{M}):g\text{ has signature }(-,+,+,+)\}$
Definition D.7 (DeWitt Metric).
The DeWitt metric [31] on the space of spatial metrics is defined by:
$\langle h,k\rangle_{g}=\int_{\Sigma}\sqrt{g}\,G^{ijkl}h_{ij}k_{kl}\,\mathrm{d}^{3}x$
where $\Sigma$ is the spatial manifold, $g$ is the spatial 3-metric (i.e., $g_{ij}$ of Axiom D.1), and:
$G^{ijkl}=\frac{1}{2}(g^{ik}g^{jl}+g^{il}g^{jk}-g^{ij}g^{kl})$
is the DeWitt supermetric (at its standard value $\lambda=-1$). More generally, the one-parameter DeWitt family is:
$\mathcal{G}^{ijkl}(\lambda)=\tfrac{1}{2}(g^{ik}g^{jl}+g^{il}g^{jk})+\lambda\,g^{ij}g^{kl}$ (D.150)
which is positive-definite for $\lambda>-1/3$ [31, 64]. The standard DeWitt value $\lambda=-1$ renders the conformal mode indefinite.
Definition D.8 (Curvature Tensors).
Given a metric $g$, define:
  1. (a)
    The Riemann tensor: $R^{\rho}{}_{\sigma\mu\nu}$
  2. (b)
    The Ricci tensor: $R_{\mu\nu}=R^{\rho}{}_{\mu\rho\nu}$
  3. (c)
    The Ricci scalar: $R=g^{\mu\nu}R_{\mu\nu}$
  4. (d)
    The Einstein tensor: $G_{\mu\nu}=R_{\mu\nu}-\frac{1}{2}Rg_{\mu\nu}$
  5. (e)
    The Kretschmann scalar [65]: $K=R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}$
D.2.3 Planck Units
Definition D.9 (Planck Units).
The Planck units [26] are defined from the fundamental constants $\hbar$ (reduced Planck constant), $G$ (gravitational constant), $c$ (speed of light), and $k_{B}$ (Boltzmann constant):
$\ell_{\mathrm{P}}$ $=\sqrt{\frac{\hbar G}{c^{3}}}\approx 1.616\times 10^{-35}\text{ m}$ (Planck length) (D.151)
$t_{\mathrm{P}}$ $=\sqrt{\frac{\hbar G}{c^{5}}}\approx 5.391\times 10^{-44}\text{ s}$ (Planck time) (D.152)
$m_{\mathrm{P}}$ $=\sqrt{\frac{\hbar c}{G}}\approx 2.176\times 10^{-8}\text{ kg}$ (Planck mass) (D.153)
$E_{\mathrm{P}}$ $=\sqrt{\frac{\hbar c^{5}}{G}}\approx 1.956\times 10^{9}\text{ J}$ (Planck energy) (D.154)
$K_{\mathrm{P}}$ $=\frac{c^{6}}{\hbar^{2}G^{2}}=\frac{1}{\ell_{\mathrm{P}}^{4}}$ (Planck curvature, Kretschmann scale) (D.155)
$\rho_{\mathrm{P}}$ $=\frac{c^{5}}{\hbar G^{2}}=\frac{m_{\mathrm{P}}}{\ell_{\mathrm{P}}^{3}}$ (Planck density) (D.156)
Proposition D.2 (Uniqueness of Planck Length).
The Planck length $\ell_{\mathrm{P}}$ is the unique length scale constructible from $\hbar$, $G$, and $c$ alone.
Proof.
Let $L=\hbar^{\alpha}G^{\beta}c^{\gamma}$ have dimensions of length. The dimensions are:
$L]=[\hbar]^{\alpha}[G]^{\beta}[c]^{\gamma}=(ML^{2}T^{-1})^{\alpha}(M^{-1}L^{3}T^{-2})^{\beta}(LT^{-1})^{\gamma}$
For $[L]=L$:
$M:$ $\alpha-\beta$ $=0$ (D.157)
$L:$ $2\alpha+3\beta+\gamma$ $=1$ (D.158)
$T:$ $-\alpha-2\beta-\gamma$ $=0$ (D.159)
The unique solution is $\alpha=\beta=1/2$, $\gamma=-3/2$, giving $L=\sqrt{\hbar G/c^{3}}=\ell_{\mathrm{P}}$. ∎
D.3 Postulates Determining the Stochastic Evolution
The canonical axioms are stated in Section D.1. Axiom A2 asserts that the metric evolves stochastically with amplitude $\sigma=\ell_{\mathrm{P}}$. This section states the five postulates from which that assertion, and the specific form of the stochastic evolution, are derived. These postulates unpack the physical content of A2; they are not independent axioms but the justification for one.
Postulate D.1 (Existence of Metric Degrees of Freedom).
The gravitational degrees of freedom are described by a spatial metric tensor $g_{ij}$ on a three-manifold $\Sigma$ (as in Axiom D.1).
Postulate D.2 (Quantum Nature of Gravity).
Gravity is a quantum phenomenon. The metric tensor is not deterministic but is subject to quantum fluctuations, and is therefore described by a stochastic process.
Postulate D.3 (Minimality).
The stochastic process governing metric fluctuations is a Wiener process (Definition D.3).
This postulate is not independent: it is a consequence of Postulates P1–P2, Axiom A1, and the Lévy–Khintchine theorem (Theorem D.3 in §D.4.1 below). It is retained as a named postulate for clarity of exposition.
Postulate D.4 (Universality).
The amplitude $\sigma$ of metric fluctuations is a universal constant: the same at all points, at all values of the ordering parameter, and for all components of the metric.
Postulate D.5 (Dimensional Consistency).
The fluctuation amplitude $\sigma$ has dimensions of length: $\sigma]=L$.
Remark D.3.
These five postulates correspond to unpacking canonical Axiom A2 (§D.1). Postulate D.1 is contained in canonical Axiom A1 (Configuration Space). Postulate D.3 (Minimality) is derivable from P1–P2 and A1 via the Lévy–Khintchine theorem (§D.4.1, Theorem D.3); the independent axiomatic content resides in P1, P2, P4, and P5. The fact that $\sigma$ is uniquely determined to be $\ell_{\mathrm{P}}$ (Theorem D.18 below) means that A2 is non-parametric: all coefficients are derived from the axioms.
Construction D.1 (Stochastic Metric Equation).
From Postulates D.1D.5, the stochastic metric equation is constructed:
$\mathrm{d}g_{ij}=\mathcal{D}_{ij}[g]\,\mathrm{d}\tau+\sigma\,\mathrm{d}W_{ij}$ (D.160)
where:
  • $\mathcal{D}_{ij}[g]$ is the drift functional (encoding gravitational dynamics; reducing to the Einstein evolution in the classical limit per canonical Axiom A3)
  • $\mathrm{d}W_{ij}$ is a symmetric tensor-valued Wiener process
  • $\sigma$ is the universal fluctuation amplitude with $\sigma]=L$
  • $\tau$ is the ordering parameter intrinsic to the stochastic process (not a background time coordinate)
This is identical in form to the canonical equation (D.145) of Axiom A2, with $\sigma$ left undetermined pending the uniqueness proof.
Remark D.4 (Dimensional conventions: fluctuation amplitude vs. Itô coefficient).
Postulate D.5 asserts $\sigma]=L$. This refers to the fluctuation amplitude, the physical scale of metric perturbations, not to the Itô diffusion coefficient of a standard-Wiener SDE. In the standard Itô convention (Definition D.4), the diffusion coefficient of a variable $X$ with $X]=L^{p}T^{q}M^{r}$ has dimensions $\sigma_{\mathrm{It\hat{o}}}]=[X]\cdot T^{-1/2}$. For a dimensionless variable such as $g_{ij}$, this would give $\sigma_{\mathrm{It\hat{o}}}]=T^{-1/2}$.
Equation (D.160) is consistent because $\mathrm{d}W_{ij}$ is not a standard unit-covariance Wiener process. Its covariance is specified by the operator $K_{ijkl}$, and the eigenvalues $\{\lambda_{n}\}$ of $K$ (Lemma D.9) absorb the dimensions needed to keep $\mathrm{d}g_{ij}$ dimensionless: for each mode, $\mathrm{d}g_{n}=\mathcal{D}_{n}\,\mathrm{d}\tau+\ell_{\mathrm{P}}\sqrt{\lambda_{n}}\,\mathrm{d}\beta_{n}$, with $\mathrm{d}\beta_{n}$ a standard (dimensionless-increment) Brownian motion. The mode variance is $\ell_{\mathrm{P}}^{2}\lambda_{n}\,\mathrm{d}\tau$, and the eigenvalues carry $\lambda_{n}]=L^{-2}T^{-1}$ so that $\ell_{\mathrm{P}}^{2}\lambda_{n}\,\mathrm{d}\tau]=1$.
The physically transparent check is the RMS dimensionless perturbation over one Planck time. Setting $\sigma=\alpha\ell_{\mathrm{P}}$ and evaluating at the Planck scale (as in §D.7), the dimensionless metric perturbation is $h_{\mathrm{rms}}=\alpha$, confirming that A2’s convention produces a well-defined dimensionless fluctuation.
If one rewrites the A2 noise in terms of a standard (unit-covariance) Wiener process $\mathrm{d}B_{ij}$ via the Karhunen–Loève expansion, the effective Itô diffusion coefficient for a length-valued variable becomes $\sigma_{L}=\ell_{\mathrm{P}}/\sqrt{t_{\mathrm{P}}}$, with $\sigma_{L}]=L\cdot T^{-1/2}$. For the dimensionless metric perturbation $h$, the corresponding coefficient is $\sigma_{h}=1/\sqrt{t_{\mathrm{P}}}$, with $\sigma_{h}]=T^{-1/2}$. Appendix C derives these values independently from the fluctuation-dissipation theorem (Corollary 5.2) and proves the equivalence:
$\sigma_{L}\cdot\sqrt{t_{\mathrm{P}}}=\frac{\ell_{\mathrm{P}}}{\sqrt{t_{\mathrm{P}}}}\cdot\sqrt{t_{\mathrm{P}}}=\ell_{\mathrm{P}}.$ (D.161)
That is, the RMS length fluctuation over one Planck time is exactly $\ell_{\mathrm{P}}$, regardless of which convention is used. The three equivalent descriptions are:
Description Symbol Dimensions Value
Fluctuation amplitude (A2) $\sigma=\ell_{\mathrm{P}}$ $L$ $1.616\times 10^{-35}$ m
[…]
Itô form (metric perturbation) $\sigma_{h}=1/\sqrt{t_{\mathrm{P}}}$ $T^{-1/2}$ $4.307\times 10^{21}$ s-1/2
[…]
A2 uses the fluctuation amplitude convention because it cleanly separates the universal scale ($\ell_{\mathrm{P}}$) from the geometric structure ($K_{ijkl}$). This convention is used throughout the superspace diffusion framework.
D.4 Mathematical Characterisation of the Stochastic Process
Postulate D.3 (Minimality) asserts that the stochastic process is a Wiener process. This subsection demonstrates that Minimality is not an independent assumption but a consequence of the remaining postulates and Axiom A1, thereby reducing the axiomatic content of the framework. It then establishes the connections between the Fokker–Planck evolution on superspace and two standard quantisation programmes: canonical quantum gravity (via the Wheeler–DeWitt equation) and the gravitational path integral (via the Martin–Siggia–Rose formalism).
D.4.1 Uniqueness of the Wiener Structure
Theorem D.3 (Wiener Process from Continuity and Independent Increments).
Let $\{X_{\tau}\}_{\tau\geq 0}$ be a stochastic process on $\mathcal{C}=\mathrm{Riem}(\Sigma)/\mathrm{Diff}(\Sigma)$ satisfying:
  1. (i)
    Stationary independent increments: $X_{\tau_{2}}-X_{\tau_{1}}$ is independent of $X_{\tau_{1}}-X_{\tau_{0}}$ for $\tau_{0}<\tau_{1}<\tau_{2}$, and the distribution of $X_{\tau+s}-X_{\tau}$ depends only on $s$.
  2. (ii)
    Path continuity: The sample paths $\tau\mapsto g_{ij}(\tau)$ are continuous in the topology induced by the DeWitt supermetric on $\mathcal{C}$.
Then $X_{\tau}$ is a Gaussian process with covariance proportional to $\tau$. That is, $X_{\tau}$ is a (possibly tensor-valued) Wiener process.
Proof.
The Lévy–Khintchine theorem [66, 68] classifies all processes with stationary independent increments (Lévy processes) on $\mathbb{R}^{n}$. The characteristic exponent of any such process $X_{\tau}$ takes the form:
$\log\mathbb{E}\bigl[\mathrm{e}^{i\langle\xi,X_{\tau}\rangle}\bigr]=\tau\Bigl(i\langle b,\xi\rangle-\tfrac{1}{2}\langle\xi,A\xi\rangle+\int_{\mathbb{R}^{n}\setminus\{0\}}\bigl(\mathrm{e}^{i\langle\xi,y\rangle}-1-i\langle\xi,y\rangle\,\mathbf{1}_{|y|\leq 1}\bigr)\,\nu(\mathrm{d}y)\Bigr),$ (D.162)
where $b\in\mathbb{R}^{n}$ is the drift, $A$ is a positive semi-definite covariance matrix (the Gaussian part), and $\nu$ is the Lévy measure characterising the jump structure. The decomposition is unique.
The process decomposes pathwise into three independent components [66]:
$X_{\tau}=b\tau+\sqrt{A}\,W_{\tau}+J_{\tau},$ (D.163)
where $W_{\tau}$ is a standard Wiener process and $J_{\tau}$ is a pure-jump process (compound Poisson for finite $\nu$, or a limit of compensated Poisson processes for infinite $\nu$).
Condition (ii) requires that the sample paths be continuous. A Lévy process has continuous paths if and only if $\nu\equiv 0$ [66, 68], i.e. the jump component vanishes identically. With $\nu=0$, Eq. (D.162) reduces to:
$\log\mathbb{E}[\mathrm{e}^{i\langle\xi,X_{\tau}\rangle}]=\tau\bigl(i\langle b,\xi\rangle-\tfrac{1}{2}\langle\xi,A\xi\rangle\bigr),$ (D.164)
which is the characteristic function of a Gaussian process with drift $b$ and covariance $A\tau$. Setting $b=\mathcal{D}_{ij}$ (the drift functional) and $A=\ell_{\mathrm{P}}^{2}\,\mathcal{G}^{ijkl}$ (the DeWitt covariance) recovers Axiom A2. ∎
Remark D.5 (Physical necessity of path continuity).
Condition (ii) is not imposed for mathematical convenience. It is a physical consequence of Axiom A1. The configuration space $\mathcal{C}$ consists of smooth Riemannian metrics on $\Sigma$. A jump discontinuity $g_{ij}(\tau)\to g_{ij}(\tau)+\Delta g_{ij}$ with $|\Delta g_{ij}|$ finite would generically:
  1. (a)
    violate positive-definiteness of $g_{ij}$ (taking a valid Riemannian metric outside $\mathrm{Riem}(\Sigma)$),
  2. (b)
    produce instantaneous topology change if the jump crosses a degenerate metric ($\det g=0$), and
  3. (c)
    generate infinite curvature ($\delta K\sim\Delta g/(\Delta\tau)^{2}\to\infty$ as $\Delta\tau\to 0$).
All three are excluded by the requirement that $g_{ij}(\tau)\in\mathrm{Riem}(\Sigma)$ for all $\tau$. Path continuity is therefore a consequence of A1, and Theorem D.3 establishes that the Wiener process is the unique Lévy process consistent with A1. Postulate D.3 (Minimality) is thus derivable from the remaining structure and may be regarded as a theorem rather than an independent axiom.
Remark D.6 (Markov property).
The Markov property (Requirement R2 of Theorem B.3) follows from the independent-increments structure established by Theorem D.3. A deeper justification is that the ordering parameter $\tau$ is not physical time: it is the filtration index of the stochastic process. Memory (non-Markovianity) requires a notion of “how long ago,” which presupposes a temporal metric. Since physical time is emergent from the quadratic variation (Theorem D.10), it is not available at the level where the process is defined. Imposing non-Markovian structure on the $\tau$-evolution would circularly presuppose the temporal structure that the framework constructs.
At the effective level, even if the fundamental process possessed $t_{\mathrm{P}}$-scale temporal correlations, Donsker’s invariance principle [69] guarantees convergence to a Wiener process after averaging over $N\sim(L/\ell_{\mathrm{P}})$ correlation lengths, with corrections of order $N^{-1/2}\sim(\ell_{\mathrm{P}}/L)$. For any observable scale $L\gg\ell_{\mathrm{P}}$, non-Markovian corrections are unmeasurably small (see SI-4, Appendix D for the explicit bound $|\mathcal{C}-1|\leq 0.13$ for the cosmological case).
Remark D.7 (Spatial locality of the noise).
The $\delta$-correlated spatial structure $\mathbb{E}[\mathrm{d}W_{ij}(x)\,\mathrm{d}W_{kl}(y)]\propto\delta(x,y)\,\mathrm{d}\tau$ follows from two considerations. First, cluster decomposition: spacelike-separated metric fluctuations must be statistically independent, since spatial correlations in the noise would transmit information acausally. Second, the absence of new length scales: any spatial correlation length $\xi$ in the noise would introduce a dimensionful parameter beyond $\{\hbar,G,c\}$. By the uniqueness of the Planck length (Proposition D.2), $\xi=\alpha\,\ell_{\mathrm{P}}$ for some dimensionless $\alpha$. At scales $L\gg\ell_{\mathrm{P}}$, the spatial central limit theorem averages over $(L/\xi)^{3}$ independent correlation volumes, rendering the effective noise $\delta$-correlated to precision $({\ell_{\mathrm{P}}}/{L})^{3/2}$. The locality assumption is thus both theoretically motivated and observationally invisible when relaxed.
D.4.2 Connection to the Wheeler–DeWitt Equation
The Fokker–Planck evolution on superspace admits a formal correspondence with the Wheeler–DeWitt equation of canonical quantum gravity. This correspondence, rooted in the Guerra–Ruggiero stochastic quantisation programme [47, 70, 6, 24], clarifies the relationship between the stochastic framework and conventional approaches to quantum gravity.
Theorem D.4 (Stochastic–Quantum Correspondence).
Let $P[g,\tau]$ satisfy the Fokker–Planck equation on superspace:
$\frac{\partial P}{\partial\tau}=-\frac{\delta}{\delta g_{ij}}\bigl(\mathcal{D}_{ij}^{\mathrm{GR}}\,P\bigr)+\frac{\ell_{\mathrm{P}}^{2}}{2}\,\mathcal{G}^{ijkl}\frac{\delta^{2}P}{\delta g_{ij}\,\delta g_{kl}}.$ (D.165)
Define the formal analytic continuation:
$\tau\;\longrightarrow\;-\frac{it}{\hbar},\qquad P[g,\tau]\;\longrightarrow\;\Psi[g,t],\qquad\frac{\ell_{\mathrm{P}}^{2}}{2}\;\longrightarrow\;-\frac{\hbar^{2}}{2M_{\mathrm{P}}^{2}},$ (D.166)
where $M_{\mathrm{P}}=\sqrt{\hbar c/G}$ is the Planck mass. Then $\Psi[g,t]$ satisfies:
$i\hbar\frac{\partial\Psi}{\partial t}=\left[-\frac{\hbar^{2}}{2M_{\mathrm{P}}^{2}}\,\mathcal{G}^{ijkl}\frac{\delta^{2}}{\delta g_{ij}\,\delta g_{kl}}+V_{\mathrm{grav}}[g]\right]\Psi,$ (D.167)
where $V_{\mathrm{grav}}[g]$ incorporates the potential term from the Hamiltonian constraint. In the limit where $\Psi$ is time-independent ($\partial\Psi/\partial t=0$), Eq. (D.167) reduces to the Wheeler–DeWitt equation $\hat{H}\Psi=0$.
Proof.
Under the substitutions (D.166), the Fokker–Planck equation (D.165) transforms term by term. The left-hand side gives:
$\frac{\partial P}{\partial\tau}\;\longrightarrow\;\frac{\partial\Psi}{\partial(-it/\hbar)}=\frac{i\hbar^{-1}}{\phantom{i}}\,\frac{\partial\Psi}{\partial t}\;\Longrightarrow\;\frac{\partial P}{\partial\tau}\to-\frac{i}{\hbar}\frac{\partial\Psi}{\partial t}.$ (D.168)
The diffusion term transforms as:
$\frac{\ell_{\mathrm{P}}^{2}}{2}\,\mathcal{G}^{ijkl}\frac{\delta^{2}P}{\delta g_{ij}\,\delta g_{kl}}\;\longrightarrow\;-\frac{\hbar^{2}}{2M_{\mathrm{P}}^{2}}\,\mathcal{G}^{ijkl}\frac{\delta^{2}\Psi}{\delta g_{ij}\,\delta g_{kl}}.$ (D.169)
The drift term contributes the gravitational potential through the identity (valid in the ADM decomposition):
[…]
where $V_{\mathrm{grav}}[g]=-\sqrt{g}\,{}^{(3)}\!R$ is the gravitational potential (the spatial scalar curvature density) and the ordering terms depend on the operator-ordering prescription. Assembling:
$\frac{i}{\hbar}\frac{\partial\Psi}{\partial t}=-\frac{\hbar^{2}}{2M_{\mathrm{P}}^{2}}\,\mathcal{G}^{ijkl}\frac{\delta^{2}\Psi}{\delta g_{ij}\,\delta g_{kl}}+V_{\mathrm{grav}}\,\Psi+\text{(ordering)},$ (D.171)
which, upon multiplication by $i\hbar$, yields Eq. (D.167). ∎
Remark D.8 (Status and limitations of the correspondence).
The substitution (D.166) is a formal analytic continuation, in the same sense as the standard Wick rotation relating Euclidean and Lorentzian path integrals. Three caveats apply:
  1. (a)
    The ordering parameter $\tau$ is not physical time; its analytic continuation to $it/\hbar$ is a mathematical operation, not a physical identification.
  2. (b)
    The operator-ordering ambiguity in Eq. (D.167) (the conformal factor ordering in the DeWitt supermetric) is not resolved by the correspondence. In the stochastic framework, the Itô convention provides a definite prescription; under Wick rotation, this maps to a specific operator ordering, but the physical significance of this choice remains an open question.
  3. (c)
    The correspondence applies at the level of the generator, not the solutions: the Fokker–Planck solution $P\geq 0$ (a probability density) and the Wheeler–DeWitt solution $\Psi$ (a wave functional) inhabit different function spaces.
The correspondence establishes that the stochastic framework is not isolated from canonical quantum gravity but is connected to it by the same analytic structure (Wick rotation) that relates Euclidean and Lorentzian formulations of quantum field theory. The stochastic framework may be regarded as the “Euclidean” version of quantum gravity in which the probability density $P$ plays the role of the Euclidean wave functional. The relationship parallels the Nelson–Guerra–Ruggiero connection between stochastic mechanics and quantum mechanics [6, 47, 71].
D.4.3 Path-Integral Formulation and Structural Consistency
The Fokker–Planck evolution (D.165) admits an equivalent path-integral representation via the Martin–Siggia–Rose / Janssen–de Dominicis (MSR/JdD) formalism [72, 73, 74]. This representation provides three structural results: a path integral for stochastic gravity, a proof that the classical constraint surface is the saddle-point locus, and a theorem on foliation independence.
Theorem D.5 (MSR Path Integral for Stochastic Gravity).
The generating functional for correlation functions of the stochastic metric evolution (Axiom A2) is:
$Z[\mathcal{J}]=\int\mathcal{D}g\;\mathcal{D}\tilde{g}\;\exp\Bigl(-S_{\mathrm{MSR}}[g,\tilde{g}]+\int\mathcal{J}^{ij}\,g_{ij}\Bigr),$ (D.172)
where $\tilde{g}_{ij}$ is the auxiliary response field and the MSR action is:
$S_{\mathrm{MSR}}[g,\tilde{g}]=\int_{0}^{T}\!\mathrm{d}\tau\int_{\Sigma}\!\mathrm{d}^{3}x\;\sqrt{g}\;\left[\tilde{g}_{ij}\Bigl(\dot{g}_{ij}-\mathcal{D}_{ij}^{\mathrm{GR}}[g]\Bigr)-\frac{\ell_{\mathrm{P}}^{2}}{2}\,\mathcal{G}^{ijkl}\,\tilde{g}_{ij}\,\tilde{g}_{kl}\right].$ (D.173)
Here $\dot{g}_{ij}\equiv\partial g_{ij}/\partial\tau$, and the integral over $\Sigma$ is with respect to the induced volume form.
Proof.
This is the standard MSR construction [72, 73, 74] applied to the Langevin equation (D.200) on superspace. The derivation proceeds in three steps.
Step 1. Write the probability of a path $\{g_{ij}(\tau)\}_{\tau\in[0,T]}$ as a functional $\delta$-function enforcing the SDE:
$\mathbb{P}[\{g\}]\propto\int\mathcal{D}\eta\;\delta\!\left[\dot{g}_{ij}-\mathcal{D}_{ij}^{\mathrm{GR}}-\ell_{\mathrm{P}}\,\eta_{ij}\right]\exp\!\left(-\frac{1}{2}\int_{0}^{T}\!\mathrm{d}\tau\int_{\Sigma}\!\mathrm{d}^{3}x\;\sqrt{g}\;\bigl(\mathcal{G}^{-1}\bigr)_{ijkl}\,\eta_{ij}\,\eta_{kl}\right),$ (D.174)
where $\eta_{ij}$ is the white-noise field with covariance $\mathcal{G}^{ijkl}$.
Step 2. Represent the $\delta$-function via its Fourier transform with respect to the auxiliary field $\tilde{g}_{ij}$:
$\delta[\dot{g}-\mathcal{D}^{\mathrm{GR}}-\ell_{\mathrm{P}}\eta]=\int\mathcal{D}\tilde{g}\;\exp\!\left(i\!\int_{0}^{T}\!\mathrm{d}\tau\int_{\Sigma}\!\mathrm{d}^{3}x\;\sqrt{g}\;\tilde{g}_{ij}\bigl(\dot{g}_{ij}-\mathcal{D}_{ij}^{\mathrm{GR}}-\ell_{\mathrm{P}}\,\eta_{ij}\bigr)\right).$ (D.175)
Step 3. Integrate out $\eta_{ij}$ (Gaussian integral). After the standard shift $\tilde{g}\to-i\tilde{g}$ (the Janssen–de Dominicis rotation) to render the action real, the result is Eq. (D.173). The Jacobian of the transformation is field-independent and absorbed into the normalisation.
The equivalence between the MSR path integral and the Fokker–Planck equation is a standard result: the $n$-point functions $\langle g_{i_{1}j_{1}}(\tau_{1})\cdots g_{i_{n}j_{n}}(\tau_{n})\rangle$ computed from $Z[\mathcal{J}]$ coincide with the moments of $P[g,\tau]$ solving Eq. (D.165[75, 76]. ∎
Lemma D.6 (Classical Constraint Surface as Saddle Point).
The saddle-point equations of $S_{\mathrm{MSR}}$ are:
$\frac{\delta S_{\mathrm{MSR}}}{\delta\tilde{g}_{ij}}=0\;\;\Longrightarrow\;\;\dot{g}_{ij}=\mathcal{D}_{ij}^{\mathrm{GR}}[g]+\ell_{\mathrm{P}}^{2}\,\mathcal{G}^{ijkl}\,\tilde{g}_{kl},$ (D.176)
$\frac{\delta S_{\mathrm{MSR}}}{\delta g_{ij}}=0\;\;\Longrightarrow\;\;\dot{\tilde{g}}_{ij}=-\frac{\delta\mathcal{D}_{kl}^{\mathrm{GR}}}{\delta g_{ij}}\,\tilde{g}_{kl}+\text{(metric variation terms)}.$ (D.177)
The classical sector is the locus $\tilde{g}_{ij}=0$, on which Eq. (D.176) reduces to the Einstein flow $\dot{g}_{ij}=\mathcal{D}_{ij}^{\mathrm{GR}}[g]$, i.e., the Hamiltonian and momentum constraints of general relativity. The MSR action evaluates to $S_{\mathrm{MSR}}=0$ on this locus.
Proof.
Setting $\tilde{g}_{ij}=0$ in Eq. (D.176) gives $\dot{g}_{ij}=\mathcal{D}_{ij}^{\mathrm{GR}}[g]$, which is the ADM evolution equation under the Einstein flow. In the Hamiltonian formulation of GR, this evolution is generated by the Hamiltonian constraint $\mathcal{H}[g,\pi]\approx 0$ and momentum constraint $\mathcal{H}_{a}[g,\pi]\approx 0$; the constraint surface is precisely the set of initial data $(g_{ij},\pi^{ij})$ for which the Einstein evolution is self-consistent [77, 53].
On the classical locus $(\tilde{g}_{ij}=0,\;\dot{g}_{ij}=\mathcal{D}_{ij}^{\mathrm{GR}})$, both terms in Eq. (D.173) vanish: the first because $\dot{g}-\mathcal{D}^{\mathrm{GR}}=0$, and the second because $\tilde{g}=0$. Hence $S_{\mathrm{MSR}}=0$ on the classical solution, confirming that the constraint surface is a global minimum of the action. ∎
Corollary D.7 (Stochastic Excursions from the Constraint Surface).
Fluctuations around the saddle point are controlled by $\ell_{\mathrm{P}}$. Expanding $S_{\mathrm{MSR}}$ to quadratic order about the classical solution $(\bar{g},0)$:
$S_{\mathrm{MSR}}^{(2)}=\int_{0}^{T}\!\mathrm{d}\tau\int_{\Sigma}\!\mathrm{d}^{3}x\;\sqrt{\bar{g}}\;\left[\delta\tilde{g}_{ij}\bigl(\delta\dot{g}_{ij}-\hat{L}_{ij}^{\;kl}\,\delta g_{kl}\bigr)-\frac{\ell_{\mathrm{P}}^{2}}{2}\,\mathcal{G}^{ijkl}\,\delta\tilde{g}_{ij}\,\delta\tilde{g}_{kl}\right],$ (D.178)
where $\hat{L}_{ij}^{\;kl}=\delta\mathcal{D}_{ij}^{\mathrm{GR}}/\delta g_{kl}|_{\bar{g}}$ is the linearised Einstein operator. The Gaussian integral over $\delta\tilde{g}$ produces:
$\langle(\delta g_{ij})^{2}\rangle\sim\ell_{\mathrm{P}}^{2}\cdot\tau,$ (D.179)
confirming that stochastic excursions from the constraint surface have amplitude $O(\ell_{\mathrm{P}})$ per unit ordering time. For an observation at scale $L$, the dimensionless departure from the constraint surface is:
$\frac{\delta g}{g}\sim\frac{\ell_{\mathrm{P}}}{L},$ (D.180)
which is $\sim 10^{-35}$ for $L=1$ m. The Fokker–Planck stationary measure concentrates on the classical constraint surface in the limit $\ell_{\mathrm{P}}/L\to 0$.
Remark D.9 (Well-posedness of the Fokker–Planck evolution).
The Fokker–Planck equation (D.165) is an infinite-dimensional parabolic PDE on $\mathcal{C}$. In the minisuperspace reduction (SI-2, §2.4), it becomes a standard finite-dimensional parabolic equation whose well-posedness follows from classical theory [44, 43]. For the full infinite-dimensional theory, the MSR path integral (Theorem D.5) provides the operative definition of the dynamics: correlation functions, expectation values, and the stationary measure are defined by the functional integral (D.172) and computed by standard perturbative and non-perturbative methods. This parallels the situation in quantum field theory, where the path integral defines the theory and the functional Schrödinger equation serves as a formal (often ill-defined) equivalent. The MSR path integral thus defines the theory operationally, making functional-analytic well-posedness of the infinite-dimensional generator a mathematical refinement rather than a physical obstruction. A rigorous functional-analytic treatment of the infinite-dimensional Fokker–Planck generator (essential self-adjointness, domain characterisation, spectral theory) is deferred to future work; the results of this paper depend only on the minisuperspace reduction and the formal properties of the full generator.
Theorem D.8 (Lapse Independence of Physical Observables).
Let $N(x)$ be a lapse function on $\Sigma$, generating a reparametrisation $\mathrm{d}\tau^{\prime}=N(x)\,\mathrm{d}\tau$ of the ordering parameter (the “many-fingered time” of ADM gravity). Then:
  1. (i)
    The MSR action transforms covariantly:
    $S_{\mathrm{MSR}}^{(N)}[g,\tilde{g}]=\int_{0}^{T}\!\mathrm{d}\tau\int_{\Sigma}\!\mathrm{d}^{3}x\;\sqrt{g}\;N(x)\left[\tilde{g}_{ij}\Bigl(\frac{1}{N}\dot{g}_{ij}-\mathcal{D}_{ij}^{\mathrm{GR}}\Bigr)-\frac{\ell_{\mathrm{P}}^{2}}{2}\,\mathcal{G}^{ijkl}\,\tilde{g}_{ij}\,\tilde{g}_{kl}\right].$ (D.181)
  2. (ii)
    For any diffeomorphism-invariant observable $\mathcal{O}[g]$, the expectation value is lapse-independent:
    $\langle\mathcal{O}[g]\rangle_{N_{1}}=\langle\mathcal{O}[g]\rangle_{N_{2}}$ (D.182)
    for any two lapse choices $N_{1}(x)$, $N_{2}(x)$.
Proof.
Part (i). Under $\mathrm{d}\tau^{\prime}=N(x)\,\mathrm{d}\tau$, the stochastic evolution becomes:
$\mathrm{d}g_{ij}=N(x)\,\mathcal{D}_{ij}^{\mathrm{GR}}\,\mathrm{d}\tau+\ell_{\mathrm{P}}\sqrt{N(x)}\,\mathrm{d}W_{ij}^{\prime},$ (D.183)
where $\mathrm{d}W^{\prime}_{ij}$ is a Wiener process with respect to the reparametrised filtration. This follows from the Dambis–Dubins–Schwarz theorem (Lemma B.1), which generalises to spatially varying $N(x)$ because the noise is spatially $\delta$-correlated (Remark D.7): at each spatial point, the DDS reparametrisation applies independently.
Repeating the MSR construction of Theorem D.5 with the reparametrised SDE yields the action (D.181).
Part (ii). The transformation $N_{1}\to N_{2}$ is implemented by the change of variables:
$\tilde{g}_{ij}\;\longrightarrow\;\frac{N_{1}(x)}{N_{2}(x)}\,\tilde{g}_{ij},\qquad\mathrm{d}\tau\;\longrightarrow\;\frac{N_{2}(x)}{N_{1}(x)}\,\mathrm{d}\tau.$ (D.184)
The Jacobian of this transformation in the path integral factorises into a product over spatial points (by ultralocality of the noise covariance):
$\mathcal{J}=\prod_{x\in\Sigma}\left(\frac{N_{1}(x)}{N_{2}(x)}\right)^{n_{\mathrm{dof}}/2},$ (D.185)
where $n_{\mathrm{dof}}=6$ is the number of independent components of $g_{ij}$ at each point. This Jacobian is field-independent (it depends only on $N_{1},N_{2}$, not on $g$ or $\tilde{g}$) and therefore cancels between numerator and denominator in the computation of:
$\langle\mathcal{O}\rangle_{N}=\frac{\displaystyle\int\mathcal{D}g\,\mathcal{D}\tilde{g}\;\mathcal{O}[g]\,\mathrm{e}^{-S_{\mathrm{MSR}}^{(N)}}}{\displaystyle\int\mathcal{D}g\,\mathcal{D}\tilde{g}\;\mathrm{e}^{-S_{\mathrm{MSR}}^{(N)}}}.$ (D.186)
The cancellation holds for any gauge-invariant $\mathcal{O}[g]$ because the field-independent Jacobian contributes the same multiplicative factor to both integrals.
Physically, the lapse $N(x)$ controls how fast the stochastic clock ticks at each spatial point, but not the statistical content of the process. Different lapse choices correspond to different parametrisations of the same diffusion on $\mathcal{C}$. The quadratic variation of the process, which defines emergent time (Theorem D.10), absorbs the lapse into the emergent temporal metric, producing the standard ADM relation $\mathrm{d}t_{\mathrm{proper}}=N\,\mathrm{d}t$ in the semiclassical limit (Lemma D.11). ∎
Remark D.10 (Distributional versus pathwise refoliation invariance).
Classical general relativity enjoys pathwise refoliation invariance: the same spacetime can be sliced by different foliations, and the hypersurface deformation algebra guarantees consistency. In the stochastic framework, pathwise refoliation invariance is broken: different lapse choices $N(x)$ generate different noise realisations $\mathrm{d}W^{\prime}_{ij}=\sqrt{N(x)}\,\mathrm{d}W_{ij}$, so the same stochastic history cannot be reconstructed from two different foliations.
Theorem D.8 establishes the stochastic analogue: distributional refoliation invariance. All gauge-invariant observables have lapse-independent expectation values. The physical content of the theory: encoded in the Fokker–Planck measure on $\mathcal{C}$ and the correlation functions computed from $Z[\mathcal{J}]$, is foliation-independent. Individual stochastic trajectories are foliation-dependent, just as individual Feynman paths are gauge-dependent in ordinary quantum field theory; only the path-integral measure is gauge-invariant.
This is the natural generalisation of the classical result: refoliation invariance is promoted from a statement about individual solutions to a statement about the statistical ensemble, consistent with the epistemic interpretation of probability (Axiom A4).
Remark D.11 (Connection to the Euclidean gravitational path integral).
Under the Wick rotation of Theorem D.4, the MSR action acquires an imaginary part. For the saddle-point sector ($\tilde{g}=0$), the Wick-rotated MSR generating functional reduces to:
$Z_{\mathrm{Eucl}}=\int\mathcal{D}g\;\exp\!\left(-\frac{1}{\ell_{\mathrm{P}}^{2}}\,S_{\mathrm{EH}}^{\mathrm{Eucl}}[g]\right),$ (D.187)
where $S_{\mathrm{EH}}^{\mathrm{Eucl}}$ is the Euclidean Einstein–Hilbert action and the prefactor $1/\ell_{\mathrm{P}}^{2}=M_{\mathrm{P}}^{2}c/\hbar$ plays the role of the inverse gravitational coupling. This establishes that the stochastic framework, viewed through its MSR path integral, contains the Euclidean gravitational path integral (as employed in the Hartle–Hawking programme [78]) as its saddle-point approximation.
The stochastic framework thus interpolates between three standard approaches to quantum gravity:
  1. (a)
    Canonical: via the Guerra–Ruggiero correspondence with the Wheeler–DeWitt equation (Theorem D.4);
  2. (b)
    Euclidean path integral: via the MSR generating functional at the saddle point (this remark);
  3. (c)
    Semiclassical GR: via the classical limit $\ell_{\mathrm{P}}/L\to 0$ (Theorem D.22).
The MSR formulation provides the unifying structure from which each is recovered as a limit or a formal transformation.
D.5 The Emergence of Physical Time
The ordering parameter $\tau$ in A2 requires careful interpretation. Diffusion is ordinarily defined as random motion as a function of time; if time itself is claimed to emerge from diffusion, one must specify diffusion with respect to what.
The stochastic process on $\mathcal{C}$ is defined as a probability measure on paths in configuration space (the Wiener measure construction), which requires only a $\sigma$-algebra and filtration, not a physical time coordinate. The index parameter $\tau$ is an abstract ordering label for the filtration $\{\mathcal{F}_{\tau}\}$.
Lemma D.9 (Spectral Decomposition of the Covariance Operator).
Let $\Sigma$ be a compact Riemannian 3-manifold. The covariance operator $K_{ijkl}[g](x,y)$ of Axiom D.2, being a continuous, symmetric, positive-definite kernel on $L^{2}(\mathrm{Sym}^{2}(T^{*}\Sigma))$ (Remark D.1), admits a Mercer decomposition:
$K_{ijkl}(x,y)\;=\;\sum_{n=1}^{\infty}\lambda_{n}\,e^{(n)}_{ij}(x)\,e^{(n)}_{kl}(y),\qquad\lambda_{n}>0,$ (D.188)
where $\{e^{(n)}_{ij}\}$ is a complete orthonormal system of eigentensors and the series converges uniformly on $\Sigma\times\Sigma$. The Wiener process admits the Karhunen–Loève expansion $\mathrm{d}W_{ij}(x)=\sum_{n}\sqrt{\lambda_{n}}\,e^{(n)}_{ij}(x)\,\mathrm{d}\beta_{n}$, with $\{\beta_{n}\}$ independent standard Brownian motions. In this basis, the quadratic form evaluates mode-by-mode:
$\langle\mathrm{d}g,\,K^{-1}\,\mathrm{d}g\rangle\;=\;\sum_{n=1}^{\infty}\frac{1}{\lambda_{n}}\,(\mathrm{d}g_{n})^{2}\;=\;\sum_{n=1}^{\infty}\frac{1}{\lambda_{n}}\bigl(\mathcal{D}_{n}\,\mathrm{d}\tau+\ell_{\mathrm{P}}\sqrt{\lambda_{n}}\,\mathrm{d}\beta_{n}\bigr)^{2}.$ (D.189)
By the Itô product rule (Remark D.2), $(\mathrm{d}\tau)^{2}=0$, $\mathrm{d}\beta_{n}\,\mathrm{d}\tau=0$, and $\mathrm{d}\beta_{m}\,\mathrm{d}\beta_{n}=\delta_{mn}\,\mathrm{d}\tau$. Therefore:
$\langle\mathrm{d}g,\,K^{-1}\,\mathrm{d}g\rangle\;=\;\ell_{\mathrm{P}}^{2}\sum_{n=1}^{\infty}1\cdot\mathrm{d}\tau\;=\;\ell_{\mathrm{P}}^{2}\,n_{\mathrm{eff}}\,\mathrm{d}\tau,$ (D.190)
where $n_{\mathrm{eff}}$ is the effective number of excited modes. The drift $\mathcal{D}_{n}$ contributes identically zero, as a consequence of the Itô product rules.
Proof.
The Mercer decomposition follows from Mercer’s theorem [79] applied to the continuous, symmetric, positive-definite kernel $K$ on the compact domain $\Sigma$ (Remark D.1 establishes positive-definiteness; compactness is assumed for $\Sigma$). The Karhunen–Loève expansion is standard. For the Itô evaluation: expanding the square in (D.189) yields three terms per mode. The first, $\mathcal{D}_{n}^{2}\,(\mathrm{d}\tau)^{2}/\lambda_{n}$, vanishes because $(\mathrm{d}\tau)^{2}=0$ in Itô calculus. The cross-term $2\ell_{\mathrm{P}}\,\mathcal{D}_{n}\,\mathrm{d}\beta_{n}\,\mathrm{d}\tau/\sqrt{\lambda_{n}}$ vanishes because $\mathrm{d}\beta_{n}\,\mathrm{d}\tau=0$. The third term is $\ell_{\mathrm{P}}^{2}\lambda_{n}\,(\mathrm{d}\beta_{n})^{2}/\lambda_{n}=\ell_{\mathrm{P}}^{2}\,\mathrm{d}\tau$, using $(\mathrm{d}\beta_{n})^{2}=\mathrm{d}\tau$. Summing over modes gives (D.190). ∎
Remark D.12 (Regularisation and the mode count).
On a compact manifold, the mode sum $n_{\mathrm{eff}}=\sum_{n=1}^{\infty}1$ is formally divergent. Two natural regularisations apply:
  1. 1.
    Planck-scale cutoff: modes with wavelength below $\ell_{\mathrm{P}}$ are unphysical (Axiom D.2 defines fluctuations at scale $\ell_{\mathrm{P}}$, not below it). On a manifold of volume $V$, this gives $n_{\mathrm{eff}}\sim V/\ell_{\mathrm{P}}^{3}$, finite.
  2. 2.
    Minisuperspace reduction: restricting to spatially homogeneous modes gives $n_{\mathrm{eff}}=1$ (FLRW) or $n_{\mathrm{eff}}=3$ (Bianchi IX), and the functional is rigorously well-defined with no regularisation needed.
In either case, the normalisation factor $n_{\mathrm{eff}}$ is absorbed into the definition of $t(\tau)$: physical time is defined as
$t(\tau)\;=\;\frac{1}{\ell_{\mathrm{P}}^{2}\,n_{\mathrm{eff}}}\int_{0}^{\tau}\langle\mathrm{d}g,\,K^{-1}\,\mathrm{d}g\rangle,$ (D.191)
so that $t(\tau)=\tau$ exactly (by Lemma D.9). All physical predictions depend on $t$, not on $n_{\mathrm{eff}}$ separately; the normalisation convention identifies the abstract ordering parameter with the physical clock.
The emergent time functional is therefore:
$t(\tau)\;=\;\frac{1}{\ell_{\mathrm{P}}^{2}\,n_{\mathrm{eff}}}\int_{0}^{\tau}\langle\mathrm{d}g,\,K^{-1}\,\mathrm{d}g\rangle\;=\;\tau,$ (D.192)
where $n_{\mathrm{eff}}$ is the (regularised) effective mode count (Remark D.12) and the final equality follows from Lemma D.9. This identifies the abstract ordering parameter $\tau$ with physical time.
Theorem D.10 (Emergence of Physical Time).
The functional $t(\tau)$ defined in (D.192) satisfies:
  1. 1.
    Strict monotone increase in $\tau$ (from positivity of $K$),
  2. 2.
    Additivity under concatenation of stochastic paths (from independent increments),
  3. 3.
    Independence of coordinate parametrisation on $\Sigma$ (from covariance of the inner product),
  4. 4.
    Exact independence of the drift $\mathcal{D}_{ij}$ (by the Itô product rule, Lemma D.9).
In the macroscopic limit $L\gg\ell_{\mathrm{P}}$, $t$ coincides with proper time along semiclassical trajectories (Lemma D.11).
Proof.
Monotonicity: by Lemma D.9, the quadratic form evaluates to $\ell_{\mathrm{P}}^{2}\,n_{\mathrm{eff}}\,\mathrm{d}\tau$ for any increment $\mathrm{d}\tau>0$, which is strictly positive since $n_{\mathrm{eff}}\geq 1$ and $\ell_{\mathrm{P}}>0$. (Equivalently: each independent Brownian motion $\beta_{n}$ has nonzero increments on every interval almost surely, and $K^{-1}$ is positive-definite by Remark D.1.) Additivity follows from the independent-increments property of the Wiener process: $t(\tau_{1}+\tau_{2})=t(\tau_{1})+t_{[\tau_{1},\tau_{1}+\tau_{2}]}$. Coordinate independence follows from the tensorial transformation properties of $g_{ij}$, $K_{ijkl}$, and the inner product under spatial diffeomorphisms. Drift independence is proven in Lemma D.9: all drift-dependent terms vanish identically by the Itô product rules $(\mathrm{d}\tau)^{2}=0$ and $\mathrm{d}\beta_{n}\,\mathrm{d}\tau=0$. The classical limit is established in Lemma D.11 below. ∎
Lemma D.11 (Classical Limit: Proper Time Recovery).
In the ADM decomposition of a semiclassical history, the emergent time $t$ coincides with proper time $t_{\mathrm{proper}}$ up to corrections of $O(\ell_{\mathrm{P}}^{2}/L^{2})$.
Proof.
In the ADM formalism [77], the spatial metric evolves as:
$\frac{\partial g_{ij}}{\partial t_{\mathrm{proper}}}=-2NK_{ij}^{\mathrm{ext}}+\nabla_{i}N_{j}+\nabla_{j}N_{i},$ (D.193)
where $N$ is the lapse function, $N_{i}$ the shift vector, and $K_{ij}^{\mathrm{ext}}$ the extrinsic curvature (superscript distinguishing it from the covariance operator $K_{ijkl}$). In the semiclassical regime ($L\gg\ell_{\mathrm{P}}$), the drift dominates the stochastic evolution, so $\mathcal{D}_{ij}\,\mathrm{d}\tau\approx\mathrm{d}g_{ij}$. Identifying with (D.193):
$\mathcal{D}_{ij}=\frac{\mathrm{d}t_{\mathrm{proper}}}{\mathrm{d}\tau}\bigl(-2NK_{ij}^{\mathrm{ext}}+\nabla_{i}N_{j}+\nabla_{j}N_{i}\bigr).$ (D.194)
But Theorem D.10 established that $t(\tau)=\tau$. In the semiclassical limit, matching the drift-dominated evolution to the ADM evolution requires:
$\frac{\mathrm{d}t_{\mathrm{proper}}}{\mathrm{d}\tau}=\frac{\mathrm{d}t_{\mathrm{proper}}}{\mathrm{d}t}=N+O(\ell_{\mathrm{P}}^{2}/L^{2}),$ (D.195)
which is the standard ADM relation $\mathrm{d}t_{\mathrm{proper}}=N\,\mathrm{d}t$ for observers following the foliation. For a comoving observer ($N_{i}=0$, $N=1$), $t_{\mathrm{proper}}=t$. Stochastic corrections enter at $O(\ell_{\mathrm{P}}^{2}/L^{2})\sim 10^{-70}$ for astrophysical scales. ∎
The emergent time theorem addresses the “problem of time” in canonical quantum gravity [80, 81]. The Wheeler-DeWitt equation $\hat{H}\Psi=0$ [31] lacks a time variable; here, time is constructed from the quadratic variation of the stochastic process on superspace. Property (4) (exact drift independence) implies that the rate of the emergent time functional is determined entirely by the diffusion (amplitude $\sigma=\ell_{\mathrm{P}}$), with the deterministic gravitational dynamics contributing nothing. The uniqueness of $\sigma$ (Theorem D.18 below) therefore determines the rate of the emergent time functional.
D.6 Uniqueness of Drift and Covariance
Theorem D.12 (Uniqueness of Drift and Diffusion).
Within the class of ultralocal, second-order, diffeomorphism-covariant stochastic evolutions on $\mathcal{C}$ satisfying dimensional consistency and classical correspondence (Axiom D.3), the dynamics is uniquely determined (up to gauge transformations):
$\mathrm{d}g_{ij}=\mathcal{D}_{ij}^{\mathrm{GR}}[g]\,\mathrm{d}\tau+\ell_{\mathrm{P}}\,\mathrm{d}W_{ij}$ (D.196)
where:
  1. 1.
    $\mathcal{D}_{ij}^{\mathrm{GR}}$ reduces in the macroscopic limit to the Einstein flow on spatial metrics,
  2. 2.
    $K_{ijkl}$ is proportional to a positive-definite member of the DeWitt family $\mathcal{G}^{ijkl}(\lambda)$ with $\lambda>-1/3$.
Proof sketch.
For the diffusion coefficient: an ultralocal, diffeomorphism-covariant, positive-definite symmetric operator on $\mathrm{Sym}^{2}(T^{*}\Sigma)$ must be constructed from $g_{ij}$ alone. The unique such operator (up to two free constants) is the DeWitt family $\mathcal{G}^{ijkl}=\tfrac{1}{2}(g^{ik}g^{jl}+g^{il}g^{jk})+\lambda\,g^{ij}g^{kl}$, where positive-definiteness requires $\lambda>-1/3$ (Definition D.7). At the standard DeWitt value $\lambda=-1$, the conformal mode is indefinite; the stochastic process is restricted to the positive-definite branch. The overall scale is fixed by $\sigma=\ell_{\mathrm{P}}$.
For the drift: the requirement that $\mathcal{D}_{ij}^{\mathrm{GR}}$ be a local, diffeomorphism-covariant functional of $g_{ij}$ and its derivatives, at most second order, that reproduces the Einstein evolution in the classical limit, invokes the Lovelock-type classification [82, 83] of such tensors adapted to the ADM formalism on 3-metrics. In three spatial dimensions, the unique such choice (up to cosmological constant) is the extrinsic-curvature evolution implied by the Einstein constraint and evolution equations. ∎
Remark D.13.
Theorem D.12 confirms the claim in Remark D.1: the covariance operator $K_{ijkl}$ is a member of the positive-definite DeWitt family. This closes the logical chain: A2 states $K$ is (semi-)definite; the uniqueness of drift and diffusion forces strict positive-definiteness; and this in turn guarantees the strict monotonicity of the emergent time functional (Theorem D.10).
D.7 Uniqueness of the Fluctuation Amplitude
The central result is now proved: the fluctuation amplitude is uniquely determined. With the emergent time theorem established (Section D.5), the physical significance is immediate: the amplitude $\sigma$ determines the rate of the emergent time functional, with the drift contributing identically zero. The value of $\sigma$ is not a tunable parameter; it is determined by self-consistency.
Lemma D.13 (Dimensional Constraint).
Let $\sigma$ be a universal constant with $\sigma]=L$, constructed from the fundamental constants of quantum gravity $\{\hbar,G,c\}$. Then $\sigma=\alpha\ell_{\mathrm{P}}$ for some dimensionless constant $\alpha$.
Proof.
By Proposition D.2, $\ell_{\mathrm{P}}$ is the unique length scale constructible from $\hbar$, $G$, and $c$. Any other length with these dimensions must be a dimensionless multiple of $\ell_{\mathrm{P}}$. ∎
Definition D.10 (Dimensionless Metric Perturbation).
Let $\delta g_{ij}$ denote the stochastic increment of the spatial metric over one ordering-parameter step $\Delta\tau$. The dimensionless metric perturbation at length scale $L$ is:
$h\;=\;\frac{|\delta g_{ij}|}{g_{ij}}\;\sim\;\frac{\sigma\sqrt{\Delta\tau}}{L},$ (D.197)
where $\sigma$ is the fluctuation amplitude (Postulate D.5) and $L$ is the characteristic length scale of the background geometry. The RMS perturbation accumulated over one Planck time ($\Delta\tau=t_{\mathrm{P}}$) at the Planck scale ($L=\ell_{\mathrm{P}}$) is:
$h_{\mathrm{rms}}=\frac{\sigma\sqrt{t_{\mathrm{P}}}}{\ell_{\mathrm{P}}}=\alpha,$ (D.198)
where $\sigma=\alpha\ell_{\mathrm{P}}/\sqrt{t_{\mathrm{P}}}$. The bounds in the subsequent lemmas apply to $h_{\mathrm{rms}}$ as defined here; specifically, inequalities such as $h_{\mathrm{rms}}\leq 1$ hold in the mean-square sense ($\langle h^{2}\rangle^{1/2}\leq 1$), not as almost-sure bounds on individual realisations.
Lemma D.14 (Curvature Fluctuation).
Let the metric fluctuate with amplitude $\sigma$ over a length scale $L$. The dimensionless metric perturbation is $h\sim\sigma/L$, and the induced fluctuation in the Kretschmann scalar is:
$\delta K\sim\frac{h^{2}}{L^{4}}=\frac{\sigma^{2}}{L^{6}}.$
Proof.
Write the metric as $g_{ij}=\bar{g}_{ij}+h_{ij}$, where $h_{ij}$ is a dimensionless perturbation. A metric fluctuation of amplitude $\sigma$ (with $\sigma]=L$) over a length scale $L$ produces a dimensionless perturbation of magnitude $h\sim\sigma/L$.
The Riemann tensor involves second derivatives of the metric:
$R^{\rho}{}_{\sigma\mu\nu}\sim\partial^{2}h\sim\frac{h}{L^{2}},$
with dimensions $R]=L^{-2}$. The Kretschmann scalar $K=R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}$ is quadratic in the Riemann tensor:
$\delta K\sim(\delta R)^{2}\sim\left(\frac{h}{L^{2}}\right)^{2}=\frac{h^{2}}{L^{4}},$
with $\delta K]=L^{-4}$ as required. Substituting $h=\sigma/L$:
$\delta K\sim\frac{\sigma^{2}}{L^{6}}.\qed$
Lemma D.15 (Self-Consistency Bound).
For the stochastic metric theory to be self-consistent, the dimensionless metric perturbation must not exceed unity at any scale:
$h_{\mathrm{rms}}\leq 1\quad\text{for all }L\geq\ell_{\mathrm{P}}.$
Equivalently, the curvature fluctuation must not exceed the Planck curvature:
$\delta K\leq K_{\mathrm{P}}=\frac{1}{\ell_{\mathrm{P}}^{4}}\quad\text{for all }L\geq\ell_{\mathrm{P}}.$
Proof.
The metric is $g_{ij}=\bar{g}_{ij}+h_{ij}$. For the perturbative decomposition to be meaningful, $|h|<|\bar{g}|$, i.e., $h_{\mathrm{rms}}\lesssim O(1)$. If $h_{\mathrm{rms}}>1$, the perturbation exceeds the background metric and the perturbative decomposition becomes inconsistent. The equivalent curvature statement follows from $\delta K\sim h^{2}/L^{4}$ (Lemma D.14): at $L=\ell_{\mathrm{P}}$ with $h=1$, $\delta K=1/\ell_{\mathrm{P}}^{4}=K_{\mathrm{P}}$. ∎
Lemma D.16 (Lower Bound from Quantum Resolution).
For metric fluctuations to resolve classical singularities (where $K\to\infty$), the fluctuation-induced curvature must reach the Planck curvature at the Planck scale:
$\delta K\geq K_{\mathrm{P}}\quad\text{at }L=\ell_{\mathrm{P}}.$
Proof.
Classical general relativity predicts singularities where $K\to\infty$ [84, 85]. For quantum effects to regularise these singularities, the fluctuation-induced curvature must become comparable to the classical curvature at the scale where quantum gravity dominates. The Planck curvature $K_{\mathrm{P}}=1/\ell_{\mathrm{P}}^{4}$ is that scale. If $\delta K\ll K_{\mathrm{P}}$ at $L=\ell_{\mathrm{P}}$, fluctuations are insufficient to modify the singularity structure, and the theory does not regularise the classical divergence. ∎
Lemma D.17 (Critical Damping at the Planck Scale).
A metric fluctuation mode of frequency $\omega$ has a gravitational self-coupling rate $\gamma(\omega)=t_{\mathrm{P}}^{2}\,\omega^{3}$ (from the equivalence principle, Planck relation, and mass-energy equivalence; see Appendix C, Theorem 8.1). The critical damping condition $\gamma(\omega_{*})=\omega_{*}$, at which a mode’s dissipation rate matches its oscillation frequency, has a unique solution:
$\omega_{*}=\frac{1}{t_{\mathrm{P}}}=\omega_{\mathrm{P}}.$ (D.199)
At this frequency, $\gamma(\omega_{\mathrm{P}})=1/t_{\mathrm{P}}$, and the corresponding diffusion coefficient gives $\alpha=1$ exactly.
Proof.
The condition $\gamma(\omega_{*})=\omega_{*}$ gives $t_{\mathrm{P}}^{2}\,\omega_{*}^{3}=\omega_{*}$. For $\omega_{*}>0$, dividing both sides by $\omega_{*}$ yields $\omega_{*}^{2}=1/t_{\mathrm{P}}^{2}$, hence $\omega_{*}=1/t_{\mathrm{P}}=\omega_{\mathrm{P}}$. This is an exact algebraic identity.
At $\omega=\omega_{\mathrm{P}}$, the self-coupling rate is $\gamma(\omega_{\mathrm{P}})=t_{\mathrm{P}}^{2}\cdot t_{\mathrm{P}}^{-3}=1/t_{\mathrm{P}}$. By the massless cancellation theorem (Appendix C, Theorem 3.1), the velocity-variable diffusion coefficient is $\sigma_{v}=c\sqrt{\gamma}=c/\sqrt{t_{\mathrm{P}}}$, corresponding to $\alpha=1$ in the proper-distance convention $\sigma_{L}=\alpha\,\ell_{\mathrm{P}}/\sqrt{t_{\mathrm{P}}}$.
Physically, the critical damping frequency separates two regimes: for $\omega<\omega_{\mathrm{P}}$, metric modes propagate coherently (classical spacetime); for $\omega>\omega_{\mathrm{P}}$, modes are overdamped (quantum foam). The stochastic theory describes fluctuations at this boundary, fixing the characteristic scale to $\alpha=1$. ∎
Theorem D.18 (Uniqueness of $\sigma$ within the A1–A4 Framework).
Under Axioms A1–A4 (§D.1) and Postulates P1–P2, P4–P5 (§D.3), with P3 (Minimality) derived from the Lévy–Khintchine theorem (Theorem D.3) and the gravitational self-coupling rate $\gamma(\omega)=t_{\mathrm{P}}^{2}\omega^{3}$ derived in Appendix C (Theorem C.16), the fluctuation amplitude is uniquely determined:
$\boxed{\sigma=\ell_{\mathrm{P}}}$
Proof.
By Lemma D.13, $\sigma=\alpha\ell_{\mathrm{P}}$ for some dimensionless $\alpha>0$. Three independent arguments determine $\alpha$:
Argument 1 (Upper bound). At scale $L=\ell_{\mathrm{P}}$, the dimensionless metric perturbation is $h=\sigma/L=\alpha\ell_{\mathrm{P}}/\ell_{\mathrm{P}}=\alpha$. By Lemma D.15, $h\leq 1$ is required for metric self-consistency, hence $\alpha\leq 1$. This is a strict bound: for $\alpha>1$, the perturbation exceeds the background metric over one Planck time, invalidating the perturbative decomposition.
Equivalently, via Lemma D.14: $\delta K\sim h^{2}/\ell_{\mathrm{P}}^{4}=\alpha^{2}/\ell_{\mathrm{P}}^{4}=\alpha^{2}K_{\mathrm{P}}$. Requiring $\delta K\leq K_{\mathrm{P}}$ gives $\alpha^{2}\leq 1$, hence $\alpha\leq 1$.
Argument 2 (Critical damping). By Lemma D.17, the critical damping condition $\gamma(\omega_{*})=\omega_{*}$ yields $\omega_{*}=\omega_{\mathrm{P}}=1/t_{\mathrm{P}}$ exactly, with the corresponding diffusion coefficient giving $\alpha=1$. This is an algebraic identity, independent of Argument 1.
Argument 3 (Lower bound). By Lemma D.14, at $L=\ell_{\mathrm{P}}$: $\delta K=\alpha^{2}/\ell_{\mathrm{P}}^{4}=\alpha^{2}K_{\mathrm{P}}$. By Lemma D.16, singularity resolution requires $\delta K\geq K_{\mathrm{P}}$, hence $\alpha^{2}\geq 1$, giving $\alpha\geq 1$.
[…]
Therefore $\sigma=\ell_{\mathrm{P}}$. ∎
Corollary D.19 (The Fundamental Equation).
The stochastic metric equation (D.160) takes the unique form:
$\boxed{\mathrm{d}g_{ij}=\mathcal{D}_{ij}[g]\,\mathrm{d}\tau+\ell_{\mathrm{P}}\,\mathrm{d}W_{ij}}$ (D.200)
This equation contains non-parametric predictions: all quantities are determined by the fundamental constants $\hbar$, $G$, and $c$.
Remark D.14 (Notation for the fundamental equation).
Equation (D.200) is written in the canonical notation of Axiom A2: spatial indices $ij$ on the 3-manifold $\Sigma$, the drift functional $\mathcal{D}_{ij}[g]$, and the abstract ordering parameter $\tau$. Three identifications, all proved within this appendix, allow this to be rewritten:
  1. 1.
    The ordering parameter $\tau$ is identified with physical time $t$ via the emergent time functional (Theorem D.10: $t(\tau)=\tau$).
  2. 2.
    The drift $\mathcal{D}_{ij}[g]$ is identified with the Einstein flow $\mathcal{D}_{ij}^{\mathrm{GR}}[g]$ (Theorem D.12).
  3. 3.
    In the semiclassical limit, emergent time $t$ reduces to proper time $t_{\mathrm{proper}}$ (Lemma D.11).
With these identifications, the equation may equivalently be written:
$\mathrm{d}g_{ij}=\mathcal{D}_{ij}^{\mathrm{GR}}[g]\,\mathrm{d}t+\ell_{\mathrm{P}}\,\mathrm{d}W_{ij}$ (D.201)
The classical limit section (§D.8) uses this identified form.
Equation (D.200) can now be read in its full context. The drift $\mathcal{D}_{ij}[g]$ is uniquely the Einstein flow (Theorem D.12). The diffusion amplitude $\ell_{\mathrm{P}}$ is uniquely determined (Theorem D.18). The ordering parameter $\tau$ is identified with physical time $t$ via the emergent time functional (Theorem D.10). The covariance operator is fixed to the positive-definite DeWitt family (Theorem D.12). No undetermined quantities remain.
D.8 The Classical Limit
The following proves that general relativity emerges in the appropriate limit, thereby verifying the consistency of the uniqueness result with Axiom A3. The notation here uses the identified form (D.201): physical time $t$ (= $\tau$ by Theorem D.10) and the drift $\mathcal{D}_{ij}^{\mathrm{GR}}[g]$ (= the Einstein flow by Theorem D.12).
Definition D.11 (Relative Fluctuation).
For a system of characteristic size $L$, the relative fluctuation is:
$\epsilon(L)=\frac{\sigma}{L}=\frac{\ell_{\mathrm{P}}}{L}$
Lemma D.20 (Mean Evolution).
The mean metric evolution satisfies classical Einstein equations:
$\mathbb{E}[\mathrm{d}g_{ij}]=\mathcal{D}_{ij}^{\mathrm{GR}}[\mathbb{E}[g]]\,\mathrm{d}t+O(\ell_{\mathrm{P}}^{2})$
Proof.
Taking expectations of the identified form (D.201):
$\mathbb{E}[\mathrm{d}g_{ij}]=\mathbb{E}[\mathcal{D}_{ij}^{\mathrm{GR}}[g]]\,\mathrm{d}t+\ell_{\mathrm{P}}\,\mathbb{E}[\mathrm{d}W_{ij}]$
Since $\mathbb{E}[\mathrm{d}W_{ij}]=0$ (property of Wiener process):
$\mathbb{E}[\mathrm{d}g_{ij}]=\mathbb{E}[\mathcal{D}_{ij}^{\mathrm{GR}}[g]]\,\mathrm{d}t$
For small fluctuations, $\mathbb{E}[\mathcal{D}_{ij}^{\mathrm{GR}}[g]]\approx\mathcal{D}_{ij}^{\mathrm{GR}}[\mathbb{E}[g]]+O(\mathrm{Var}[g])=\mathcal{D}_{ij}^{\mathrm{GR}}[\mathbb{E}[g]]+O(\ell_{\mathrm{P}}^{2})$. ∎
Lemma D.21 (Variance Scaling).
For a single mode of the stochastic process (Remark D.12), the variance of the dimensionless metric perturbation accumulated over emergent time $T$ is:
$\mathrm{Var}[h]=\frac{T}{t_{\mathrm{P}}},$
since $\sigma_{h}=1/\sqrt{t_{\mathrm{P}}}$ is the per-mode Itô coefficient for the dimensionless metric perturbation (Remark D.4). For a macroscopic observable probing scale $L$, the relevant quantity is not the bare per-mode variance but the relative fluctuation $\epsilon=\ell_{\mathrm{P}}/L$ (Definition D.11), which characterises the dimensionless perturbation amplitude at that scale.
Theorem D.22 (Classical Limit).
In the limit $L\gg\ell_{\mathrm{P}}$ (equivalently $\epsilon\to 0$), the identified form (D.201) reduces to classical general relativity:
$\boxed{\lim_{\ell_{\mathrm{P}}/L\to 0}\left(\mathrm{d}g_{ij}=\mathcal{D}_{ij}^{\mathrm{GR}}\,\mathrm{d}t+\ell_{\mathrm{P}}\,\mathrm{d}W_{ij}\right)=\mathrm{d}g_{ij}=\mathcal{D}_{ij}^{\mathrm{GR}}\,\mathrm{d}t}$ (D.202)
Proof.
The proof proceeds by showing that stochastic corrections to the Einstein equations are controlled by $\epsilon=\ell_{\mathrm{P}}/L\to 0$.
Step 1: The perturbation amplitude vanishes. By Definition D.11 and Lemma D.14, the dimensionless metric perturbation induced by stochastic fluctuations at scale $L$ is $h\sim\ell_{\mathrm{P}}/L=\epsilon$. For $L\gg\ell_{\mathrm{P}}$, $h\ll 1$: the metric is well-approximated by a smooth background $\bar{g}_{ij}$ with perturbations of order $\epsilon$.
Step 2: The mean evolution is the Einstein flow. By Lemma D.20, $\mathbb{E}[\mathrm{d}g_{ij}]=\mathcal{D}_{ij}^{\mathrm{GR}}[\mathbb{E}[g]]\,\mathrm{d}t+O(\epsilon^{2})$. The $O(\epsilon^{2})$ correction arises from the nonlinearity of $\mathcal{D}^{\mathrm{GR}}$: expanding $\mathcal{D}^{\mathrm{GR}}[g]$ about $\mathbb{E}[g]$ gives a correction proportional to $\mathrm{Var}[g]\sim\epsilon^{2}$.
Step 3: Observable fluctuations vanish. The curvature fluctuation at scale $L$ is $\delta K\sim\epsilon^{2}/L^{4}$ (Lemma D.14 with $h=\epsilon$), while the classical background curvature for a system of scale $L$ is $K_{\mathrm{cl}}\sim 1/L^{4}$. The ratio is $\delta K/K_{\mathrm{cl}}\sim\epsilon^{2}=\ell_{\mathrm{P}}^{2}/L^{2}\to 0$. All curvature observables converge to their classical values.
As $\epsilon\to 0$, the stochastic term becomes negligible, the mean evolution dominates, and fluctuations become unobservable. The limiting equation is Einstein’s equation of general relativity. ∎
Corollary D.23 (Determinism as Approximation).
Classical determinism is not fundamental but emergent. It is an excellent approximation for $L\gg\ell_{\mathrm{P}}$, which includes all macroscopic and astronomical phenomena. For an object of size $L=1$ m, the relative fluctuation is $\epsilon=\ell_{\mathrm{P}}/L=1.6\times 10^{-35}$; quantum gravitational corrections are suppressed by a factor of $10^{-35}$, consistent with the empirical success of classical general relativity across all observed scales.
D.9 Summary of the Logical Chain
The complete deductive structure of this appendix is:
  1. 1.
    Axiom A1 (Configuration Space, §D.1): The arena is Wheeler superspace $\mathcal{C}=\mathrm{Riem}(\Sigma)/\mathrm{Diff}(\Sigma)$.
  2. 2.
    Axiom A2 (Stochastic Evolution, §D.1): The metric undergoes a stochastic process with $\sigma=\ell_{\mathrm{P}}$ and (semi-)definite covariance $K_{ijkl}$.
  3. 3.
    Postulates P1–P5D.3): Unpack the content of A2 into five postulates: existence of metric degrees of freedom, quantum nature of gravity, minimality (Wiener process), universality of $\sigma$, and dimensional consistency ($\sigma]=L$). P3 (Minimality) is derived from P1–P2 and A1 via the Lévy–Khintchine theorem (Theorem D.3), reducing the independent axiomatic content to four postulates.
  4. 4.
    Theorem D.3 (Wiener Uniqueness, §D.4.1): The Lévy–Khintchine classification, combined with path continuity (a consequence of A1), uniquely selects the Wiener process. The Markov property and spatial locality follow (Remarks D.6, D.7).
  5. 5.
    Proposition D.2 (Uniqueness of $\ell_{\mathrm{P}}$): Dimensional analysis proves $\ell_{\mathrm{P}}$ is the unique length constructible from $\{\hbar,G,c\}$.
  6. 6.
    Remark D.1 (Positive-definiteness of $K$): Three independent requirements force $K_{ijkl}$ to be strictly positive-definite.
  7. 7.
    Lemma D.9 (Mercer Decomposition): The covariance operator admits a spectral decomposition; the quadratic variation evaluates to $\ell_{\mathrm{P}}^{2}\,n_{\mathrm{eff}}\,\mathrm{d}\tau$ with the drift contributing exactly zero.
  8. 8.
    Theorem D.10 (Emergent Time): Physical time is the quadratic variation of the stochastic process. It is strictly monotone, additive, coordinate-independent, and exactly drift-independent. The rate of the emergent time functional is determined by $\sigma=\ell_{\mathrm{P}}$.
  9. 9.
    Lemma D.11 (Proper Time Recovery): In the semiclassical limit, the emergent time reduces to proper time via the ADM relation $\mathrm{d}t_{\mathrm{proper}}=N\,\mathrm{d}t$.
  10. 10.
    Theorem D.12 (Drift and Covariance Uniqueness): Diffeomorphism covariance and classical correspondence uniquely fix the drift to the Einstein flow and the covariance to the positive-definite DeWitt family.
  11. 11.
    Lemma D.13 (Dimensional Constraint): $\sigma=\alpha\ell_{\mathrm{P}}$ for dimensionless $\alpha>0$.
  12. 12.
    Lemma D.15 (Self-Consistency): The dimensionless perturbation $h=\sigma/L$ must satisfy $h\leq 1$; equivalently $\delta K\leq K_{\mathrm{P}}=1/\ell_{\mathrm{P}}^{4}$. At $L=\ell_{\mathrm{P}}$: $\alpha\leq 1$ (strict bound).
  13. 13.
    Lemma D.17 (Critical Damping): The self-coupling rate $\gamma(\omega)=t_{\mathrm{P}}^{2}\omega^{3}$ equals $\omega$ at exactly $\omega_{*}=1/t_{\mathrm{P}}$, giving $\alpha=1$ (algebraic identity; Appendix C).
  14. 14.
    Lemma D.16 (Quantum Resolution): Singularity resolution requires $\delta K\geq K_{\mathrm{P}}$ at $L=\ell_{\mathrm{P}}$. Since $\delta K=\alpha^{2}K_{\mathrm{P}}$, this gives $\alpha\geq 1$.
  15. 15.
    Theorem D.18 (Uniqueness): $\alpha\leq 1\;\wedge\;\alpha=1\;\wedge\;\alpha\geq 1\;\Longrightarrow\;\alpha=1$, giving $\boxed{\sigma=\ell_{\mathrm{P}}}$. Three independent arguments yield the same value.
  16. 16.
    Corollary D.19 (Fundamental Equation): The stochastic metric equation $\mathrm{d}g_{ij}=\mathcal{D}_{ij}[g]\,\mathrm{d}\tau+\ell_{\mathrm{P}}\,\mathrm{d}W_{ij}$ contains non-parametric, first-principles predictions: drift, covariance, amplitude, and time are all uniquely determined.
  17. 17.
    Theorem D.4 (Stochastic–Quantum Correspondence, §D.4.2): Formal Wick rotation maps the Fokker–Planck equation on superspace to the Wheeler–DeWitt equation of canonical quantum gravity.
  18. 18.
    Theorem D.5 (MSR Path Integral, §D.4.3): The Fokker–Planck evolution admits an equivalent path-integral representation; the classical constraint surface is the saddle-point locus (Lemma D.6), with $O(\ell_{\mathrm{P}})$ stochastic excursions (Corollary D.7).
  19. 19.
    Theorem D.8 (Lapse Independence, §D.4.3): Physical observables are independent of the foliation (lapse) choice; distributional refoliation invariance replaces pathwise refoliation invariance (Remark D.10).
  20. 20.
    Theorem D.22 (Classical Limit, §D.8): Stochastic corrections vanish as $\epsilon=\ell_{\mathrm{P}}/L\to 0$, recovering Einstein’s equations. The physical time $t$ in the proof is the emergent time functional of Theorem D.10, and the drift is the Einstein flow of Theorem D.12.
  21. 21.
    Axiom A4 (Single Realisation, §D.1): Constrains interpretation: probability is epistemic, not ontological.
In summary, the fluctuation amplitude $\sigma=\ell_{\mathrm{P}}$ is the unique value consistent with dimensional analysis, curvature self-consistency, critical damping, and singularity resolution. The drift is uniquely the Einstein flow; the covariance is uniquely the positive-definite DeWitt family; physical time is uniquely the quadratic variation of the diffusion. The Wiener structure itself is uniquely selected by the Lévy–Khintchine theorem. The framework is non-parametric: all coefficients are derived from first principles. Classical general relativity is recovered as an emergent approximation for $L\gg\ell_{\mathrm{P}}$. The framework connects to canonical quantum gravity via the Guerra–Ruggiero correspondence (Theorem D.4), to the gravitational path integral via the MSR formalism (Theorem D.5), and exhibits distributional refoliation invariance under changes of foliation (Theorem D.8). The notation $\sigma=\ell_{\mathrm{P}}$ (fluctuation amplitude, $\sigma]=L$) used throughout this appendix is equivalent to the proper-distance Itô coefficient $\sigma_{L}=\ell_{\mathrm{P}}/\sqrt{t_{\mathrm{P}}}$ derived independently in Appendix C (Remark D.4).